### v1

# The Role of Privacy Guarantees in Voluntary Donation of Private Data for Altruistic Goals

Abstract
Voluntary donation of private information for altruistic purposes, such as advancing research, is common. However, concerns about data misuse and leakage may deter individuals from donating their information. While prior research has indicated that Privacy Enhancement Technologies (PETs) can alleviate these concerns, the extent to which these techniques influence willingness to donate data remains unclear.
This study conducts a vignette survey ($N=485$) to examine people’s willingness to donate medical data for developing new treatments under four privacy guarantees: data expiration, anonymization, use restriction, and access control. The study explores two mechanisms for verifying these guarantees: self-auditing and expert auditing, and evaluates the impact on two types of data recipient entities: for-profit and non-profit institutions.
Our findings reveal that the type of entity collecting data strongly influences respondents’ privacy expectations, which in part influence their willingness to donate data. Respondents have such high expectations of the privacy provided by non-profit entities that explicitly stating the privacy protections provided makes little adjustment to those expectations. In contrast, statements about privacy bring respondents’ expectations of the privacy provided by for-profit entities nearly in-line with non-profit expectations. We highlight the risks of these respective results as well as the need for future research to better align technical community and end-user perceptions about the effectiveness of auditing PETs and to effectively set expectations about the efficacy of PETs in the face of end-user concerns about data breaches.
1 Introduction
The altruistic analysis of personal health data can bring tremendous societal benefit. For instance, the public health sector could leverage individuals’ medical history and health data to analyze epidemiological patterns, support disease surveillance and risk prediction, diagnose rare and emerging diseases, and improve the development of new treatments [18, 5, 14, 42, 61, 106]. The potential of these applications entails economic benefits: for example, the EU Commission believes that innovations supporting data sharing and data availability may save billions of Euros in the EU health sector [28]. Consequently, the need for personal data in the health sector is ever-increasing [117, 18] — especially with the recent boom of large language models (LLMs), whose capabilities depend on the data they are fed.
Personal health data is already being collected by care providers [14, 61] and personal mobile apps and devices [42]; then, third parties may get access to this data through data sharing agreements that regulate data donation [106]. However, people are often unwilling to donate such sensitive personal data due to privacy concerns [88, 112, 97, 42, 61]. For example, people may be uncomfortable with hospital policies that allow data sharing agreements with third parties without the data owners’ explicit consent [64]. In general, privacy concerns significantly influence (i) people’s adoption of, engagement with, and ability to benefit from technology [15, 89] as well as (ii) the regulations suggested by a country’s legislators [27].
In an effort to foster data collection and analysis while protecting the privacy of data contributors, there has been an increasing focus on deploying privacy-enhancing technologies (PETs) for data storage, data processing, and machine learning (ML) training. For example, encrypted databases such as CryptDB [87] or BlindSeer [81] store user data in encrypted form; hardware-secured platforms [8, 12, 16] additionally rely on Trusted Execution Environments (TEEs) to process data in a secure hardware compartment, so that the confidentiality and integrity of this data (and the code processing it) is guaranteed even during computation; secure Multi-Party Computation (MPC) platforms [44, 25] allow distinct entities to pool their encrypted data and jointly analyze it, without any single entity learning other entities’ data, even indirectly; finally, systems combining all these approaches [119, 32] expand the set of guarantees achievable by any single approach. In addition, many PETs implement differential privacy (DP) [7] to prevent the results of statistical computations on a given sensitive dataset from leaking the presence (or absence) of a specific user’s data.
Despite growing research and deployment of PETs, people feel less control over their data than ever [9]. Arguably, the effort towards strengthening and widening the spectrum of PETs’ privacy-preserving guarantees is insufficient if people’s privacy concerns are not mitigated in practice.
Prior work [30, 2, 6, 10, 31, 33, 3, 118, 120, 75] has sought to meaningfully explain such guarantees—or the technologies that implement them—to the people they aim to protect. We build on this work to evaluate whether such guarantees effectively alleviate the people’s privacy concerns, so that they would be more willing to donate their personal data. While much of the prior work that shares our aim has studied specific, singular technologies or guarantees (often DP [30, 19, 75] or end-to-end encryption (E2EE) [33, 11, 2]), we focus on the more general security and privacy guarantees recurring in PETs:
  1. PG(1):
    Anonymization [74, 96]: data is not linkable to its owner (as defined in [24]);
  2. PG(2):
    Access control [42, 51, 76, 105]: data is accessible only by authorized people, specified in the data collection agreement;
  3. PG(3):
    Data expiration [86, 93, 82]: data is discarded or inaccessible after a given expiration time, specified in the data collection agreement;
  4. PG(4):
    Purpose restriction [32, 20]: data is only used for the stated collection purpose, specified in the data collection agreement;
Additionally, and distinct from prior work, we investigate the role of auditing or verification processes increasingly introduced to produce public auditing results for PETs, specifically:
  1. AG(1):
    Expert auditing [65, 113]: engaging an aggregator-selected external advisory board to audit the system to verify that the PET functions as described in the data collection agreement;
  2. AG(2):
    Self auditing [49, 23]: granting anyone, including donors or external advisors appointed by them, the ability to perform such audits.
Specifically, we investigate the role of these general PETs guarantees and audits of them in shaping privacy expectations and data sharing intentions in the context of medical data donation, asking:
  1. RQ(1):
    How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?
  2. RQ(2):
    How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?
To address these questions, we conduct a vignette survey ($n=485$) following best practice methodology established by prior work [30, 2, 120]. Each survey respondent is presented with a hypothetical opportunity to donate their health data to help develop a treatment for a specific chronic disease, along with information on how their data will be protected (PG(1)–PG(4) enforced (or not) by AG(1)–AG(2)). Furthermore, we isolate the effect of PG(1)–PG(4) and AG(1)–AG(2) from external confounding factors identified by prior work:
  • Egocentricity [104, 85, 43]: people making decisions depending on personal gain, i.e., whether they (or people close to them) are personally affected by the disease that will benefit from data donation;
  • The data-collection entity [106, 48]: whether the collecting entity is a for-profit or a non-profit organization;
  • Socio-demographics [47, 34, 106]: age, gender, education, donation history, and technical background.
We find that even when told nothing about PETs implemented by the entity, participants are 23% more likely, on average, to expect a non-profit to implement PG(1)–PG(4) and AG(1)–AG(2). As a result of these already high privacy expectations for non-profit organizations, we find that mentioning a specific privacy protection in the survey does not significantly enhance people’s willingness to donate towards non-profit entities: even when no privacy protection is explicitly mentioned, 89% of the participants are willing to donate to a non-profit entity. In contrast, for-profit entities need to effectively prove their privacy protections; indeed, explicitly mentioning privacy protections in the survey does increase privacy expectations of for-profit entities from 50% to the level of non-profit entities. Privacy expectations, in turn, influence willingness to donate.
Furthermore, while the technical community has suggested external audits as a mechanism to enhance trust in PET implementation, our initial inquiry suggests that more work is needed to align expert and end-user perceptions of the effectiveness of such audits. In fact, the effect of audit statements on people’s willingness to donate is limited to a specific scenario involving for-profit entities and auditing to check that purpose restriction (PG(4)) is correctly implemented.
We argue that it is thus essential that non-profit entities correctly implement data privacy, as any future data leak would result in a severe loss of trust. In fact, prior work [4] highlights that users value maintaining expected privacy more than obtaining privacy they did not already expect. In contrast, we highlight the risk of for-profit entities using statements about PETs to engage in ‘privacy washing’ [26] to raise privacy expectations in order to collect data. Simultaneously, we note our respondent’s savviness to the gaps in general PETs guarantees to protect against e.g., data breaches and point to the need for future work to investigate how to effectively set expectations about stronger emerging PETs guarantees in the face of end-user skepticism.
2 Related Work
In this section, we examine prior work on people’s willingness to donate personal data, privacy concerns, the role of privacy-enhancing technologies (PETs), and educational or explanatory strategies related to PETs. Our review primarily focuses on literature specific to personal health data.
[…]
Data donation.Data donationData donation. The analysis of personal health data is crucial for medical research, particularly in diagnosing emerging or rare diseases and developing new treatments. Indeed, the COVID-19 pandemic has further intensified the demand for personal health data [94]. Despite the importance of this data, its sensitive nature poses significant obstacles to its collection [92, 21, 22, 71, 17, 35]. Prior work indicates that individuals are generally willing to donate their data for altruistic purposes [98, 5, 84, 5, 95], though there is a notable reluctance when it comes to their health data specifically [70, 72]. However, willingness to donate personal health data increases if the donor or their close family members are directly affected by the disease under investigation [41, 14].
[…]
Privacy concerns.Privacy concernsPrivacy concerns. Concerns about the privacy and the misuse of donated data are prominent among potential donors [72]. These concerns include the risk of being identified, discriminated against, or having their personal sensitive data misused or leaked [116, 5, 61, 68]. A common source of these concerns is distrust of the receiving entity [116, 91, 50, 5, 100], often due to fears that the entity might violate consent terms by sharing data with unauthorized third parties [54, 70]. This distrust is intensified when participants are unfamiliar with the recipient entity [116, 91, 50], especially if it is a governmental or for-profit organization [5, 100, 37, 48, 60]. Beyond this intentional data misuse, there are concerns about data leakage caused by hackers or unintentional mishandling [55, 114]. People are also reluctant to interact with entities with any history of data breaches, doubting their ability to safeguard sensitive information [108].
[…]
Privacy guarantees (PGs) and auditing.Privacy guarantees (PGs) and auditingPrivacy guarantees (PGs) and auditing. People typically expect the recipient entity to offer some level of privacy which depends on the type of data [55]. Consequently, security experts propose that recipient entities use PETs that implement one or more PGs and an auditing process. We focus on a specific set of 4 PGs: data expiration, data anonymization, use restriction, access control, and auditing (see section 1). In particular, we examine whether offering a given PG influences people’s willingness to donate health data to the recipient entity. Furthermore, we focus on two different auditing processes: expert auditing and self auditing.
To the best of our knowledge, no prior work surveys the effect of auditing (AG(1)–AG(2)) on people’s willingness to donate personal data, and only limited prior work investigates the effect of PG(1)–PG(4). For instance, Leon et al. [62] explored how data access, retention, and usage scope influence willingness to allow behavioral data collection, finding that retention and scope of use impact willingness to share personal information but that people were unwilling to pay more for sites that offered the preferred protections. Valdez and Ziefle [109] used conjoint analysis to study German participants’ willingness to donate health data to different recipients, with different quantified privacy risk, and differing purposes of data collection. They ran two conjoint analysis experiments, one expressing quantified risk of re-identification (for k-anonymity) and the other expressing privacy risk as a combination of sample size and exceptionalism (appropriate for differential privacy). They find similar results across both PETs: that the probability of identification and data recipient entity are the two factors that most influence data donation. Weiss et al. [115] examined participants’ willingness to share sensitive data based on varying epsilon thresholds of differential privacy, finding that the willingness to share is influenced by the purpose of the donation and amount of data subjects held by the entity. Other studies have examined public understanding of data retention and deletion [73], general data anonymization [45], and willingness to donate medical data if it is allowed to be used for secondary purposes [53].
We build upon prior work to present a comprehensive overview of current public attitudes towards mainstream privacy guarantees in health data donation. Specifically, we investigate whether PG(1)–PG(4) and AG(1)–AG(2) influence users’ willingness to donate health data.
[…]
Mental models on PETs.Mental models on PETsMental models on PETs. PETs aim to address user privacy concerns. While various PETs are available for health data donation environments [52], prior work indicates that only participants with online privacy literacy tend to have more trusting beliefs in PETs [46]. Studies focusing on sensitive personal data, including health data, reveal similar findings. For instance, Dechand et al. [33] found that most participants do not trust encryption because they do not understand how it works: e.g., many believe that service providers can bypass encryption to gain unauthorized data access. Furthermore, Lerner et al. [63] noted that some participants doubt the existence of true privacy.
Non-experts may misunderstand privacy technologies [118, 78, 33], e.g., viewing encryption merely as restrictive access control [118] or confusing it with data encoding [3]. Additionally, non-expert end users may struggle to understand the consequences of inadequate privacy protections [103, 111, 2, 40, 83, 110, 55]. As a result, even when privacy protections are present, people may have risk expectations that are misaligned with reality. We investigate the alignment between stated privacy guarantees and people’s expectations for how their data will be protected as part of RQ1.
[…]
Explaining PETs.Explaining PETsExplaining PETs. To address these misunderstandings, technologists and researchers have worked to explain PETs to the public to address their unfamiliarity with privacy concepts [29]. Prior work includes efforts to explain secure communication, such as end-to-end encrypted messaging, and differential privacy [2, 6, 10, 30, 31, 6, 33, 3, 118, 120, 75]. These efforts remain ongoing, as effectively and scalable setting privacy expectations remains a challenge. Methods found effective to explain PETs include visualizations [101], mental models [99, 66], nutrition labels [59], metaphors [56], short statements [80], and privacy games [107, 57]. In this work we use short statements to briefly explain the relevant guarantees in order to evaluate the impact of these guarantees on participants’ privacy expectations and willingness to donate medical data.
3 Methodology
As mentioned in section 1, we address research questions RQ1 and RQ2 through a user survey. This section details our survey’s design (subsection 3.1) and the cognitive interview process used to verify respondents’ comprehension of the presented scenarios (subsection 3.2). We also describe our participant selection strategy (subsection 3.3) and the limitations of our methodology (subsection 3.5).
3.1 Survey Structure
Our survey design is shown in Figure 1.
Figure 1: Survey design. Each participant is presented with a donation scenario with a survey introduction with two variations: recipient entity and disease. statements (PG(1)–PG(4) and AG(1)–AG(2)) are also presented in the donation scenario. Participants are then asked five sets of questions (survey questions). When the control statements are presented, the trust (dashed) is omitted. The willingness to donate and trust questions (dark shaded) are associated with an open-text question. The scenario understanding is used as a filter. Participants who did not understand the donation scenario were excluded from the analysis.
First, we present to the respondents a suvery introduction, which we detail in Figure 2.
Survey introduction
Imagine that an entity wants to develop a new treatment for disease. They need medical data from people with and without disease to develop the treatment. They ask you to donate your medical record to help develop the treatment. Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
Figure 2: Survey Introduction. The entity type (for-profit or non-profit) and the disease (selected from a list of common chronic diseases) are randomly selected.
We present a donation scenario using this introduction: we want to assess respondents’ willingness to donate their health data to a recipient entity developing a new treatment for a specific disease. The type of entity (for-profit or non-profit) and the disease (cancer, diabetes, heart failure, high blood pressure, and stroke) are randomly selected; the disease options are taken from a list of common chronic diseases published by the Centers for Medicare and Medicaid Services (CMS) [36].
Then, we present to the respondents either a control (no statement) or an experimental statement. Experimental statements are composed of either one privacy statement, alone, or one privacy statement and one auditing statement. Privacy statements are uniformly selected at random from a pool of four privacy statements; auditing statements are selected at random from two auditing statements. We have written these statements to align with PG(1)–PG(4) (see section 1), and we list them in Figure 3. We further distinguish the auditing guarantee into two different auditing statements to align with AG(1)–AG(2), and we list them in Figure 4.
PG(1): Anonymization — personally identifiable data will not be collected.
The privacy-preserving technology removes any personal identifiable information at the time of data collection, so that the data stored by the recipient entity is not linkable to its data owner (i.e., the data is anonymous).
PG(2): Access control — only relevant scientists can read data.
The privacy-preserving technology restricts data access to the authorized scientists (within the recipient entity) which are working on the treatment for the given disease.
PG(3): Data expiration — data will be deleted later.
The privacy-preserving technology discards the donated data, or makes it inaccessible after a given expiration time, which can be chosen by the data donor in the data collection agreement.
PG(4): Purpose restriction — data will not be used for other purposes.
The privacy-preserving technology ensures that the recipient entity can only use the donated data to develop a treatment for the given disease, and not in the context of any other research they may be working on (e.g., different disease).
Baseline: Control
No privacy-preserving guarantee is at all mentioned.
Figure 3: Privacy statements PG(1)–PG(4). One of the five statements (including the control statement) is randomly presented in the donation scenario immediately after the survey introduction.
AG(1): Expert auditing — the implementation of PET will be verified by experts.
An external advisory board of scientists and software engineers appointed by the recipient entity will regularly verify that the privacy-preserving technology is working as described. The results of this verification will be made public.
AG(2): Self-auditing — the implementation of PET can be verified by anyone.
Anyone interested, including the respondent and the experts the respondent trusts, will be able to verify that the privacy-preserving technology is working as described. Anyone, including the respondent and the experts the respondent trusts, can make their verification results public.
Baseline: Control
No auditing is at all mentioned.
Figure 4: Auditing statements AG(1)–AG(2). One of the three statements (including the control statement) is randomly presented in the donation scenario after the privacy statement.
Our survey questions assess respondents’ self-reported:
  • scenario understanding, used to filter out respondents that report not understanding the scenario (see Section 3.3): “How would you rate your understanding of the above scenario?” (4 point Likert scale: Fully understand - Not understand).
  • willingness to donate their health data to the recipient entity: “In this scenario how likely would you be to donate your medical record?” (4 point Likert scale: Very likely - Very unlikely).
  • privacy expectations regarding the specific privacy guarantees we investigate, measured by their agreement with the statements presented in Figure 5 using the same Likert scale as willingness to donate. [^1] This question included an additional attention check with a sixth statement: “If I donate my data, I will meet Albert Einstein.” Respondents who did not answer ‘Very Unlikely’ were removed from the experiment, as detailed in Section 3.3. We assess participants’ expectations about all guarantees, regardless of what statement they were presented.
  • trust that the recipient entity will implement the protection described by the privacy statement: “I trust the entity will handle my data as described.” (4-point Likert scale Strongly agree - Strongly disagree).
  • demographics and experiences, including age, gender, education level, donation history, and technical background.
Anonymization
My full name or other personal identifiable information will be linked to the donated medical record.
Access control
Any employee at the recipient entity will be able to access the donated medical records.
Data expiration
The donated medical record will be deleted at a set point in time.
Purpose restriction
The donated medical records will be used for another purpose without my consent
Expert auditing
A group of independent experts will verify whether the privacy-preserving technology works and publish a report on their findings.
Self-auditing
I will be able to hire someone to verify that my medical record is protected as described.
Figure 5: To measure privacy expectations respondents reported their agreement with each statement listed above on a 4-point Likert Scale from “Very Likely” to “Very Unlikely”.
We also ask respondents to explain their responses to the willingness to donate and trust questions.These explanations are collected through open-ended questions: “Why your are willing (or unwilling) to share your medical record with this entity?” and “Please explain why you do (or do not) trust that the entity will handle your data as described.” (We note that the scenario understanding and trust questions are only asked when a non-control privacy statement is presented).
3.2 Questionnaire Refinement
The final questionnaire was refined through multiple iterations. Before launching the survey, we prescreened 49 users on the Prolific online platform [1] by asking them to complete a short demographics questionnaire and indicate their willingness to participate in a cognitive interview. Then, we conducted 17 cognitive interviews to ensure participants understood the questionnaire correctly. Each interview involved presenting the questionnaire to participants and actively seeking feedback on their interpretation of the data donation scenario, and the privacy and auditing statements. We additionally presented all statements to the interviewees and asked for their understanding and perception of each of them. We continued refining the questionnaire through cognitive interviews until no further constructive feedback was received. In the final interview, we observe a clear understanding of the statements and the donation scenario. For example, the participant described anonymization (PG(1)) as "eliminating any sort of demographic or personal information affiliated with your data that could be associated with you if the data were to somehow be leaked" and access control (PG(2)) as "some sort of database that only certain individuals have access to by using a passcode and certain credentials to to log in." Each Prolific user who completed a cognitive interview was paid \$15.
3.3 Survey Sampling
Following best practice in security research [102], we recruited 560 respondents via Prolific. We restricted the respondents to adults residing in the U.S. to obtain focused results. Then, we ensured a balanced gender distribution: 49.11% men, 48.57% women, 1.6% non-binary, and 0.7% unwilling to disclose. Each participant received $0.83\$$ upon completing the questionnaire; the median completion time was around 7 minutes.
Out of these 560 respondents, we sequentially excluded the following from analysis: 13 that reported non-binary gender [^2] As we lack statistical power to draw meaningful conclusions about the privacy sentiments of the non-binary community specifically, we do not include these data in our analysis. However, this in no way diminishes the importance of understanding the unique and valuable insights of this population, which should be explored with either larger samples (yielding statistical power) or targeted research specific to their viewpoints and perspectives [39, 38].; 31 that submitted incomplete responses; 20 that failed our attention check (subsection 3.1); and 11 that indicated they did not understand the survey scenario (using scenario understand). Thus, our final dataset comprises 485 respondents, whose demographics are reported in Table I. How many respondents were assigned to each condition is reported in Table II.
Description Category $n$
Age $\geq$ 40 219 45.15%
< 40 166 54.85%
Gender Woman 254 52.37%
Man 231 47.63%
Education w/ BSc degree 352 72.58%
w/o BSc degree 133 27.42%
Technical Background Yes 128 26.39%
No 357 73.61%
Donation History Yes 55 11.34%
No 430 88.66%
TABLE I: Participant demographics. We note that the second category of each demographical attribute (e.g., no technical background) is considered the baseline scenario during analysis.
For-Profit Non-Profit
AG(1) AG(2) Ctrl. AG(1) AG(2) Ctrl.
PG(1) 22 18 19 19 19 18
PG(2) 17 18 19 18 18 19
PG(3) 18 20 18 18 17 17
PG(4) 21 18 18 20 18 19
Ctrl. - - 19 - - 20
TABLE II: Number of respondents assigned to each condition (i.e., who saw each privacy statement and auditing statement, or no privacy statement (bottom row) or no auditing statement (columns 4 and 6).
3.4 Analysis
We analyzed the open-text questions about willingness to donate and trust (see subsection 3.1) using inductive-thematic open coding. Two researchers independently coded each entry and generated a codebook from a random sample of at least 100 (20.6%) responses. Then, they composed a final codebook and double-coded all responses. Since all responses were double-coded and inconsistencies were resolved, we do not report inter-rater reliability (IRR) [69].
For all quantitative analyses, we binarized the responses to the willingness to donate, privacy expectations, and trust questions (see subsection 3.1)) to get ‘True’ if the respondents selected ‘Likely’ or ‘Very Likely’, and ‘False’ otherwise.
In addition to presenting a descriptive analysis detailing the distribution of responses on the survey items that address our research questions, we construct logistic regression models to analyze factors related to two dependent variables: privacy expectations (RQ1) and willingness to donate (RQ2). For RQ1, the independent variables were the presence of a given privacy and/or auditing statement. The dependent variable was the privacy expectation corresponding to the privacy statement presented to a given respondent. Using this model we compare the privacy expectations of respondents in the experimental conditions (those shown a privacy statement) with those of the control groups. Responses from respondents in an experimental conditions are only modeled in the analysis of the privacy statement they were shown.
For RQ2, the dependent variable was willingness to donate, and the independent variables were the presence of a given privacy statement, the privacy expectation for each guarantee, as well as demographics and experiences. We categorized education into two groups: with and without a bachelor’s degree. We also categorized age into two groups and cut to the age of forty. We took “no bachelor’s degree” and “less than forty years old” as the reference categories, respectively. For gender, we took man as the reference category. Technical background and donation history are binary factors, and we took the negative response as the reference category. We built separate regression models for the two recipient entities: for-profit and non-profit.
3.5 Limitations
Although the four privacy-preserving guarantees examined in this study (section 1) span many PETs, they are not fully comprehensive and do not capture the full complexity of PETs nor of real-world threats, which may involve privacy issues beyond the scope of our investigation. Moreover, despite our efforts to mitigate misunderstandings through rigorous cognitive interviews, filtering out respondents who reported not fully understanding the scenario, and controlling for privacy expectations in our statistical analyses, the brief descriptions of each guarantee might have led to partial or incorrect understanding among respondents, potentially influencing our results. While this study employs vignette scenarios that can effectively mirror real-world behaviors [90], discrepancy between participants’ expressed privacy concerns and their actual behaviors may exist (i.e., privacy paradox [13]).
Furthermore, statistically-significant results do not inherently guarantee the absence of underlying biases that could skew the findings. In fact, various unmeasured factors that were not controlled for could influence the interpretations of the results. Finally, our respondents were recruited via Prolific, and thus, may not accurately reflect the diversity of the entire U.S. population, let alone the varied perspectives of individuals from other countries. This limitation challenges the generalizability of our findings.
4 Results
4.1 RQ1: Understanding of PGs
Next, we answer the first research question (RQ1, see section 1): “How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?". We analyze and report the survey responses and evaluate the rationales for the responses.
[…]
Analysis Set Up.Analysis Set UpAnalysis Set Up.
We compare privacy expectations on PG(1)—PG(4) and AG(1)–AG(2) among groups of respondents who received different privacy statements (listed in Figure 3) and auditing statements (listed in Figure 4). In Table III, we show the privacy expectations of respondents who were shown different privacy statements.
To assess the statistical significance of the descriptive quantitative results presented above, we used binomial logistic regression to analyze the relationship between the presence of a privacy statement in the scenario and respondents’ privacy expectations. We summarize results, which we distinguish for non-profit and for-profit entities, in Table IV.
Relationship between Privacy Statement Shown and Privacy Expectation Statement Shown Overall For-Profit Non-Profit PG(1) Anonymization 0.68 0.68 0.67 + AG(1) Expert audit 0.63 0.59 0.68 + AG(2) Self audit 0.65 0.67 0.63 Control 0.38 0.26 0.50 PG(2) Access control 0.50 0.58 0.42 + AG(1) Expert audit 0.51 0.35 0.67 + AG(2) Self audit 0.50 0.56 0.44 Control 0.49 0.37 0.60 PG(3) Data expiration 0.83 0.67 1.00 + AG(1) Expert audit 0.92 0.83 1.00 + AG(2) Self audit 0.86 0.85 0.88 Control 0.41 0.37 0.45 PG(4) Purpose restriction 0.78 0.72 0.84 + AG(1) Expert audit 0.71 0.62 0.80 + AG(2) Self audit 0.81 0.67 0.94 Control 0.56 0.37 0.75 AG(1) Expert audit 0.82 0.78 0.87 Control 0.62 0.58 0.66 AG(2) Self audit 0.53 0.50 0.57 Control 0.26 0.25 0.28
TABLE III: Percentage of respondents who had a positive privacy expectation when shown a particular privacy statement in their donation scenario. The values in the table are the percentage of respondents in a given condition who had the privacy expectation (see Figure 5) that corresponded to the privacy statement they were presented. The 2nd column (“overall") reports results across both entities (for-profit and non-profit), the 3rd column reports the results of the subset in the for-profit entity conditions, and the 4th column reports the results of the subset in the non-profit entity conditions. (Recall that agreement has been binarized as discussed in subsection 3.4). For example, the left-most and top-most numerical cell indicates that 68% of participants in both entity scenarios who were shown PG(1) alone – with no auditing statement – expected that their data would be anonymized.
Statement Shown OR 95% CI $p$-value
For-Profit PG(1) Anonymization 5.07 [1.69, 17.52] 0.006
PG(2) Access control 1.71 [0.6, 5.23] 0.325
PG(3) Data expiration 6.29 [2.09, 20.44] 0.001
PG(4) Purpose restriction 3.43 [1.19, 10.59] 0.026
AG(1) Expert audit 2.59 [1.33, 5.20] 0.006
AG(2) Self audit 3.04 [1.59, 5.93] <0.001
Non-Profit PG(1) Anonymization 1.95 [0.69, 5.57] 0.208
PG(2) Access control 1.27 [0.45, 3.61] 0.651
PG(3) Data expiration 16.7 [3.64, 120.49] <0.001
PG(4) Purpose restriction 2.04 [0.55, 7.11] 0.266
AG(1) Expert audit 3.41 [1.59, 7.85] 0.002
AG(2) Self audit 3.41 [1.79, 6.61] <0.001
TABLE IV: Effect of presenting different privacy statements on respondents’ privacy expectations toward the two recipient entities. OR: odd ratio; CI: confidence interval.
[…]
Response Distribution and Statistical Analysis Results.Response Distribution and Statistical Analysis ResultsResponse Distribution and Statistical Analysis Results.
We find that privacy expectations differ based on the data-collecting entity. In the control group, which was not shown any privacy statement, 26%-37% of respondents expected for-profit entities to provide anonymization, data expiration, access control, use restriction to a specific purpose, expert auditing, and self-auditing. In contrast, a higher percentage of respondents (45%-75%) expected non-profit entities to employ these mechanisms. For instance, we note that respondent P430 (who received a for-profit scenario) expected privacy protections even though they were in the control group (and thus received no statement about privacy protection): "I also assume that the data is looked at in the aggregate and likely no one at the company knows me.". In the same situation, respondent P53 (who received a non-profit scenario instead) made a stronger assumption: "I trust that the organization will uphold strict privacy and ethical standards".
In the for-profit scenario, all privacy statements, except for the access-control statement, significantly increase the corresponding privacy expectations ($p$-value < $0.05$). In the non-profit scenario, where expectations are already high, only the privacy statement about data expiration significantly increases expectations (from 60% to 100%). Comparing the for-profit and non-profit subsets of the descriptive results (2nd and 3rd column of Table III), we observe that presenting a privacy statement was particularly beneficial in for-profit scenarios. Indeed, the presence of a privacy statement raised respondents’ privacy expectations to nearly the same level as non-profit scenarios for anonymization (68% and 67%) and access control (58% and 42%). It also reduced the gap in expectations for data expiration (67% and 100%)) and purpose restriction (72% and 84%)).
In both scenarios, the auditing statements increase the expectation that a given privacy guarantee will be audited ($p$-value < $0.05$). However, we note that incorporating an audit statement does not markedly change privacy expectations. The statistical significance of the auditing statements in Table IV indicates that respondents shown the audit statement correctly expected the corresponding type of audit, while respondents that were not shown the auditing statement did not expect this protection. As observed in  Table III, presenting an auditing statement does not increase the number of respondents who expect the privacy guarantee implied by the privacy statement they were shown, despite the stronger assurances that audits imply.
[…]
Qualitative Analysis of Privacy Expectations via Open-Answer Responses.Qualitative Analysis of Privacy Expectations via Open-Answer ResponsesQualitative Analysis of Privacy Expectations via Open-Answer Responses.
Across for-profit and non-profit scenarios, 69% and 85% of respondents, on average, expressed that their positive privacy expectations were formed based on the presence of the privacy statement, belief in privacy obligations of the entity, trust in the entity, or trust in the auditing process.
Satisfaction with the privacy guarantees. 16.3% and 20.8% of respondents in the for-profit and non-profit scenarios, respectively, were convinced by the received privacy statement. Some respondents, like P56, vaguely stated that the policy looks promising: "I see their policy, and they have to follow their own policy.". Others felt explicitly safer and trusted the statement. For example, P494 noted: "Trust is very important when it comes to medical data. I believe the organization has privacy policies that outline that they collect and will use my data. I also believe the organization will employ security measures to safeguard data". Similarly, P555 highlighted the presence of the privacy statement: "(…) purposely states there is software in place to conserve privacy." and P516 felt their information was safe because of it: "(…) with the privacy protection in place, they are isolating the data they need while basically "throwing out" the rest by putting in under that protection. In essence, my information is safe, and they’re only using what they said they’d use". P208 also felt safer contributing to research knowing the privacy mechanisms were in place: "I want to be able to contribute to research to better improve cancer treatment, and I feel safe if my data is protected through the mechanisms above".
Belief in legal and reputational obligations. 16.3% and 10.6% of respondents in the for-profit and non-profit scenarios, respectively, believed that entities are forced to protect the donated data due to legislative requirements and reputational concerns. For example, P28 stated regarding for-profits: "A for-profit organization wouldn’t want to violate HIPAA, HITECH laws". Additionally, P117 wrote "I would trust them to do the right thing so they won’t face lawsuits". Regarding non-profits, P141 noted: "I trust that the law will restrict any data leaks to third parties".
Trust in the recipient entity. 13.9% and 17.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed a general trust in the recipient entity without specifying reasons. For instance, P70 succinctly said "I feel they are reliable and trustworthy" and P50 stated that "I assume they take their research seriously, so they would handle the data carefully". Other respondents, like P145, reported having no reason not to trust it: "I have no reason to think they would do anything nefarious with my medical data".
Trust in the auditing process. 7.8% and 7.9% of respondents in the for-profit and non-profit scenarios, respectively, found the auditing statement to add at least some reliability. For example, P202 showed some reservation but found confidence in the public nature of audits: "I don’t fully trust them, but I somewhat do, particularly if audits and verification of results are made public. That said, claiming that the advisory board is external is only partly reassuring, as it’s appointed by the institute." Additionally, P540 had trust in the entity’s data handling because "they let outsiders audit them" and P51 (who received the expert-auditing statement) because of "safelocks and checks in place". On the other hand, one interviewee doubted the expert auditing process and stated: "somebody else says something doesn’t mean that it’s real.". No questionnaire respondent explicitly raised concern towards the auditing process.
Across for-profit and non-profit scenarios, 31% and 15% of respondents, on average, expressed negative privacy expectations because they were skeptical of the privacy statement or doubted whether the recipient entity would actually employ PETs as stated. Their qualitative responses offer insights on the underlying reasons:
Skepticism on the privacy statement and limits of privacy-preserving guarantees. 21.2% and 10.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed general distrust in the feasibility of the privacy statement. In the non-profit scenario, P34 stated that "no privacy technology is foolproof". Similarly, in the for-profit scenario, P45 wrote: "I don’t trust that the privacy-preserving tech would work." More precisely, privacy statements were considered too ideal to be fully enforced. Some respondents believed that unauthorized employee access would be unavoidable. For instance, P71 said: "Why would I trust someone other than my doctor with my medical records? These days especially, I don’t trust anyone. There could be a breach or simply people I don’t know from a hole in the wall will then have access to all my med records. Insane!". Others felt that data breaches and cyberattacks were inevitable. For instance, P83 stated that "the primary reason for my distrust is due to past news of companies being hacked by people and their data getting leaked", while P124 noted that "corporate data breaches are very common", and P80 echoed this sentiment, saying: "I believe the [intentions] will be good, but data can be hacked". We observe that the auditing statement did not substantially instill trust on all accounts — with the exception of expert auditing for access control in non-profit scenario, and data expiration in for-profit scenario. For example, P474 mentioned: "They can check their privacy technology all they want but when there is a breach it is done and info is stolen. After it fails then they say sorry and offer monitoring but the info is still stolen".
Doubt on the recipient entity’s motivation to employ PETs. 8.6% and 3.3% of the for- and non-profit scenarios, respectively, claimed that these entities are inherently self-serving and lack the motivation to uphold privacy-preserving guarantees or implement such measures at all. Most criticism was directed at for-profit entities. For instance, P55 noted that for-profit entities "will do what is profitable and not much more than that". Similarly, P119 stated "(…) because it is a for-profit organization. I expect them to cut corners" and P32 wrote that "for-profit organization have low standard of morality". However, some respondents also expressed concerns about non-profit entities. For example, P190 in the non-profit scenario remarked: "Medicine has become a business. My data is only useful to them if it helps them make more money. Money comes first before the actual well being of humans".
4.2 RQ2: Willingness to donate health data
Next, we answer the second research question (RQ2, see section 1): “How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?". As in the previous subsection, we statistically analyze the responses and then report the qualitative rationales collected from the open-text questions.
Statement Shown Overall For-Profit Non-Profit
PG(1) - Anonymization 0.68 0.63 0.73
+ AG(1) - Expert audit 0.63 0.59 0.68
+ AG(2) - Self audit 0.70 0.67 0.74
PG(2) - Access control 0.68 0.67 0.68
+ AG(1) - Expert audit 0.69 0.59 0.78
+ AG(2) - Self audit 0.58 0.61 0.56
PG(3) - Data expiration 0.73 0.67 0.80
+ AG(1) - Expert audit 0.69 0.67 0.72
+ AG(2) - Self audit 0.68 0.70 0.65
PG(4) - Purpose restriction 0.72 0.53 0.89
+ AG(1) - Expert audit 0.71 0.71 0.70
+ AG(2) - Self audit 0.72 0.72 0.72
Control 0.65 0.60 0.68
TABLE V: Percentage of respondents willing to donate their personal health data to the recipient entity in each scenario.(e.g., the left-most and top-most numerical cell indicates that 68% of participants in both entity scenarios that with PG(1) presented but no auditing statement were willing to donate.)
[…]
Analysis Set Up.Analysis Set UpAnalysis Set Up.
Table V summarizes the respondents’ willingness to donate their health data to a recipient entity. The 1st column (“overall") reports results for all entity types, the 2nd column reports the results of the subset with for-profit entity, and the 3rd column reports the results of the subset with non-profit entity. The values represent the percentage of respondents willing to donate their heath data to the recipient entity.
We constructed a logistic regression model to understand the factors that influence willingness to donate. We first analyze and confirm the influence of the recipient entity and then separate the participants based on the recipient entities and analyze them in separate models. The results are summarized in Table VI.
In Figure 6, we visualize the overlap in experimental group respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described.
Entity Factor Levels OR 95% CI $p$-value
Both Entity - FP 0.66 [0.45, 0.98] 0.038
For-Profit Privacy Statement
   PG(1) 1.14 [0.39, 3.24] 0.809
   PG(2) 1.24 [0.42, 3.58] 0.696
   PG(3) 1.54 [0.52, 4.47] 0.432
   PG(4) 1.35 [0.45, 3.88] 0.584
Privacy Expectation
   PG(1) 3.07 [1.21, 8.06] 0.020
   PG(2) 2.64 [1.00, 7.33] 0.054
   PG(3) 6.06 [2.15, 18.27] <0.001
   PG(4) 6.33 [2.34, 18.49] <0.001
   AG(1) 5.34 [2.71, 10.82] <0.001
   AG(2) 2.94 [1.46, 6.24] 0.003
Demographics & Experiences
   Education 0.75 [0.33, 1.63] 0.466
   Age 0.68 [0.33, 1.39] 0.290
   Gender 0.80 [0.40, 1.61] 0.535
   Tech Background 1.33 [0.57, 3.15] 0.512
   Egocentrism 0.41 [0.20, 0.84] 0.016
   Donation History 9.47 [2.22, 54.68] 0.005
Non-Profit Privacy Statement
   PG(1) 1.17 [0.36, 3.52] 0.783
   PG(2) 0.88 [0.27, 2.60] 0.823
   PG(3) 1.16 [0.35, 3.55] 0.794
   PG(4) 1.45 [0.44, 4.45] 0.522
Privacy Expectation
   PG(1) 7.32 [2.48, 24.13] <0.001
   PG(2) 2.03 [0.76, 5.64] 0.163
   PG(3) 3.13 [0.78, 12.76] 0.102
   PG(4) 12.15 [3.31, 52.60] <0.001
   AG(1) 3.19 [1.49, 6.82] 0.003
   AG(2) 3.74 [1.72, 8.89] 0.001
Demographics & Experiences
   Education 0.37 [0.15, 0.86] 0.025
   Age 1.28 [0.61, 2.70] 0.516
   Gender 0.83 [0.38, 1.78] 0.630
   Tech Background 0.39 [0.16, 0.89] 0.026
   Egocentrism 1.79 [0.85, 3.96] 0.136
   Donation History 4.11 [1.23, 17.58] 0.034
TABLE VI: Influences of privacy expectations and demographics on willingness to donate of the three participant groups. OR: odd ratio. CI: confidence interval.
Figure 6: Overlap in respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described (respondents in the control – no privacy statement – are excluded). The numbers outside each circle summarize the total number of respondents who were e.g., willing to donate their data (305). $N=456$.
[…]
Response Distribution and Statistical Analysis Results.Response Distribution and Statistical Analysis ResultsResponse Distribution and Statistical Analysis Results.
There was no direct relationship between showing a privacy statement and respondents’ willingness to donate. However, several privacy expectations positively correlate with willingness to donate for both non-profit and for-profit entities, except in specific cases like access control for both for- and non-profit and data expiration for the non-profit.Visualized in Figure 6, we observe high alignment among the three explored constructs, where 234 (51.3%) respondents gave all positive responses and 70 (15.4%) respondents gave all negative responses. Combined with the results of RQ(1), the findings suggest that willingness to donate is influenced by privacy expectations, which are, in turn, impacted by privacy statements.
Privacy statements and expectations are not the only reason for donation intentions however. While participants were generally less willing to donate to for-profit entities (OR=0.66, $p=0.038$), when analyzing only control group participants, we observe that the absolute difference in willingness to donate between for-profit and non-profit entities was 8.8% was smaller than the 20.3% difference in privacy expectations between the two entities. This indicates that other non-privacy-related factors, such as the perceived impact or benefits of the donations, also influence participants’ willingness to donate.
Furthermore, we observe that respondents with egocentric relationships to the donation subject are less willing to donate to a for-profit entity (OR = 0.41, $p=0.016$). This is potentially related to their existing (negative) experiences with for-profit entities in healthcare. For example, P136 claimed they have heard a number of times that the entity says they won’t sell your information but can get around it by calling it something else, thus technically selling the information. P244 also mentioned that their information has been leaked before by a clinician. An alternative explanation could be the perceived imbalance in cost and profit between the donor and the recipient entity, where the donor provides data for free while the recipient entity profits from its use. For instance, participants P495, P463, and P459 all share this perception. However, we note that some individual participants alternatively were motivated to donate due to egocentricity, even to for-profit entities. For example, P45 "just lost my father to heart failure thirteen days ago. (…) I would be more than willing to donate my medical records if it helps develop medicines to make hearts function better.
Respondents with bachelor’s degrees are less willing to donate to a non-profit entity (OR = 0.37, $p=0.025$), which may be related to their greater awareness of the complexities and potential vulnerabilities with privacy statement. They have heightened privacy concerns that need to be addressed. For example, P42 wanted to verify that the PET worked visually; P97 thought that only the stored data was encrypted, but the data transaction was not protected the same way; and P236 argued that: “privacy-preserving technology that works today may not work in a few years.”
Prior donors are more willing to donate to both entities. Respondents who have donated before are more willing to donate to both entities, with the OR value of 9.47 ($p=0.005$) and 4.11 ($p=0.034$) for the for and non-profit entities, respectively.These respondents expressed strong willingness of supporting science and recognized the importance of medical data in developing new treatments. For example, P40 claimed: “I feel that it is important information to share in hopes that they can find better ways to deal with diabetes." and P208 responded: “Breakthroughs in science and moving forward in knowledge, our greatly benefited by such Data Collection."
[…]
Qualitative Analysis of Donation Intention Via Open-Answer Responses.Qualitative Analysis of Donation Intention Via Open-Answer ResponsesQualitative Analysis of Donation Intention Via Open-Answer Responses.
Many respondents reasons for donating (or not) included a desire to support research, ego-centric connections, or desire for personal reward. In the for-profit scenario, 50.2% of respondents’ reasons for donating (or not) fell into these categories, in the non-profit scenario, 53.3% did.
One non-privacy-related reason for donating included supporting research: e.g., P107 expressed that "there are a number of things I could share to research such as money, time, data, and more. I would be willing to share my medical records to contribute to cancer research in hopes that people will be healthier and to enhance research future medicines". Similarly, P97 stated: "even if they are for-profit (which I don’t like), any closer we get to developing better treatments and/or a potential cure to cancer is something I’d be willing to assist in".
Ego-centric reasons, particularly familial connections, also motivated donations. P154 mentioned donating because "my dad died when his heart failed on him, and I’d have rather that not had happened", and P233 noted that they "would be willing to help to prevent someone else’s mom from dying of a stroke like [theirs] did".
Some respondents desired personal reward for donating and thus were not interested in donating in our scenario: e.g., P92 explained that "I would give them my data if I got paid for it. I would think twice if they want me to donate the data and make no money from it while the researchers will make money off the of research they compiled".
Respondents expressed negative sentiments toward for-profit entities collecting data for research. The desire for personal reward related closely to lack of donation intent for for-profit organizations. In the for-profit scenario, 11.4% explicitly refused due to negative perceptions of such organizations. For instance, P191 and P202 respectively stated: "I have a negative connotation with for profit organization" and "I feel like this type of organization already profits significantly off a large number of people (occasionally off of me as well) and as such, I do not want to give them direct permission to profit further off of me." Furthermore, respondents felt that for-profits already benefit enough and were unwilling to contribute further. For instance, P36 wrote: "For profit organization makes profit using the data. I’m never going to donate it. They can BUY it from me for a reasonable amount."
Alternatively, a small group of 1.3% respondents explicitly mentioned that they were willing to donate data to entities of non-profit nature, but even they expressed concerns about potential misuse or unethical behavior within these organizations. For instance, P122 mentions: "it’s always going to be in the back of my head that there’s a possibility that it’s being sold or taken or used in some way I’m not okay with. Scares me a bit". Then, P282 notes: "it is difficult to police everyone in a non-profit organization. I have seen leaders in organizations act unethically at times, which makes me think that even a well-intended organization cannot fully control the actions of every employee or volunteer they have."
Privacy-related donation considerations focused on data sensitivity and leakage. Concerns for not donating were mostly about data leakage, with respondents expressing that they 12.9% and 13.9% participants in the for-profit and non-profit scenario, respectively, were protective of their data and hence would not donate at all, as P545 noted that "there are so many data leaks from so called safe places. No one can anticipate what hackers can do in the future." Others cited previous data leak experiences as to why they would not donate, as P290 said "my medical data has been breached in the past by a clinician and I had to file a formal grievance against the health care system that employed her. For this reason, because my medical history was abused, I no longer have any trust and will never voluntarily consent to my medical history/information being shared with others." Some respondents who would not donate data also expressed that they would only possibly donate to specific entities that have already proven trustworthiness to them, as P357 explained that "regardless of privacy policy, the probability of a leak or misuse is high. unless it is an organization I have had personal interaction with, or am very familiar with, it is unlikely I would donate my medical records.". Another reason was that the participant was concerned that their medical data included too much or too specific information. P155 highlighted this concern: "I would not feel comfortable with so much of my personal health information being shared with an organization that is not involved with my direct medical care, regardless of whether there is an advisory board or not"; as did P229: "It’s a lot of detailed information that I’m worried if it somehow gets in the wrong hands, it could reveal a lot of private information about me." Contrastingly, some respondents felt comfortable sharing data due to a lack of sensitive information (e.g., "I don’t have mayor illnesses or nothing to hide so Im ok with that") or because they already share information with other entities (e.g., "Considering I already disclose this information to other organizations (for example, data collection on phone, data compiled through search history), other agencies likely have the information on hand already and so another organization having it is no different").
5 Discussion
Analyzing the effect of privacy statements on users’ perceptions, trust, and willingness to donate in the context of a survey is particularly useful because the survey mimics the actual process of consenting that recipient entities use when requesting data. These entities typically provide an introduction, present privacy guarantees, and ask for consent to donate data, much like our survey.Our study reveals several key findings about privacy expectations, their influence on willingness to donate health data, and the implications for both researchers and regulators.
[…]
Pre-existing Privacy Expectations.Pre-existing Privacy ExpectationsPre-existing Privacy Expectations. We find that people have strong pre-existing privacy expectations even when no specific privacy statements are shown. This expectation is particularly high for non-profit entities, which are perceived as more ethical, thus prioritizing public good over profit. In contrast, for-profit entities are seen as driven by financial gain, leading to skepticism on their commitment to protecting user privacy. Non-profit entities foster greater trust and confidence in their data handling practices to the point that respondents often expect more stringent privacy measures, such as purpose use restriction. However, such measures may not be technically enforced or commonly practiced; indeed, many health organizations employ broad (not purpose restricted) consent [58]. This gap between expectations and reality can lead to privacy expectation violations, which are a significant source of privacy concerns according to the theory of contextual integrity [77].
[…]
Impact of Privacy Statements.Impact of Privacy StatementsImpact of Privacy Statements. Privacy statements play a role similar to advertising, such as those used by companies like Apple [79]. For for-profit entities, where baseline privacy expectations are lower, privacy statements raise these expectations. (In contrast, for non-profits, the impact is minimal because the expectations are already high.) Thus, we argue that privacy expectations of for-profits can be effectively influenced by transparently implementing and communicating about PETs.
[…]
Privacy Expectations and Willingness to Donate.Privacy Expectations and Willingness to DonatePrivacy Expectations and Willingness to Donate. Privacy expectations significantly influence people’s willingness to donate health data: individuals are more willing to donate when their privacy expectations are met or exceeded, as presented in Figure 6. Thus, we argue that entities wishing to receive people’s data should not only focus on employing PETs but also on effectively communicating these measures to potential data donors. However, there is a risk of "privacy washing" [67, 26], where information about PETs and “vague claims” are used to raise expectations and mislead individuals into donating their personal data.
[…]
The Role of Auditing.The Role of AuditingThe Role of Auditing. Auditing had a limited effect on raising privacy expectations, except in the case of purpose use restrictions in the for-profit scenario. Participants recognize that PETs are not flawless and that auditing does not practically mitigate the risk of data breaches. This result highlights a disconnect between the expert view, which often sees auditing as a fundamental guarantee to offer, and the participant view, which is more concerned with the overall protection of data rather than specific PETs implementations (e.g., a respondent writes “they can guarantee privacy all they want, things still get hacked”).
[…]
Recommendations for Researchers.Recommendations for ResearchersRecommendations for Researchers. If researchers aim to foster trust in entities that collect personal health data, they may want to focus on effectively communicating about PETs that protect against data breaches. Technologies like multi-party computation, although not yet scalable, could enhance trust in data protection: they specifically distribute sensitive information across different entities to ensure that no single entity can reconstruct the information on its own, thereby minimizing the impact of a breach. However, deploying such systems requires careful engineering to prevent distributed data breaches on all the entities involved. We believe that convincing the public of the effectiveness of these measures presents an even greater challenge, as there is existing doubt and any failure of these systems could severely damage trust.
[…]
Recommendations for Regulators.Recommendations for RegulatorsRecommendations for Regulators. Our findings suggest the necessity of policies that ensure privacy practices align with public expectations. Given the strong assumptions people have about privacy protections, especially for non-profit entities, regulations should be designed to ensure that the protections assumed are consistently applied.

### v2

@@
-The Role of Privacy Guarantees in Voluntary Donation of Private Data for Altruistic Goals
-Abstract
-Voluntary donation of private information for altruistic purposes, such as advancing research, is common.
-However, concerns about data misuse and leakage may deter individuals from donating their information.
-While prior research has indicated that Privacy Enhancement Technologies (PETs) can alleviate these concerns, the extent to which these techniques influence willingness to donate data remains unclear.
-This study conducts a vignette survey ($N=485$) to examine people’s willingness to donate medical data for developing new treatments under four privacy guarantees: data expiration, anonymization, use restriction, and access control.
-The study explores two mechanisms for verifying these guarantees: self-auditing and expert auditing, and evaluates the impact on two types of data recipient entities: for-profit and non-profit institutions.
-Our findings reveal that the type of entity collecting data strongly influences respondents’ privacy expectations, which in part influence their willingness to donate data.
-Respondents have such high expectations of the privacy provided by non-profit entities that explicitly stating the privacy protections provided makes little adjustment to those expectations.
-In contrast, statements about privacy bring respondents’ expectations of the privacy provided by for-profit entities nearly in-line with non-profit expectations.
-We highlight the risks of these respective results as well as the need for future research to better align technical community and end-user perceptions about the effectiveness of auditing PETs and to effectively set expectations about the efficacy of PETs in the face of end-user concerns about data breaches.
-1 Introduction
-The altruistic analysis of personal health data can bring tremendous societal benefit.
-For instance, the public health sector could leverage individuals’ medical history and health data to analyze epidemiological patterns, support disease surveillance and risk prediction, diagnose rare and emerging diseases, and improve the development of new treatments [18, 5, 14, 42, 61, 106].
-The potential of these applications entails economic benefits: for example, the EU Commission believes that innovations supporting data sharing and data availability may save billions of Euros in the EU health sector [28].
-Consequently, the need for personal data in the health sector is ever-increasing [117, 18] — especially with the recent boom of large language models (LLMs), whose capabilities depend on the data they are fed.
-Personal health data is already being collected by care providers [14, 61] and personal mobile apps and devices [42]; then, third parties may get access to this data through data sharing agreements that regulate data donation [106].
-However, people are often unwilling to donate such sensitive personal data due to privacy concerns [88, 112, 97, 42, 61].
-For example, people may be uncomfortable with hospital policies that allow data sharing agreements with third parties without the data owners’ explicit consent [64].
-In general, privacy concerns significantly influence (i) people’s adoption of, engagement with, and ability to benefit from technology [15, 89] as well as (ii) the regulations suggested by a country’s legislators [27].
-In an effort to foster data collection and analysis while protecting the privacy of data contributors, there has been an increasing focus on deploying privacy-enhancing technologies (PETs) for data storage, data processing, and machine learning (ML) training.
-For example, encrypted databases such as CryptDB [87] or BlindSeer [81] store user data in encrypted form; hardware-secured platforms [8, 12, 16] additionally rely on Trusted Execution Environments (TEEs) to process data in a secure hardware compartment, so that the confidentiality and integrity of this data (and the code processing it) is guaranteed even during computation; secure Multi-Party Computation (MPC) platforms [44, 25] allow distinct entities to pool their encrypted data and jointly analyze it, without any single entity learning other entities’ data, even indirectly; finally, systems combining all these approaches [119, 32] expand the set of guarantees achievable by any single approach.
-In addition, many PETs implement differential privacy (DP) [7] to prevent the results of statistical computations on a given sensitive dataset from leaking the presence (or absence) of a specific user’s data.
-Despite growing research and deployment of PETs, people feel less control over their data than ever [9].
-Arguably, the effort towards strengthening and widening the spectrum of PETs’ privacy-preserving guarantees is insufficient if people’s privacy concerns are not mitigated in practice.
-Prior work [30, 2, 6, 10, 31, 33, 3, 118, 120, 75] has sought to meaningfully explain such guarantees—or the technologies that implement them—to the people they aim to protect.
-We build on this work to evaluate whether such guarantees effectively alleviate the people’s privacy concerns, so that they would be more willing to donate their personal data.
-While much of the prior work that shares our aim has studied specific, singular technologies or guarantees (often DP [30, 19, 75] or end-to-end encryption (E2EE) [33, 11, 2]), we focus on the more general security and privacy guarantees recurring in PETs:
+The Role of Privacy Guarantees in Voluntary Donation of Private Health Data for Altruistic Goals
+Abstract.
+The voluntary donation of private health information for altruistic purposes, such as supporting research advancements, is a common practice.
+However, concerns about data misuse and leakage may deter people from donating their information.
+Privacy Enhancement Technologies (PETs) aim to alleviate these concerns and in turn allow for safe and private data sharing.
+This study conducts a vignette survey ($N=494$) with participants recruited from Prolific to examine the willingness of US-based people to donate medical data for developing new treatments under four general guarantees offered across PETs: data expiration, anonymization, purpose restriction, and access control.
+The study explores two mechanisms for verifying these guarantees: self-auditing and expert auditing, and controls for the impact of confounds including demographics and two types of data collectors: for-profit and non-profit institutions.
+Our findings reveal that respondents hold such high expectations of privacy from non-profit entities a priori that explicitly outlining privacy protections has little impact on their overall perceptions.
+In contrast, offering privacy guarantees elevates respondents’ expectations of privacy for for-profit entities, bringing them nearly in line with those for non-profit organizations.
+Further, while the technical community has suggested audits as a mechanism to increase trust in PET guarantees, we observe limited effect from transparency about such audits.
+We emphasize the risks associated with these findings and underscore the critical need for future interdisciplinary research efforts to bridge the gap between the technical community’s and end-users’ perceptions regarding the effectiveness of auditing PETs.
+Privacy Guarantee, Health Data Donation, Privacy Expectation, Auditing
+ccs: Security and privacy Social aspects of security and privacy
+1. Introduction
+The altruistic use of personal health data holds immense potential for societal benefit.
+For example, the public health sector can leverage individuals’ medical histories and health data to analyze epidemiological trends, enhance disease surveillance and risk prediction, diagnose rare and emerging diseases, and accelerate the development of new treatments (Bietz et al., 2015; Aitken et al., 2016; Bartlett et al., 2018; Greene et al., 2019; Kim et al., 2015; Trinidad et al., 2020).
+Beyond these applications, there are significant economic advantages.
+The European Commission estimates that innovations enabling greater data sharing and availability could save billions of euros within the EU health sector (Commission, [n.
+d.]).
+As a result, the demand for personal health data continues to grow (Wilbanks and Friend, 2016; Bietz et al., 2015), particularly in light of the recent surge in large language models (LLMs), whose effectiveness heavily depends on the data they are trained on.
+Upon the acceptance of data-sharing agreements, personal health data is routinely collected by healthcare providers (Bartlett et al., 2018; Kim et al., 2015) as well as personal mobile apps and devices (Greene et al., 2019).
+This data is often accessed by third parties through data-sharing agreements that govern data donation practices (Trinidad et al., 2020).
+However, people frequently self-report that privacy concerns deter them from donating their sensitive information (Redmiles, 2020; Voigt et al., 2020; Silber et al., 2022; Greene et al., 2019; Kim et al., 2015).
+For instance, many individuals express discomfort with hospital policies that permit data-sharing agreements with third parties without obtaining explicit consent from the data owners (Lewis, 2019).
+In an effort to foster data collection and analysis while protecting the privacy of data contributors, there has been an increasing focus on deploying privacy-enhancing technologies (PETs) for data storage (Popa et al., 2011; Pappas et al., 2014), data processing (Arasu et al., 2013; Bajaj and Sion, 2011; Baumann et al., 2015; Gupta et al., 2019; Chowdhury et al., 2019; Al-Rubaie and Chang, 2019; Gaballah et al., 2024), and machine learning (ML) training (Wu et al., [n.
+d.]; De Viti et al., 2022).
+For example, CoVault (De Viti et al., 2022) enables multiparty analytics for epidemics-related datasets while preserving data confidentiality, and Anonify (Gaballah et al., 2024) supports decentralized, dual-level anonymity for secure medical data donation.
+However, in order for such PETs to impact the availability of data, the guarantees they offer must impact user privacy concerns significantly enough to alter their willingness to donate their data.
+In order to evaluate whether the guarantees offered by PETs ultimately impact people’s willingness to donate their data, at least three steps are necessary.
+First, we must be able to effectively explain guarantees (either general guarantees or those of a particular PET) to the people they aim to protect: end-users.
+Prior work has focused on doing so for particular PETs, chiefly differential privacy (DP) and end-to-end encryption (E2EE) (Cummings et al., 2021; Abu-Salma et al., 2018; Akgul et al., 2021; Bai, 2019; De Luca et al., 2016; Dechand et al., 2019; Abu-Salma et al., 2017; Wu and Zappala, 2018; Xiong et al., 2020; Nanayakkara et al., 2023).
+Second, using these explanations, we must evaluate whether the guarantees impact people’s privacy concerns (or expectations), and in turn, their willingness to share data.
+Prior work has done such evaluations using self-report studies focused on individual PETs, again primarily DP (Cummings et al., 2021; Bullek et al., 2017; Nanayakkara et al., 2023) or E2EE (Dechand et al., 2019; Bai et al., 2020; Abu-Salma et al., 2018)).
+Third, while rare, we must validate those results in the field to confirm that the effects observed replicate when studying actual vs.
+intended behavior (e.g., (Dooley et al., 2022b; Acquisti et al., 2013)).
+Instead of focusing on a singular PET and it’s accompanying guarantees, we focus on addressing steps one and two for the general security and privacy guarantees that recur across PETs:
   1. PG(1):
-    Anonymization [74, 96]: data is not linkable to its owner (as defined in [24]);
+    Anonymization (Murthy et al., 2019; Shahid et al., 2022): data is not linkable to its owner (as defined in (Chevrier et al., 2019));
   2. PG(2):
-    Access control [42, 51, 76, 105]: data is accessible only by authorized people, specified in the data collection agreement;
+    Access control (Greene et al., 2019; Hu et al., 2006; Ni et al., 2010; Tourani et al., 2017): data is accessible only by authorized people, specified in the data collection agreement;
   3. PG(3):
-    Data expiration [86, 93, 82]: data is discarded or inaccessible after a given expiration time, specified in the data collection agreement;
+    Data expiration (Politou et al., 2018; Sarkar et al., 2018; Perlman, 2005): data is discarded or inaccessible after a given expiration time, specified in the data collection agreement;
   4. PG(4):
-    Purpose restriction [32, 20]: data is only used for the stated collection purpose, specified in the data collection agreement;
-Additionally, and distinct from prior work, we investigate the role of auditing or verification processes increasingly introduced to produce public auditing results for PETs, specifically:
+    Purpose restriction (De Viti et al., 2022; Byun et al., 2005): data is only used for the stated collection purpose, specified in the data collection agreement;
+While previous research has explored how the combined presence of guarantees PG(1)-PG(4) affects data-sharing intentions (Aitken et al., 2016; Mori, 2016), our study takes a different approach by focusing on the impact of each guarantee individually.
+Our objective is to identify which specific protections or factors most significantly influence privacy expectations and, in turn, reported data-sharing intentions.
+Additionally, we investigate the role of auditing or verification processes increasingly introduced to produce public auditing results for PETs, specifically:
   1. AG(1):
-    Expert auditing [65, 113]: engaging an aggregator-selected external advisory board to audit the system to verify that the PET functions as described in the data collection agreement;
+    Expert auditing (Li et al., 2015; Wang et al., 2011): engaging an aggregator-selected external advisory board to audit the system to verify that the PET functions as described in the data collection agreement;
   2. AG(2):
-    Self auditing [49, 23]: granting anyone, including donors or external advisors appointed by them, the ability to perform such audits.
-Specifically, we investigate the role of these general PETs guarantees and audits of them in shaping privacy expectations and data sharing intentions in the context of medical data donation, asking:
+    Self auditing (Holt and Seamons, 2005; Chen et al., 2022): granting anyone, including donors or external advisors appointed by them, the ability to perform such audits.
+We examine the role of these PET guarantees and their audits in shaping privacy expectations and influencing data-sharing intentions within a specific context used in prior work (see e.g., (Cummings et al., 2021; Nanayakkara et al., 2023)): medical data donation.
+We ask:
   1. RQ(1):
@@
     How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?
-To address these questions, we conduct a vignette survey ($n=485$) following best practice methodology established by prior work [30, 2, 120].
+To address these questions, we conduct a vignette survey ($n=494$) following the best practice methodology established by prior work (Cummings et al., 2021; Abu-Salma et al., 2018; Xiong et al., 2020).
 Each survey respondent is presented with a hypothetical opportunity to donate their health data to help develop a treatment for a specific chronic disease, along with information on how their data will be protected (PG(1)–PG(4) enforced (or not) by AG(1)–AG(2)).
 Furthermore, we isolate the effect of PG(1)–PG(4) and AG(1)–AG(2) from external confounding factors identified by prior work:
-  • Egocentricity [104, 85, 43]: people making decisions depending on personal gain, i.e., whether they (or people close to them) are personally affected by the disease that will benefit from data donation;
-  • The data-collection entity [106, 48]: whether the collecting entity is a for-profit or a non-profit organization;
-  • Socio-demographics [47, 34, 106]: age, gender, education, donation history, and technical background.
+  • The data-collection entity (Trinidad et al., 2020; Hendricks-Sturrup and Lu, 2023): whether the collecting entity is a for-profit or a non-profit organization;
+  • Socio-demographics (Hargittai et al., 2020; Dooley et al., 2022a; Trinidad et al., 2020): age, gender, education, donation history, and technical background.
 We find that even when told nothing about PETs implemented by the entity, participants are 23% more likely, on average, to expect a non-profit to implement PG(1)–PG(4) and AG(1)–AG(2).
 As a result of these already high privacy expectations for non-profit organizations, we find that mentioning a specific privacy protection in the survey does not significantly enhance people’s willingness to donate towards non-profit entities: even when no privacy protection is explicitly mentioned, 89% of the participants are willing to donate to a non-profit entity.
-In contrast, for-profit entities need to effectively prove their privacy protections; indeed, explicitly mentioning privacy protections in the survey does increase privacy expectations of for-profit entities from 50% to the level of non-profit entities.
-Privacy expectations, in turn, influence willingness to donate.
-Furthermore, while the technical community has suggested external audits as a mechanism to enhance trust in PET implementation, our initial inquiry suggests that more work is needed to align expert and end-user perceptions of the effectiveness of such audits.
+In contrast, for-profit entities need to effectively demonstrate their privacy protections; indeed, explicitly mentioning privacy protections in the survey does increase privacy expectations of for-profit entities from 50% to the level of non-profit entities.
+Privacy expectations, in turn, influence the willingness to donate.
+Furthermore, while the technical community has suggested external audits as a mechanism to increase trust in PET implementation, our initial inquiry suggests that more work is needed to align the perceptions of the experts and end users about the effectiveness of such audits.
 In fact, the effect of audit statements on people’s willingness to donate is limited to a specific scenario involving for-profit entities and auditing to check that purpose restriction (PG(4)) is correctly implemented.
-We argue that it is thus essential that non-profit entities correctly implement data privacy, as any future data leak would result in a severe loss of trust.
-In fact, prior work [4] highlights that users value maintaining expected privacy more than obtaining privacy they did not already expect.
-In contrast, we highlight the risk of for-profit entities using statements about PETs to engage in ‘privacy washing’ [26] to raise privacy expectations in order to collect data.
-Simultaneously, we note our respondent’s savviness to the gaps in general PETs guarantees to protect against e.g., data breaches and point to the need for future work to investigate how to effectively set expectations about stronger emerging PETs guarantees in the face of end-user skepticism.
-2 Related Work
+We argue that it is critical for non-profit entities to rigorously implement data privacy measures, as any future data leak could lead to a significant loss of trust.
+Prior research (Acquisti et al., 2013) underscores that users place greater value on maintaining their expected privacy than on gaining additional privacy they did not initially anticipate.
+In contrast, we highlight the risk of for-profit entities engaging in ”privacy washing” (Cirucci, 2024), where statements about PETs are used to artificially raise privacy expectations to encourage data collection.
+At the same time, our findings reveal that respondents are perceptive of the limitations in general PET guarantees, particularly concerning protections against data breaches.
+This underscores the need for future research to explore how to effectively communicate the potential of stronger, emerging PET guarantees and address the skepticism of end-users.
+2. Related Work
 In this section, we examine prior work on people’s willingness to donate personal data, privacy concerns, the role of privacy-enhancing technologies (PETs), and educational or explanatory strategies related to PETs.
-Our review primarily focuses on literature specific to personal health data.
 […]
@@
 The analysis of personal health data is crucial for medical research, particularly in diagnosing emerging or rare diseases and developing new treatments.
-Indeed, the COVID-19 pandemic has further intensified the demand for personal health data [94].
-Despite the importance of this data, its sensitive nature poses significant obstacles to its collection [92, 21, 22, 71, 17, 35].
-Prior work indicates that individuals are generally willing to donate their data for altruistic purposes [98, 5, 84, 5, 95], though there is a notable reluctance when it comes to their health data specifically [70, 72].
-However, willingness to donate personal health data increases if the donor or their close family members are directly affected by the disease under investigation [41, 14].
+Indeed, the COVID-19 pandemic has further intensified the demand for personal health data (Scott et al., 2020).
+Despite the importance of this data, its sensitive nature poses significant obstacles to its collection (Fahey and Hino, 2020).
+Prior work indicates that individuals are generally willing to donate their data for altruistic purposes (Pilgrim and Bohnet-Joschko, 2022; Aitken et al., 2016; Seberger and Patil, 2021), though there is a notable reluctance when it comes to their health data specifically (Middleton et al., 2020; Mori, 2016).
+However, willingness to donate personal health data increases if the donor or their close family members are directly affected by the disease under investigation (Goodman et al., 2017; Bartlett et al., 2018).
 […]
 Privacy concerns.Privacy concernsPrivacy concerns.
-Concerns about the privacy and the misuse of donated data are prominent among potential donors [72].
-These concerns include the risk of being identified, discriminated against, or having their personal sensitive data misused or leaked [116, 5, 61, 68].
-A common source of these concerns is distrust of the receiving entity [116, 91, 50, 5, 100], often due to fears that the entity might violate consent terms by sharing data with unauthorized third parties [54, 70].
-This distrust is intensified when participants are unfamiliar with the recipient entity [116, 91, 50], especially if it is a governmental or for-profit organization [5, 100, 37, 48, 60].
-Beyond this intentional data misuse, there are concerns about data leakage caused by hackers or unintentional mishandling [55, 114].
-People are also reluctant to interact with entities with any history of data breaches, doubting their ability to safeguard sensitive information [108].
-[…]
-Privacy guarantees (PGs) and auditing.Privacy guarantees (PGs) and auditingPrivacy guarantees (PGs) and auditing.
-People typically expect the recipient entity to offer some level of privacy which depends on the type of data [55].
-Consequently, security experts propose that recipient entities use PETs that implement one or more PGs and an auditing process.
-We focus on a specific set of 4 PGs: data expiration, data anonymization, use restriction, access control, and auditing (see section 1).
-In particular, we examine whether offering a given PG influences people’s willingness to donate health data to the recipient entity.
-Furthermore, we focus on two different auditing processes: expert auditing and self auditing.
-To the best of our knowledge, no prior work surveys the effect of auditing (AG(1)–AG(2)) on people’s willingness to donate personal data, and only limited prior work investigates the effect of PG(1)–PG(4).
-For instance, Leon et al.
-[62] explored how data access, retention, and usage scope influence willingness to allow behavioral data collection, finding that retention and scope of use impact willingness to share personal information but that people were unwilling to pay more for sites that offered the preferred protections.
-Valdez and Ziefle [109] used conjoint analysis to study German participants’ willingness to donate health data to different recipients, with different quantified privacy risk, and differing purposes of data collection.
-They ran two conjoint analysis experiments, one expressing quantified risk of re-identification (for k-anonymity) and the other expressing privacy risk as a combination of sample size and exceptionalism (appropriate for differential privacy).
-They find similar results across both PETs: that the probability of identification and data recipient entity are the two factors that most influence data donation.
-Weiss et al.
-[115] examined participants’ willingness to share sensitive data based on varying epsilon thresholds of differential privacy, finding that the willingness to share is influenced by the purpose of the donation and amount of data subjects held by the entity.
-Other studies have examined public understanding of data retention and deletion [73], general data anonymization [45], and willingness to donate medical data if it is allowed to be used for secondary purposes [53].
-We build upon prior work to present a comprehensive overview of current public attitudes towards mainstream privacy guarantees in health data donation.
-Specifically, we investigate whether PG(1)–PG(4) and AG(1)–AG(2) influence users’ willingness to donate health data.
+Concerns about the privacy and the misuse of donated data are prominent among potential donors (Mori, 2016).
+These concerns include the risk of being identified, discriminated against, or having their personal sensitive data misused or leaked (Weitzman et al., 2012; Aitken et al., 2016; Kim et al., 2015; McCormick and Hopkins, 2021).
+A common source of these concerns is distrust of the receiving entity (Weitzman et al., 2012; Howe et al., 2018; Aitken et al., 2016; Stockdale et al., 2019), often due to fears that the entity might violate consent terms by sharing data with unauthorized third parties (Kacsmar et al., 2022; Middleton et al., 2020).
+This distrust is intensified when participants are unfamiliar with the recipient entity (Weitzman et al., 2012; Howe et al., 2018), especially if it is a governmental or for-profit organization (Aitken et al., 2016; Stockdale et al., 2019; Garrison et al., 2016; Hendricks-Sturrup and Lu, 2023; Khanra, 2023; American Medical Association, 2022).
+Beyond this intentional data misuse, there are concerns about data leakage caused by hackers or unintentional mishandling (Seh et al., 2020).
+People are also reluctant to interact with entities with any history of data breaches, doubting their ability to safeguard sensitive information (Seh et al., 2020).
 […]
@@
 PETs aim to address user privacy concerns.
-While various PETs are available for health data donation environments [52], prior work indicates that only participants with online privacy literacy tend to have more trusting beliefs in PETs [46].
+While various PETs are available for health data donation environments (Jordan et al., 2022), prior work indicates that participants with greater online privacy literacy tend to have more trusting beliefs in PETs (Harborth and Pape, 2020).
 Studies focusing on sensitive personal data, including health data, reveal similar findings.
 For instance, Dechand et al.
-[33] found that most participants do not trust encryption because they do not understand how it works: e.g., many believe that service providers can bypass encryption to gain unauthorized data access.
+(Dechand et al., 2019) found that most participants do not trust encryption because they do not understand how it works: e.g., many believe that service providers can bypass encryption to gain unauthorized data access.
 Furthermore, Lerner et al.
-[63] noted that some participants doubt the existence of true privacy.
-Non-experts may misunderstand privacy technologies [118, 78, 33], e.g., viewing encryption merely as restrictive access control [118] or confusing it with data encoding [3].
-Additionally, non-expert end users may struggle to understand the consequences of inadequate privacy protections [103, 111, 2, 40, 83, 110, 55].
+(Lerner et al., 2022) noted that some participants doubt the existence of true privacy.
+Non-experts may misunderstand privacy technologies (Wu and Zappala, 2018; Oates et al., 2018; Dechand et al., 2019), e.g., viewing encryption merely as restrictive access control (Wu and Zappala, 2018) or confusing it with data encoding (Abu-Salma et al., 2017).
+Additionally, non-expert end users may struggle to understand the consequences of inadequate privacy protections (Tang et al., 2021; Velykoivanenko et al., 2022; Abu-Salma et al., 2018; Gerber et al., 2019; Pham and Castro, 2019; Vaniea et al., 2014).
 As a result, even when privacy protections are present, people may have risk expectations that are misaligned with reality.
@@
 Explaining PETs.Explaining PETsExplaining PETs.
-To address these misunderstandings, technologists and researchers have worked to explain PETs to the public to address their unfamiliarity with privacy concepts [29].
-Prior work includes efforts to explain secure communication, such as end-to-end encrypted messaging, and differential privacy [2, 6, 10, 30, 31, 6, 33, 3, 118, 120, 75].
+To address these misunderstandings and evaluate the impact of PETs on downstream factors such as privacy concerns (or expectations) and willingness to share data, technologists and researchers have worked to explain PETs to the public to address their unfamiliarity with privacy concepts (Cranor et al., 2006).
+Prior work focuses heavily on E2EE and DP (Abu-Salma et al., 2018; Bai, 2019; Cummings et al., 2021; De Luca et al., 2016; Akgul et al., 2021; Dechand et al., 2019; Abu-Salma et al., 2017; Wu and Zappala, 2018; Xiong et al., 2020; Nanayakkara et al., 2023).
 These efforts remain ongoing, as effectively and scalable setting privacy expectations remains a challenge.
-Methods found effective to explain PETs include visualizations [101], mental models [99, 66], nutrition labels [59], metaphors [56], short statements [80], and privacy games [107, 57].
+Methods found effective to explain PETs include visualizations (Stransky et al., 2021), mental models (Stewart and Lacey, 2012; Lin et al., 2012), nutrition labels (Kelley et al., 2010), metaphors (Karegar et al., 2022), short statements (Papacharissi and Fernback, 2005), and privacy games (Tupsamudre et al., 2018; Karl and Tao, 2023).
 In this work we use short statements to briefly explain the relevant guarantees in order to evaluate the impact of these guarantees on participants’ privacy expectations and willingness to donate medical data.
-3 Methodology
-As mentioned in section 1, we address research questions RQ1 and RQ2 through a user survey.
-This section details our survey’s design (subsection 3.1) and the cognitive interview process used to verify respondents’ comprehension of the presented scenarios (subsection 3.2).
-We also describe our participant selection strategy (subsection 3.3) and the limitations of our methodology (subsection 3.5).
-3.1 Survey Structure
-Our survey design is shown in Figure 1.
-Figure 1: Survey design.
-Each participant is presented with a donation scenario with a survey introduction with two variations: recipient entity and disease.
+[…]
+Privacy guarantees (PGs) and auditing.Privacy guarantees (PGs) and auditingPrivacy guarantees (PGs) and auditing.
+Prior work on the effect of PETs and their accompanying PGs typically focuses on the effect of a combination of guarantees offered by a specific PET (e.g., DP (Valdez and Ziefle, 2019; Cummings et al., 2021) or E2EE (Dechand et al., 2019)) on privacy concerns .
+A more limited subset of studies evaluate the impact on willingness to share data (Valdez and Ziefle, 2019).
+Instead of focusing on a specific PET and it’s combined accompanying PGs we seek to examine and compare four PGs that are shared across many PETs: data expiration, data anonymization, use restriction, access control, and auditing (see § 1).
+In particular, we examine whether offering a given PG influences people’s willingness to donate health data to the recipient entity.
+Furthermore, we focus on the effect of two different auditing processes – expert auditing and self auditing – on enhancing privacy expectations and willingness to share data.
+Other studies have examined public understanding of data retention (Murillo et al., 2018) and general data anonymization (Haddow et al., 2011), and one study examines willingness to donate medical data if it is allowed to be used for secondary purposes (Juga et al., 2021).
+To our knowledge, no prior work focuses on comparing the effects of the four core PGs we study nor the effects of auditing PETs on user privacy expectations or willingness to donate data.
+3. Methodology
+As mentioned in § 1, we address research questions RQ1 and RQ2 through a user survey.
+In this section, we discuss the ethical considerations of our study, outline the survey design (§ 3.2), and describe the cognitive interview process used to ensure participants’ understanding of the presented scenarios (§ 3.3).
+We also describe our participant selection strategy (§ 3.4) and the limitations of our methodology (§ 3.6).
+3.1. Ethics
+Before we began our research, our ethics review board approved all of the study methods, including the cognitive interviews and the survey.
+A consent form detailing the data storage, handling, and survey compensation is presented at the beginning of the survey.
+Participants can opt out of disclosing any demographic data without influencing their compensation in the survey.
+Each participant received $1.20\$$ upon completing the survey.
+To determine this figure, we ran a test of 20 participants, who took 5m56s (median) to complete it.
+Thus, the actual compensation was $12\$$/hr, which aligns with Prolific recommendations.
+For the signup survey of the cognitive interview, we set the study name to “Sign Up for Paid Interview on Data Donation” with a payment of $1.25\$$ ($13.68\$$/hr based on a loose completion estimate of 5 minutes).
+However, most respondents refused our interview request.
+Thus, in order to encourage participants interested in attending the interview to complete the screener, we lowered the compensation for the screening survey to $9.96\$$/hr.
+The interview reward is $15\$$, and all interviews lasted at most thirty minutes.
+The survey is conducted through Qualtrics (Qualtrics, 2020), with the IP address collection feature disabled to protect respondents’ privacy.
+Interviews are conducted via Zoom (zoom, 2024), and participants are instructed to avoid using their real names when joining the meeting to ensure anonymity.
+Research participants are anonymously recruited through Prolific (Prolific, 2024), as described in greater detail below.
+3.2. Survey Structure
+Figure 1.
+Survey design.
+Each participant is presented with a donation scenario and a survey introduction with two variations: recipient entity and disease.
 statements (PG(1)–PG(4) and AG(1)–AG(2)) are also presented in the donation scenario.
@@
 When the control statements are presented, the trust (dashed) is omitted.
-The willingness to donate and trust questions (dark shaded) are associated with an open-text question.
+The willingness to donate and trust questions (dark-shaded) are associated with an open-text question.
 The scenario understanding is used as a filter.
 Participants who did not understand the donation scenario were excluded from the analysis.
-First, we present to the respondents a suvery introduction, which we detail in Figure 2.
+A sample screen that participants would encounter is provided in Figure 6.
+Our survey design is shown in Figure 1, and the full survey can be found in Appendix A. First, we present to the respondents a suvery introduction, which we detail in Figure 2.
 Survey introduction
@@
 Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
-Figure 2: Survey Introduction. The entity type (for-profit or non-profit) and the disease (selected from a list of common chronic diseases) are randomly selected.
+Figure 2. Survey Introduction. The entity type (for-profit or non-profit) and the disease (selected from a list of common chronic diseases) are randomly selected.
 We present a donation scenario using this introduction: we want to assess respondents’ willingness to donate their health data to a recipient entity developing a new treatment for a specific disease.
-The type of entity (for-profit or non-profit) and the disease (cancer, diabetes, heart failure, high blood pressure, and stroke) are randomly selected; the disease options are taken from a list of common chronic diseases published by the Centers for Medicare and Medicaid Services (CMS) [36].
+The type of entity (for-profit or non-profit) and the disease (cancer, diabetes, heart failure, high blood pressure, and stroke) are randomly selected; the disease options are taken from a list of common chronic diseases published by the Centers for Medicare and Medicaid Services (CMS) (for Medicare and Services, 2024).
 Then, we present to the respondents either a control (no statement) or an experimental statement.
@@
 Privacy statements are uniformly selected at random from a pool of four privacy statements; auditing statements are selected at random from two auditing statements.
-We have written these statements to align with PG(1)–PG(4) (see section 1), and we list them in Figure 3.
+We have written these statements to align with PG(1)–PG(4) (see § 1), and we list them in Figure 3.
 We further distinguish the auditing guarantee into two different auditing statements to align with AG(1)–AG(2), and we list them in Figure 4.
@@
 No privacy-preserving guarantee is at all mentioned.
-Figure 3: Privacy statements PG(1)–PG(4). One of the five statements (including the control statement) is randomly presented in the donation scenario immediately after the survey introduction.
+Figure 3. Privacy statements PG(1)–PG(4). One of the five statements (including the control statement) is randomly presented in the donation scenario immediately after the survey introduction.
 AG(1): Expert auditing — the implementation of PET will be verified by experts.
@@
 No auditing is at all mentioned.
-Figure 4: Auditing statements AG(1)–AG(2). One of the three statements (including the control statement) is randomly presented in the donation scenario after the privacy statement.
+Figure 4. Auditing statements AG(1)–AG(2). One of the three statements (including the control statement) is randomly presented in the donation scenario after the privacy statement.
 Our survey questions assess respondents’ self-reported:
@@
   • trust that the recipient entity will implement the protection described by the privacy statement: “I trust the entity will handle my data as described.” (4-point Likert scale Strongly agree - Strongly disagree).
-  • demographics and experiences, including age, gender, education level, donation history, and technical background.
+  • demographics and experiences, including age, gender, education level, donation history, technical background (i.e., education background and job field relationships with computer technologies), and egocentricity(i.e., whether respondents or their close relatives have the disease, as defined in (Grgić-Hlača et al., 2020; Plane et al., 2017)).
 Anonymization
@@
 I will be able to hire someone to verify that my medical record is protected as described.
-Figure 5: To measure privacy expectations respondents reported their agreement with each statement listed above on a 4-point Likert Scale from “Very Likely” to “Very Unlikely”.
-We also ask respondents to explain their responses to the willingness to donate and trust questions.These explanations are collected through open-ended questions: “Why your are willing (or unwilling) to share your medical record with this entity?” and “Please explain why you do (or do not) trust that the entity will handle your data as described.” (We note that the scenario understanding and trust questions are only asked when a non-control privacy statement is presented).
-3.2 Questionnaire Refinement
-The final questionnaire was refined through multiple iterations.
-Before launching the survey, we prescreened 49 users on the Prolific online platform [1] by asking them to complete a short demographics questionnaire and indicate their willingness to participate in a cognitive interview.
-Then, we conducted 17 cognitive interviews to ensure participants understood the questionnaire correctly.
-Each interview involved presenting the questionnaire to participants and actively seeking feedback on their interpretation of the data donation scenario, and the privacy and auditing statements.
+Figure 5. To measure privacy expectations respondents reported their agreement with each statement listed above on a 4-point Likert Scale from “Very Likely” to “Very Unlikely”.
+We also ask respondents to explain their responses to the willingness to donate and trust questions.
+These explanations are collected through open-ended questions: “Why you are willing (or unwilling) to share your medical record with this entity?” and “Please explain why you do (or do not) trust that the entity will handle your data as described.” We note that the scenario understanding and trust questions are only asked when a non-control privacy statement is presented.
+A sample screen that participants would encounter is provided in Figure 6.
+Imagine that a non-profit organization (e.g., university, non-profit research institute) wants to develop a new treatment for Hypertension (High blood pressure).
+They need medical data from people with and without Hypertension (High blood pressure) to develop the treatment.
+They ask you to donate your medical record to help develop the treatment.
+Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
+The non-profit organization uses a privacy-preserving technology to ensure that your data is deleted after a set period of time.
+You will be able to choose from a list of time period options how long your data will be stored.
+Anyone interested, including you and experts you trust, will be able to verify that the privacy-preserving technology is working as described.
+Anyone, including you and experts you trust, can post their verification results publically.
+How would you rate your understanding of the above scenario? Fully understand Mostly understand Partially understand Not understand
+Figure 6.
+Sample survey screen that composed of three segments.
+The first segment presents the survey introduction, which varies the disease and the entity among all participants.
+The second segment includes the randomized privacy statement and auditing statement, with color coding applied to emphasize these elements.
+The third segment contains the survey question that participants are required to answer.
+3.3. Survey Refinement
+The final survey was refined through multiple iterations.
+Before launching the survey, we prescreened 49 users on the Prolific online platform (Prolific, 2024) by asking them to complete a short demographics survey and indicate their willingness to participate in a cognitive interview.
+Then, we conducted 17 cognitive interviews to ensure participants understood the survey correctly.
+Each interview involved presenting the survey to participants and actively seeking feedback on their interpretation of the data donation scenario, and the privacy and auditing statements.
 We additionally presented all statements to the interviewees and asked for their understanding and perception of each of them.
-We continued refining the questionnaire through cognitive interviews until no further constructive feedback was received.
+We continued refining the survey through cognitive interviews until no further constructive feedback was received.
 In the final interview, we observe a clear understanding of the statements and the donation scenario.
-For example, the participant described anonymization (PG(1)) as "eliminating any sort of demographic or personal information affiliated with your data that could be associated with you if the data were to somehow be leaked" and access control (PG(2)) as "some sort of database that only certain individuals have access to by using a passcode and certain credentials to to log in." Each Prolific user who completed a cognitive interview was paid \$15.
-3.3 Survey Sampling
-Following best practice in security research [102], we recruited 560 respondents via Prolific.
-We restricted the respondents to adults residing in the U.S.
-to obtain focused results.
-Then, we ensured a balanced gender distribution: 49.11% men, 48.57% women, 1.6% non-binary, and 0.7% unwilling to disclose.
-Each participant received $0.83\$$ upon completing the questionnaire; the median completion time was around 7 minutes.
-Out of these 560 respondents, we sequentially excluded the following from analysis: 13 that reported non-binary gender [^2] As we lack statistical power to draw meaningful conclusions about the privacy sentiments of the non-binary community specifically, we do not include these data in our analysis.
-However, this in no way diminishes the importance of understanding the unique and valuable insights of this population, which should be explored with either larger samples (yielding statistical power) or targeted research specific to their viewpoints and perspectives [39, 38].; 31 that submitted incomplete responses; 20 that failed our attention check (subsection 3.1); and 11 that indicated they did not understand the survey scenario (using scenario understand).
-Thus, our final dataset comprises 485 respondents, whose demographics are reported in Table I.
-How many respondents were assigned to each condition is reported in Table II.
+For example, the participant described anonymization (PG(1)) as ”eliminating any sort of demographic or personal information affiliated with your data that could be associated with you if the data were to somehow be leaked” and access control (PG(2)) as ”some sort of database that only certain individuals have access to by using a passcode and certain credentials to to log in.” Our final survey can be found in Appendix A.
+3.4. Survey Sampling
+Following the recommendation of (Tang et al., 2022), we recruited participants through Prolific, gathering responses from 560 individuals.
+To ensure focused results, we restricted respondents to adults residing in the U.S.
+and requested a gender-balanced distribution.
+Our recruitment included 272 men, 275 women, 9 nonbinary individuals, and 4 participants who chose not to disclose their gender.
+Of these 560 respondents, we sequentially excluded the following from the analysis: 35 that submitted incomplete responses; 20 that failed our attention check (§ 3.2); and 11 that indicated they did not understand the survey scenario (using scenario understand).
+Thus, our final dataset comprises 494 respondents.
+To enhance statistical power, we limit the levels of age and education to binary groups based on the data distribution among our participants.
+We report their demographics in Table 1 and the number of respondents assigned to each condition in Table 2.
 Description Category $n$
-Age $\geq$ 40 219 45.15%
-< 40 166 54.85%
-Gender Woman 254 52.37%
-Man 231 47.63%
-Education w/ BSc degree 352 72.58%
-w/o BSc degree 133 27.42%
-Technical Background Yes 128 26.39%
-No 357 73.61%
-Donation History Yes 55 11.34%
-No 430 88.66%
-TABLE I: Participant demographics. We note that the second category of each demographical attribute (e.g., no technical background) is considered the baseline scenario during analysis.
+Age 18 - 29 120 24.3%
+30 - 49 222 44.9%
+50 - 64 111 22.5%
+65+ 41 8.3%
+Gender Woman 254 51.42%
+Man 231 46.76%
+Others 9 1.82%
+Education B.S. or above 358 72.47%
+Up to H.S. 136 27.53%
+Technical Background Yes 129 26.11%
+No 365 73.89%
+Donation History Yes 56 11.34%
+No 438 88.66%
+Egocentricity Yes 201 40.69%
+No 293 59.31%
+Table 1.
+Participant demographics.
+We note that the second category of each demographical attribute (e.g., no technical background) is considered the baseline scenario during analysis, except for “age”, which is treated numerically.
 For-Profit Non-Profit
 AG(1) AG(2) Ctrl. AG(1) AG(2) Ctrl.
-PG(1) 22 18 19 19 19 18
-PG(2) 17 18 19 18 18 19
-PG(3) 18 20 18 18 17 17
-PG(4) 21 18 18 20 18 19
-Ctrl. - - 19 - - 20
-TABLE II: Number of respondents assigned to each condition (i.e., who saw each privacy statement and auditing statement, or no privacy statement (bottom row) or no auditing statement (columns 4 and 6).
-3.4 Analysis
-We analyzed the open-text questions about willingness to donate and trust (see subsection 3.1) using inductive-thematic open coding.
-Two researchers independently coded each entry and generated a codebook from a random sample of at least 100 (20.6%) responses.
+PG(1) 22 18 19 19 20 18
+PG(2) 17 21 19 18 18 19
+PG(3) 18 20 18 19 19 17
+PG(4) 21 18 19 20 18 19
+Ctrl. - - 19 - - 21
+Table 2. Number of respondents assigned to each condition (i.e., who saw each privacy statement and auditing statement, or no privacy statement (bottom row) or no auditing statement (columns 4 and 7).
+3.5. Analysis
+We analyzed the open-text questions about willingness to donate and trust (see § 3.2) using inductive-thematic open coding.
+Two researchers independently coded each entry and generated a codebook from a random sample of at least 100 (20.2%) responses.
 Then, they composed a final codebook and double-coded all responses.
-Since all responses were double-coded and inconsistencies were resolved, we do not report inter-rater reliability (IRR) [69].
-For all quantitative analyses, we binarized the responses to the willingness to donate, privacy expectations, and trust questions (see subsection 3.1)) to get ‘True’ if the respondents selected ‘Likely’ or ‘Very Likely’, and ‘False’ otherwise.
+Since all responses were double-coded and inconsistencies were resolved, we do not report inter-rater reliability (IRR) (McDonald et al., 2019).
 In addition to presenting a descriptive analysis detailing the distribution of responses on the survey items that address our research questions, we construct logistic regression models to analyze factors related to two dependent variables: privacy expectations (RQ1) and willingness to donate (RQ2).
@@
 For RQ2, the dependent variable was willingness to donate, and the independent variables were the presence of a given privacy statement, the privacy expectation for each guarantee, as well as demographics and experiences.
+We collapsed the independent variables into a binary measure, designating responses of “Likely” or “Very Likely” as True and all others as False.
+We did so to avoid the ambiguity introduced by intermediate scale points and to simplify the statistical model.
 We categorized education into two groups: with and without a bachelor’s degree.
-We also categorized age into two groups and cut to the age of forty.
 We took “no bachelor’s degree” and “less than forty years old” as the reference categories, respectively.
@@
 We built separate regression models for the two recipient entities: for-profit and non-profit.
-3.5 Limitations
-Although the four privacy-preserving guarantees examined in this study (section 1) span many PETs, they are not fully comprehensive and do not capture the full complexity of PETs nor of real-world threats, which may involve privacy issues beyond the scope of our investigation.
-Moreover, despite our efforts to mitigate misunderstandings through rigorous cognitive interviews, filtering out respondents who reported not fully understanding the scenario, and controlling for privacy expectations in our statistical analyses, the brief descriptions of each guarantee might have led to partial or incorrect understanding among respondents, potentially influencing our results.
-While this study employs vignette scenarios that can effectively mirror real-world behaviors [90], discrepancy between participants’ expressed privacy concerns and their actual behaviors may exist (i.e., privacy paradox [13]).
-Furthermore, statistically-significant results do not inherently guarantee the absence of underlying biases that could skew the findings.
-In fact, various unmeasured factors that were not controlled for could influence the interpretations of the results.
-Finally, our respondents were recruited via Prolific, and thus, may not accurately reflect the diversity of the entire U.S.
-population, let alone the varied perspectives of individuals from other countries.
-This limitation challenges the generalizability of our findings.
-4 Results
-4.1 RQ1: Understanding of PGs
-Next, we answer the first research question (RQ1, see section 1): “How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?".
+3.6. Limitations
+Although the four privacy-preserving guarantees examined in this study (§ 1) cover a broad range of PETs, they are not fully comprehensive and do not capture the full complexity of PETs or real-world threats, which may involve privacy issues beyond the scope of our investigation.
+The randomly presented disease may also influence the results.
+Furthermore, our approach may not fully reflect the intricacies of real-world donation scenarios, where multiple guarantees often coexist.
+While such complexities lie outside the primary focus of our research, qualitative investigations have been conducted in prior work (Aitken et al., 2016; Mori, 2016), and quantitative exploration of these aspects remains an interesting direction for future research.
+Despite our efforts to mitigate misunderstandings—through rigorous cognitive interviews, filtering respondents who reported not fully understanding the scenarios, and controlling for privacy expectations in our statistical analyses—the brief descriptions of each guarantee might still have led to partial or incorrect understanding among respondents.
+This limitation may have influenced our results.
+Additionally, while our study employs vignette scenarios that effectively mirror real-world behaviors (Redmiles et al., 2017), there may still be discrepancies between participants’ expressed privacy concerns and their actual behaviors, a phenomenon commonly referred to as the privacy paradox (Barth and de Jong, 2017).
+Furthermore, statistically significant results do not inherently eliminate the risk of underlying biases that could skew the findings.
+Various unmeasured factors not accounted for in our analysis might have influenced the interpretations of the results.
+Lastly, our respondents were recruited via Prolific, which may limit the generalizability of our findings.
+The sample might not fully represent the diversity of the U.S.
+population, nor does it capture the perspectives of individuals from other countries, highlighting a key limitation of this work.
+4. Results
+4.1. RQ1: Understanding of PGs
+Next, we answer the first research question (RQ1, see § 1): “How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?”.
 We analyze and report the survey responses and evaluate the rationales for the responses.
-[…]
-Analysis Set Up.Analysis Set UpAnalysis Set Up.
+4.1.1. Analysis Setup
 We compare privacy expectations on PG(1)—PG(4) and AG(1)–AG(2) among groups of respondents who received different privacy statements (listed in Figure 3) and auditing statements (listed in Figure 4).
-In Table III, we show the privacy expectations of respondents who were shown different privacy statements.
+In Figure 7, we show the privacy expectations of respondents who were shown different privacy statements.
 To assess the statistical significance of the descriptive quantitative results presented above, we used binomial logistic regression to analyze the relationship between the presence of a privacy statement in the scenario and respondents’ privacy expectations.
-We summarize results, which we distinguish for non-profit and for-profit entities, in Table IV.
-Relationship between Privacy Statement Shown and Privacy Expectation Statement Shown Overall For-Profit Non-Profit PG(1) Anonymization 0.68 0.68 0.67 + AG(1) Expert audit 0.63 0.59 0.68 + AG(2) Self audit 0.65 0.67 0.63 Control 0.38 0.26 0.50 PG(2) Access control 0.50 0.58 0.42 + AG(1) Expert audit 0.51 0.35 0.67 + AG(2) Self audit 0.50 0.56 0.44 Control 0.49 0.37 0.60 PG(3) Data expiration 0.83 0.67 1.00 + AG(1) Expert audit 0.92 0.83 1.00 + AG(2) Self audit 0.86 0.85 0.88 Control 0.41 0.37 0.45 PG(4) Purpose restriction 0.78 0.72 0.84 + AG(1) Expert audit 0.71 0.62 0.80 + AG(2) Self audit 0.81 0.67 0.94 Control 0.56 0.37 0.75 AG(1) Expert audit 0.82 0.78 0.87 Control 0.62 0.58 0.66 AG(2) Self audit 0.53 0.50 0.57 Control 0.26 0.25 0.28
-TABLE III: Percentage of respondents who had a positive privacy expectation when shown a particular privacy statement in their donation scenario.
+We summarize results, which we distinguish for non-profit and for-profit entities, in Figure 8.
+Figure 7.
+Percentage of respondents who had a positive privacy expectation when shown a particular privacy statement in their donation scenario.
 The values in the table are the percentage of respondents in a given condition who had the privacy expectation (see Figure 5) that corresponded to the privacy statement they were presented.
-The 2nd column (“overall") reports results across both entities (for-profit and non-profit), the 3rd column reports the results of the subset in the for-profit entity conditions, and the 4th column reports the results of the subset in the non-profit entity conditions.
-(Recall that agreement has been binarized as discussed in subsection 3.4).
-For example, the left-most and top-most numerical cell indicates that 68% of participants in both entity scenarios who were shown PG(1) alone – with no auditing statement – expected that their data would be anonymized.
-Statement Shown OR 95% CI $p$-value
-For-Profit PG(1) Anonymization 5.07 [1.69, 17.52] 0.006
-PG(2) Access control 1.71 [0.6, 5.23] 0.325
-PG(3) Data expiration 6.29 [2.09, 20.44] 0.001
-PG(4) Purpose restriction 3.43 [1.19, 10.59] 0.026
-AG(1) Expert audit 2.59 [1.33, 5.20] 0.006
-AG(2) Self audit 3.04 [1.59, 5.93] <0.001
-Non-Profit PG(1) Anonymization 1.95 [0.69, 5.57] 0.208
-PG(2) Access control 1.27 [0.45, 3.61] 0.651
-PG(3) Data expiration 16.7 [3.64, 120.49] <0.001
-PG(4) Purpose restriction 2.04 [0.55, 7.11] 0.266
-AG(1) Expert audit 3.41 [1.59, 7.85] 0.002
-AG(2) Self audit 3.41 [1.79, 6.61] <0.001
-TABLE IV: Effect of presenting different privacy statements on respondents’ privacy expectations toward the two recipient entities. OR: odd ratio; CI: confidence interval.
-[…]
-Response Distribution and Statistical Analysis Results.Response Distribution and Statistical Analysis ResultsResponse Distribution and Statistical Analysis Results.
+The 3rd column (“overall”) reports results across both entities.
+For example, the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios who were shown PG(1) alone – with no auditing statement – expected their data to be anonymized.
+The agreements are binarized here to present the overall tendency.
+Figure 8.
+Odd ratios and 95% confidence intervals of presenting different privacy statements on the respondents’ privacy expectations toward the two kinds of recipient entities, for-profit and non-profit.
+4.1.2. Response Distribution and Statistical Analysis Results
 We find that privacy expectations differ based on the data-collecting entity.
 In the control group, which was not shown any privacy statement, 26%-37% of respondents expected for-profit entities to provide anonymization, data expiration, access control, use restriction to a specific purpose, expert auditing, and self-auditing.
-In contrast, a higher percentage of respondents (45%-75%) expected non-profit entities to employ these mechanisms.
-For instance, we note that respondent P430 (who received a for-profit scenario) expected privacy protections even though they were in the control group (and thus received no statement about privacy protection): "I also assume that the data is looked at in the aggregate and likely no one at the company knows me.".
-In the same situation, respondent P53 (who received a non-profit scenario instead) made a stronger assumption: "I trust that the organization will uphold strict privacy and ethical standards".
-In the for-profit scenario, all privacy statements, except for the access-control statement, significantly increase the corresponding privacy expectations ($p$-value < $0.05$).
-In the non-profit scenario, where expectations are already high, only the privacy statement about data expiration significantly increases expectations (from 60% to 100%).
-Comparing the for-profit and non-profit subsets of the descriptive results (2nd and 3rd column of Table III), we observe that presenting a privacy statement was particularly beneficial in for-profit scenarios.
-Indeed, the presence of a privacy statement raised respondents’ privacy expectations to nearly the same level as non-profit scenarios for anonymization (68% and 67%) and access control (58% and 42%).
-It also reduced the gap in expectations for data expiration (67% and 100%)) and purpose restriction (72% and 84%)).
-In both scenarios, the auditing statements increase the expectation that a given privacy guarantee will be audited ($p$-value < $0.05$).
+In contrast, a higher percentage of respondents (43%-76%) expected non-profit entities to employ these mechanisms, except for self-auditing (AG2).
+For instance, we note that respondent P430 (who received a for-profit scenario) expected privacy protections even though they were in the control group (and thus received no statement about privacy protection): ”I also assume that the data is looked at in the aggregate and likely no one at the company knows me”.
+In the same situation, respondent P53 (who received a non-profit scenario instead) made a stronger assumption: ”I trust that the organization will uphold strict privacy and ethical standards”.
+In the for-profit scenario, all privacy statements, except for the access-control statement (PG2), significantly increase the corresponding privacy expectations ($p$-value ¡ $0.05$).
+In the non-profit scenario, where expectations are already high, only the privacy statement about data expiration (PG3) significantly increases expectations (from 57% to 100%).
+Comparing the for-profit and non-profit subsets of the descriptive results (1st and 2nd column of Figure 7), we observe that presenting a privacy statement was particularly beneficial in for-profit scenarios.
+Indeed, the presence of a privacy statement raised respondents’ privacy expectations to nearly the same level as non-profit scenarios for anonymization (PG1, 68% and 67%) and access control (PG2, 58% and 42%).
+It also reduced the gap in expectations for data expiration (PG3, 67% and 100%)) and purpose restriction (PG4, 74% and 84%)).
+In both scenarios, the auditing statements increase the expectation that a given privacy guarantee will be audited ($p$-value ¡ $0.05$).
 However, we note that incorporating an audit statement does not markedly change privacy expectations.
-The statistical significance of the auditing statements in Table IV indicates that respondents shown the audit statement correctly expected the corresponding type of audit, while respondents that were not shown the auditing statement did not expect this protection.
-As observed in  Table III, presenting an auditing statement does not increase the number of respondents who expect the privacy guarantee implied by the privacy statement they were shown, despite the stronger assurances that audits imply.
+The statistical significance of the auditing statements in Figure 8 indicates that respondents shown the audit statement correctly expected the corresponding type of audit, while respondents that were not shown the auditing statement did not expect this protection.
+As observed in  Figure 7, presenting an auditing statement does not increase the number of respondents who expect the privacy guarantee implied by the privacy statement they were shown, despite the stronger assurances that audits imply.
+4.1.3. Qualitative Analysis of Privacy Expectations via Open-Answer Responses
+Across for-profit and non-profit scenarios, 69% and 85% of respondents, on average, expressed that their positive privacy expectations were formed based on the presence of the privacy statement, belief in privacy obligations of the entity, trust in the entity, or trust in the auditing process.
 […]
-Qualitative Analysis of Privacy Expectations via Open-Answer Responses.Qualitative Analysis of Privacy Expectations via Open-Answer ResponsesQualitative Analysis of Privacy Expectations via Open-Answer Responses.
-Across for-profit and non-profit scenarios, 69% and 85% of respondents, on average, expressed that their positive privacy expectations were formed based on the presence of the privacy statement, belief in privacy obligations of the entity, trust in the entity, or trust in the auditing process.
-Satisfaction with the privacy guarantees.
+Satisfaction with the privacy guarantees..Satisfaction with the privacy guarantees.Satisfaction with the privacy guarantees..
 16.3% and 20.8% of respondents in the for-profit and non-profit scenarios, respectively, were convinced by the received privacy statement.
-Some respondents, like P56, vaguely stated that the policy looks promising: "I see their policy, and they have to follow their own policy.".
+Some respondents, like P56, vaguely stated that the policy looks promising: ”I see their policy, and they have to follow their own policy”.
 Others felt explicitly safer and trusted the statement.
-For example, P494 noted: "Trust is very important when it comes to medical data.
+For example, P494 noted: ”Trust is very important when it comes to medical data.
 I believe the organization has privacy policies that outline that they collect and will use my data.
-I also believe the organization will employ security measures to safeguard data".
-Similarly, P555 highlighted the presence of the privacy statement: "(…) purposely states there is software in place to conserve privacy." and P516 felt their information was safe because of it: "(…) with the privacy protection in place, they are isolating the data they need while basically "throwing out" the rest by putting in under that protection.
-In essence, my information is safe, and they’re only using what they said they’d use".
-P208 also felt safer contributing to research knowing the privacy mechanisms were in place: "I want to be able to contribute to research to better improve cancer treatment, and I feel safe if my data is protected through the mechanisms above".
-Belief in legal and reputational obligations.
+I also believe the organization will employ security measures to safeguard data”.
+Similarly, P555 highlighted the presence of the privacy statement: ”(…) purposely states there is software in place to conserve privacy”.
+and P516 felt their information was safe because of it: ”(…) with the privacy protection in place, they are isolating the data they need while basically ”throwing out” the rest by putting in under that protection.
+In essence, my information is safe, and they’re only using what they said they’d use”.
+P208 also felt safer contributing to research knowing the privacy mechanisms were in place: ”I want to be able to contribute to research to better improve cancer treatment, and I feel safe if my data is protected through the mechanisms above”.
+[…]
+Belief in legal and reputational obligations..Belief in legal and reputational obligations.Belief in legal and reputational obligations..
 16.3% and 10.6% of respondents in the for-profit and non-profit scenarios, respectively, believed that entities are forced to protect the donated data due to legislative requirements and reputational concerns.
-For example, P28 stated regarding for-profits: "A for-profit organization wouldn’t want to violate HIPAA, HITECH laws".
-Additionally, P117 wrote "I would trust them to do the right thing so they won’t face lawsuits".
-Regarding non-profits, P141 noted: "I trust that the law will restrict any data leaks to third parties".
-Trust in the recipient entity.
+For example, P28 stated regarding for-profits: ”A for-profit organization wouldn’t want to violate HIPAA, HITECH laws”.
+Additionally, P117 wrote ”I would trust them to do the right thing so they won’t face lawsuits”.
+Regarding non-profits, P141 noted: ”I trust that the law will restrict any data leaks to third parties”.
+[…]
+Trust in the recipient entity..Trust in the recipient entity.Trust in the recipient entity..
 13.9% and 17.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed a general trust in the recipient entity without specifying reasons.
-For instance, P70 succinctly said "I feel they are reliable and trustworthy" and P50 stated that "I assume they take their research seriously, so they would handle the data carefully".
-Other respondents, like P145, reported having no reason not to trust it: "I have no reason to think they would do anything nefarious with my medical data".
-Trust in the auditing process.
+For instance, P70 succinctly said ”I feel they are reliable and trustworthy” and P50 stated that ”I assume they take their research seriously, so they would handle the data carefully”.
+Other respondents, like P145, reported having no reason not to trust it: ”I have no reason to think they would do anything nefarious with my medical data”.
+[…]
+Trust in the auditing process..Trust in the auditing process.Trust in the auditing process..
 7.8% and 7.9% of respondents in the for-profit and non-profit scenarios, respectively, found the auditing statement to add at least some reliability.
-For example, P202 showed some reservation but found confidence in the public nature of audits: "I don’t fully trust them, but I somewhat do, particularly if audits and verification of results are made public.
-That said, claiming that the advisory board is external is only partly reassuring, as it’s appointed by the institute." Additionally, P540 had trust in the entity’s data handling because "they let outsiders audit them" and P51 (who received the expert-auditing statement) because of "safelocks and checks in place".
-On the other hand, one interviewee doubted the expert auditing process and stated: "somebody else says something doesn’t mean that it’s real.".
-No questionnaire respondent explicitly raised concern towards the auditing process.
+For example, P202 showed some reservation but found confidence in the public nature of audits: ”I don’t fully trust them, but I somewhat do, particularly if audits and verification of results are made public.
+That said, claiming that the advisory board is external is only partly reassuring, as it’s appointed by the institute”.
+Additionally, P540 had trust in the entity’s data handling because ”they let outsiders audit them” and P51 (who received the expert-auditing statement) because of ”safelocks and checks in place”.
+On the other hand, one interviewee doubted the expert auditing process and stated: ”somebody else says something doesn’t mean that it’s real”.
+No survey respondent explicitly raised concern towards the auditing process.
 Across for-profit and non-profit scenarios, 31% and 15% of respondents, on average, expressed negative privacy expectations because they were skeptical of the privacy statement or doubted whether the recipient entity would actually employ PETs as stated.
 Their qualitative responses offer insights on the underlying reasons:
-Skepticism on the privacy statement and limits of privacy-preserving guarantees.
+[…]
+Skepticism on the privacy statement and limits of privacy-preserving guarantees..Skepticism on the privacy statement and limits of privacy-preserving guarantees.Skepticism on the privacy statement and limits of privacy-preserving guarantees..
 21.2% and 10.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed general distrust in the feasibility of the privacy statement.
-In the non-profit scenario, P34 stated that "no privacy technology is foolproof".
-Similarly, in the for-profit scenario, P45 wrote: "I don’t trust that the privacy-preserving tech would work." More precisely, privacy statements were considered too ideal to be fully enforced.
+In the non-profit scenario, P34 stated that ”no privacy technology is foolproof”.
+Similarly, in the for-profit scenario, P45 wrote: ”I don’t trust that the privacy-preserving tech would work.” More precisely, privacy statements were considered too ideal to be fully enforced.
 Some respondents believed that unauthorized employee access would be unavoidable.
-For instance, P71 said: "Why would I trust someone other than my doctor with my medical records?
+For instance, P71 said: ”Why would I trust someone other than my doctor with my medical records?
 These days especially, I don’t trust anyone.
 There could be a breach or simply people I don’t know from a hole in the wall will then have access to all my med records.
-Insane!".
-Others felt that data breaches and cyberattacks were inevitable.
-For instance, P83 stated that "the primary reason for my distrust is due to past news of companies being hacked by people and their data getting leaked", while P124 noted that "corporate data breaches are very common", and P80 echoed this sentiment, saying: "I believe the [intentions] will be good, but data can be hacked".
+Insane!”.
+Some others felt that data breaches and cyberattacks were inevitable.
+For instance, P83 stated that ”the primary reason for my distrust is due to past news of companies being hacked by people and their data getting leaked”, while P124 noted that ”corporate data breaches are very common”, and P80 echoed this sentiment, saying: ”I believe the [intentions] will be good, but data can be hacked”.
 We observe that the auditing statement did not substantially instill trust on all accounts — with the exception of expert auditing for access control in non-profit scenario, and data expiration in for-profit scenario.
-For example, P474 mentioned: "They can check their privacy technology all they want but when there is a breach it is done and info is stolen.
-After it fails then they say sorry and offer monitoring but the info is still stolen".
-Doubt on the recipient entity’s motivation to employ PETs.
+For example, P474 mentioned: ”They can check their privacy technology all they want but when there is a breach it is done and info is stolen.
+After it fails then they say sorry and offer monitoring but the info is still stolen”.
+[…]
+Doubt on the recipient entity’s motivation to employ PETs..Doubt on the recipient entity’s motivation to employ PETs.Doubt on the recipient entity’s motivation to employ PETs..
 8.6% and 3.3% of the for- and non-profit scenarios, respectively, claimed that these entities are inherently self-serving and lack the motivation to uphold privacy-preserving guarantees or implement such measures at all.
 Most criticism was directed at for-profit entities.
-For instance, P55 noted that for-profit entities "will do what is profitable and not much more than that".
-Similarly, P119 stated "(…) because it is a for-profit organization.
-I expect them to cut corners" and P32 wrote that "for-profit organization have low standard of morality".
+For instance, P55 noted that for-profit entities ”will do what is profitable and not much more than that”.
+Similarly, P119 stated ”(…) because it is a for-profit organization.
+I expect them to cut corners” and P32 wrote that ”for-profit organization have low standard of morality”.
 However, some respondents also expressed concerns about non-profit entities.
-For example, P190 in the non-profit scenario remarked: "Medicine has become a business.
+For example, P190 in the non-profit scenario remarked: ”Medicine has become a business.
 My data is only useful to them if it helps them make more money.
-Money comes first before the actual well being of humans".
-4.2 RQ2: Willingness to donate health data
-Next, we answer the second research question (RQ2, see section 1): “How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?".
+Money comes first before the actual well being of humans”.
+4.2. RQ2: Willingness to donate health data
+Next, we answer the second research question (RQ2, see § 1): “How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?”.
 As in the previous subsection, we statistically analyze the responses and then report the qualitative rationales collected from the open-text questions.
-Statement Shown Overall For-Profit Non-Profit
-PG(1) - Anonymization 0.68 0.63 0.73
-+ AG(1) - Expert audit 0.63 0.59 0.68
-+ AG(2) - Self audit 0.70 0.67 0.74
-PG(2) - Access control 0.68 0.67 0.68
-+ AG(1) - Expert audit 0.69 0.59 0.78
-+ AG(2) - Self audit 0.58 0.61 0.56
-PG(3) - Data expiration 0.73 0.67 0.80
-+ AG(1) - Expert audit 0.69 0.67 0.72
-+ AG(2) - Self audit 0.68 0.70 0.65
-PG(4) - Purpose restriction 0.72 0.53 0.89
-+ AG(1) - Expert audit 0.71 0.71 0.70
-+ AG(2) - Self audit 0.72 0.72 0.72
-Control 0.65 0.60 0.68
-TABLE V: Percentage of respondents willing to donate their personal health data to the recipient entity in each scenario.(e.g., the left-most and top-most numerical cell indicates that 68% of participants in both entity scenarios that with PG(1) presented but no auditing statement were willing to donate.)
-[…]
-Analysis Set Up.Analysis Set UpAnalysis Set Up.
-Table V summarizes the respondents’ willingness to donate their health data to a recipient entity.
-The 1st column (“overall") reports results for all entity types, the 2nd column reports the results of the subset with for-profit entity, and the 3rd column reports the results of the subset with non-profit entity.
+4.2.1. Analysis Set Up
+In Figure 9, we summarize the respondents’ willingness to donate their health data to a recipient entity.
+The 1st column (“overall”) reports results for all entity types, the 2nd column reports the results of the subset with for-profit entity, and the 3rd column reports the results of the subset with non-profit entity.
 The values represent the percentage of respondents willing to donate their heath data to the recipient entity.
@@
 We first analyze and confirm the influence of the recipient entity and then separate the participants based on the recipient entities and analyze them in separate models.
-The results are summarized in Table VI.
+The results are summarized in Table 3.
 In Figure 10, we visualize the overlap in experimental group respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described.
+4.2.2. Response Distribution and Statistical Analysis Results
+Figure 9.
+Percentage of respondents willing to donate their personal health data to the recipient entity in each scenario (e.g., the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios that with PG(1) presented but no auditing statement were willing to donate.) The agreements are binarized here to present the overall tendency.
+Figure 10.
+Overlap in respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described (respondents in the control – no privacy statement – are excluded).
+The numbers outside each circle summarize the total number of respondents who were e.g., willing to donate their data (310).
+$N=454$.
 Entity Factor Levels OR 95% CI $p$-value
-Both Entity - FP 0.66 [0.45, 0.98] 0.038
+Both Entity - FP 0.38 [0.24, 0.59] <0.001
 For-Profit Privacy Statement
-   PG(1) 1.14 [0.39, 3.24] 0.809
-   PG(2) 1.24 [0.42, 3.58] 0.696
-   PG(3) 1.54 [0.52, 4.47] 0.432
-   PG(4) 1.35 [0.45, 3.88] 0.584
+   PG(1) 0.65 [0.25, 1.70] 0.381
+   PG(2) 0.68 [0.27, 1.75] 0.429
+   PG(3) 0.86 [0.32, 2.28] 0.756
+   PG(4) 0.62 [0.22, 1.70] 0.352
 Privacy Expectation
-   PG(1) 3.07 [1.21, 8.06] 0.020
-   PG(2) 2.64 [1.00, 7.33] 0.054
-   PG(3) 6.06 [2.15, 18.27] <0.001
-   PG(4) 6.33 [2.34, 18.49] <0.001
-   AG(1) 5.34 [2.71, 10.82] <0.001
-   AG(2) 2.94 [1.46, 6.24] 0.003
+   PG(1) 1.09 [0.66, 1.80] 0.741
+   PG(2) 1.60 [0.96, 2.69] 0.073
+   PG(3) 1.70 [1.01, 2.88] 0.047
+   PG(4) 3.25 [1.78, 5.87] <0.001
+   AG(1) 2.46 [1.40, 4.27] 0.002
+   AG(2) 1.32 [0.76, 2.27] 0.317
 Demographics & Experiences
-   Education 0.75 [0.33, 1.63] 0.466
-   Age 0.68 [0.33, 1.39] 0.290
-   Gender 0.80 [0.40, 1.61] 0.535
-   Tech Background 1.33 [0.57, 3.15] 0.512
-   Egocentrism 0.41 [0.20, 0.84] 0.016
-   Donation History 9.47 [2.22, 54.68] 0.005
+   Education 0.73 [0.43, 1.26] 0.264
+   Age 1.00 [0.98, 1.01] 0.765
+   Gender 1.30 [0.80, 2.12] 0.304
+   Tech Background 1.57 [0.88, 2.80] 0.127
+   Egocentricity 0.69 [0.42, 1.14] 0.144
+   Donation History 3.25 [1.31, 8.01] 0.011
 Non-Profit Privacy Statement
-   PG(1) 1.17 [0.36, 3.52] 0.783
-   PG(2) 0.88 [0.27, 2.60] 0.823
-   PG(3) 1.16 [0.35, 3.55] 0.794
-   PG(4) 1.45 [0.44, 4.45] 0.522
+   PG(1) 1.27 [0.52, 3.13] 0.603
+   PG(2) 0.94 [0.37, 2.38] 0.901
+   PG(3) 1.17 [0.45, 3.00] 0.746
+   PG(4) 1.54 [0.61, 3.90] 0.369
 Privacy Expectation
-   PG(1) 7.32 [2.48, 24.13] <0.001
-   PG(2) 2.03 [0.76, 5.64] 0.163
-   PG(3) 3.13 [0.78, 12.76] 0.102
-   PG(4) 12.15 [3.31, 52.60] <0.001
-   AG(1) 3.19 [1.49, 6.82] 0.003
-   AG(2) 3.74 [1.72, 8.89] 0.001
+   PG(1) 1.72 [1.02, 2.89] 0.043
+   PG(2) 1.46 [0.87, 2.46] 0.152
+   PG(3) 1.27 [0.73, 2.25] 0.400
+   PG(4) 5.21 [2.61, 10.38] <0.001
+   AG(1) 2.59 [1.35, 4.90] 0.004
+   AG(2) 1.17 [0.68, 2.03] 0.563
 Demographics & Experiences
-   Education 0.37 [0.15, 0.86] 0.025
-   Age 1.28 [0.61, 2.70] 0.516
-   Gender 0.83 [0.38, 1.78] 0.630
-   Tech Background 0.39 [0.16, 0.89] 0.026
-   Egocentrism 1.79 [0.85, 3.96] 0.136
-   Donation History 4.11 [1.23, 17.58] 0.034
-TABLE VI: Influences of privacy expectations and demographics on willingness to donate of the three participant groups. OR: odd ratio. CI: confidence interval.
-Figure 6: Overlap in respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described (respondents in the control – no privacy statement – are excluded).
-The numbers outside each circle summarize the total number of respondents who were e.g., willing to donate their data (305).
-$N=456$.
-[…]
-Response Distribution and Statistical Analysis Results.Response Distribution and Statistical Analysis ResultsResponse Distribution and Statistical Analysis Results.
+   Education 0.46 [0.26, 0.80] 0.007
+   Age 1.00 [0.98, 1.02] 0.897
+   Gender 1.16 [0.69, 1.97] 0.560
+   Tech Background 0.71 [0.40, 1.26] 0.237
+   Egocentricity 1.36 [0.83, 2.23] 0.217
+   Donation History 2.25 [1.06, 4.75] 0.033
+Table 3. Influences of privacy expectations and demographics on willingness to donate of the three participant groups. OR: odd ratio. CI: confidence interval.
 There was no direct relationship between showing a privacy statement and respondents’ willingness to donate.
-However, several privacy expectations positively correlate with willingness to donate for both non-profit and for-profit entities, except in specific cases like access control for both for- and non-profit and data expiration for the non-profit.Visualized in Figure 6, we observe high alignment among the three explored constructs, where 234 (51.3%) respondents gave all positive responses and 70 (15.4%) respondents gave all negative responses.
+However, several privacy expectations positively correlate with willingness to donate for both non-profit and for-profit entities, except in specific cases like access control for both for- and non-profit and data expiration for the non-profit.Visualized in Figure 10, we observe high alignment among the three explored constructs, where 236 (52.0%) respondents gave all positive responses and 63 (13.9%) respondents gave all negative responses.
 Combined with the results of RQ(1), the findings suggest that willingness to donate is influenced by privacy expectations, which are, in turn, impacted by privacy statements.
 Privacy statements and expectations are not the only reason for donation intentions however.
-While participants were generally less willing to donate to for-profit entities (OR=0.66, $p=0.038$), when analyzing only control group participants, we observe that the absolute difference in willingness to donate between for-profit and non-profit entities was 8.8% was smaller than the 20.3% difference in privacy expectations between the two entities.
+While participants were generally less willing to donate to for-profit entities (OR=0.38, $p=\textless 0.001$), when analyzing only control group participants, we observe that the absolute difference in willingness to donate between for-profit and non-profit entities was 8.8% was smaller than the 20.3% difference in privacy expectations between the two entities.
 This indicates that other non-privacy-related factors, such as the perceived impact or benefits of the donations, also influence participants’ willingness to donate.
-Furthermore, we observe that respondents with egocentric relationships to the donation subject are less willing to donate to a for-profit entity (OR = 0.41, $p=0.016$).
-This is potentially related to their existing (negative) experiences with for-profit entities in healthcare.
+Furthermore, we observe that respondents with bachelor’s degrees are less willing to donate to a non-profit entity (OR = 0.46, $p=0.007$), which may be related to their greater awareness of the complexities and potential vulnerabilities with privacy statement.
+They have heightened privacy concerns that need to be addressed.
+For example, P42 wanted to verify that the PET worked visually; P97 thought that only the stored data was encrypted, but the data transaction was not protected the same way; and P236 argued that: “privacy-preserving technology that works today may not work in a few years.”
+Finally, we found that prior donors are more willing to donate to both entities.
+Respondents who have donated before are more willing to donate to both entities, with the OR value of 3.25 ($p=0.011$) and 2.25 ($p=0.033$) for the for and non-profit entities, respectively.These respondents expressed strong willingness of supporting science and recognized the importance of medical data in developing new treatments.
+For example, P40 claimed: “I feel that it is important information to share in hopes that they can find better ways to deal with diabetes.” and P208 responded: “Breakthroughs in science and moving forward in knowledge, our greatly benefited by such Data Collection”.
+4.2.3. Qualitative Analysis of Donation Intention Via Open-Answer Responses
+Many respondents reasons for donating (or not) included a desire to support research, ego-centric connections, or desire for personal reward.
+In the for-profit scenario, 50.2% of respondents’ reasons for donating (or not) fell into these categories, in the non-profit scenario, 53.3% did.
+One non-privacy-related reason for donating included supporting research: e.g., P107 expressed that ”there are a number of things I could share to research such as money, time, data, and more.
+I would be willing to share my medical records to contribute to cancer research in hopes that people will be healthier and to enhance research future medicines.” Similarly, P97 stated: ”even if they are for-profit (which I don’t like), any closer we get to developing better treatments and/or a potential cure to cancer is something I’d be willing to assist in.”
+Ego-centric reasons, particularly familial connections, also motivated donations.
+P154 mentioned donating because ”my dad died when his heart failed on him, and I’d have rather that not had happened”, and P233 noted that they ”would be willing to help to prevent someone else’s mom from dying of a stroke like [theirs] did.”
+However, their existing (negative) experiences with for-profit healthcare entities also deter some of them from making donations.
 For example, P136 claimed they have heard a number of times that the entity says they won’t sell your information but can get around it by calling it something else, thus technically selling the information.
 P244 also mentioned that their information has been leaked before by a clinician.
+Some respondents desired personal reward for donating and thus were not interested in donating in our scenario: e.g., P92 explained that ”I would give them my data if I got paid for it.
+I would think twice if they want me to donate the data and make no money from it while the researchers will make money off the of research they compiled.”
 An alternative explanation could be the perceived imbalance in cost and profit between the donor and the recipient entity, where the donor provides data for free while the recipient entity profits from its use.
@@
 However, we note that some individual participants alternatively were motivated to donate due to egocentricity, even to for-profit entities.
-For example, P45 "just lost my father to heart failure thirteen days ago.
+For example, P45 ”just lost my father to heart failure thirteen days ago.
 (…) I would be more than willing to donate my medical records if it helps develop medicines to make hearts function better.
-Respondents with bachelor’s degrees are less willing to donate to a non-profit entity (OR = 0.37, $p=0.025$), which may be related to their greater awareness of the complexities and potential vulnerabilities with privacy statement.
-They have heightened privacy concerns that need to be addressed.
-For example, P42 wanted to verify that the PET worked visually; P97 thought that only the stored data was encrypted, but the data transaction was not protected the same way; and P236 argued that: “privacy-preserving technology that works today may not work in a few years.”
-Prior donors are more willing to donate to both entities.
-Respondents who have donated before are more willing to donate to both entities, with the OR value of 9.47 ($p=0.005$) and 4.11 ($p=0.034$) for the for and non-profit entities, respectively.These respondents expressed strong willingness of supporting science and recognized the importance of medical data in developing new treatments.
-For example, P40 claimed: “I feel that it is important information to share in hopes that they can find better ways to deal with diabetes." and P208 responded: “Breakthroughs in science and moving forward in knowledge, our greatly benefited by such Data Collection."
-[…]
-Qualitative Analysis of Donation Intention Via Open-Answer Responses.Qualitative Analysis of Donation Intention Via Open-Answer ResponsesQualitative Analysis of Donation Intention Via Open-Answer Responses.
-Many respondents reasons for donating (or not) included a desire to support research, ego-centric connections, or desire for personal reward.
-In the for-profit scenario, 50.2% of respondents’ reasons for donating (or not) fell into these categories, in the non-profit scenario, 53.3% did.
-One non-privacy-related reason for donating included supporting research: e.g., P107 expressed that "there are a number of things I could share to research such as money, time, data, and more.
-I would be willing to share my medical records to contribute to cancer research in hopes that people will be healthier and to enhance research future medicines".
-Similarly, P97 stated: "even if they are for-profit (which I don’t like), any closer we get to developing better treatments and/or a potential cure to cancer is something I’d be willing to assist in".
-Ego-centric reasons, particularly familial connections, also motivated donations.
-P154 mentioned donating because "my dad died when his heart failed on him, and I’d have rather that not had happened", and P233 noted that they "would be willing to help to prevent someone else’s mom from dying of a stroke like [theirs] did".
-Some respondents desired personal reward for donating and thus were not interested in donating in our scenario: e.g., P92 explained that "I would give them my data if I got paid for it.
-I would think twice if they want me to donate the data and make no money from it while the researchers will make money off the of research they compiled".
 Respondents expressed negative sentiments toward for-profit entities collecting data for research.
@@
 In the for-profit scenario, 11.4% explicitly refused due to negative perceptions of such organizations.
-For instance, P191 and P202 respectively stated: "I have a negative connotation with for profit organization" and "I feel like this type of organization already profits significantly off a large number of people (occasionally off of me as well) and as such, I do not want to give them direct permission to profit further off of me." Furthermore, respondents felt that for-profits already benefit enough and were unwilling to contribute further.
-For instance, P36 wrote: "For profit organization makes profit using the data.
+For instance, P191 and P202 respectively stated: ”I have a negative connotation with for profit organization” and ”I feel like this type of organization already profits significantly off a large number of people (occasionally off of me as well) and as such, I do not want to give them direct permission to profit further off of me.” Furthermore, respondents felt that for-profits already benefit enough and were unwilling to contribute further.
+For instance, P36 wrote: ”For profit organization makes profit using the data.
 I’m never going to donate it.
-They can BUY it from me for a reasonable amount."
+They can BUY it from me for a reasonable amount.”
 Alternatively, a small group of 1.3% respondents explicitly mentioned that they were willing to donate data to entities of non-profit nature, but even they expressed concerns about potential misuse or unethical behavior within these organizations.
-For instance, P122 mentions: "it’s always going to be in the back of my head that there’s a possibility that it’s being sold or taken or used in some way I’m not okay with.
-Scares me a bit".
-Then, P282 notes: "it is difficult to police everyone in a non-profit organization.
-I have seen leaders in organizations act unethically at times, which makes me think that even a well-intended organization cannot fully control the actions of every employee or volunteer they have."
+For instance, P122 mentions: ”it’s always going to be in the back of my head that there’s a possibility that it’s being sold or taken or used in some way I’m not okay with.
+Scares me a bit.” Then, P282 notes: ”it is difficult to police everyone in a non-profit organization.
+I have seen leaders in organizations act unethically at times, which makes me think that even a well-intended organization cannot fully control the actions of every employee or volunteer they have.”
 Privacy-related donation considerations focused on data sensitivity and leakage.
-Concerns for not donating were mostly about data leakage, with respondents expressing that they 12.9% and 13.9% participants in the for-profit and non-profit scenario, respectively, were protective of their data and hence would not donate at all, as P545 noted that "there are so many data leaks from so called safe places.
-No one can anticipate what hackers can do in the future." Others cited previous data leak experiences as to why they would not donate, as P290 said "my medical data has been breached in the past by a clinician and I had to file a formal grievance against the health care system that employed her.
-For this reason, because my medical history was abused, I no longer have any trust and will never voluntarily consent to my medical history/information being shared with others." Some respondents who would not donate data also expressed that they would only possibly donate to specific entities that have already proven trustworthiness to them, as P357 explained that "regardless of privacy policy, the probability of a leak or misuse is high.
-unless it is an organization I have had personal interaction with, or am very familiar with, it is unlikely I would donate my medical records.".
+Concerns for not donating were mostly about data leakage, with respondents expressing that they 12.9% and 13.9% participants in the for-profit and non-profit scenario, respectively, were protective of their data and hence would not donate at all, as P545 noted that ”there are so many data leaks from so called safe places.
+No one can anticipate what hackers can do in the future.” Others cited previous data leak experiences as to why they would not donate, as P290 said ”my medical data has been breached in the past by a clinician and I had to file a formal grievance against the health care system that employed her.
+For this reason, because my medical history was abused, I no longer have any trust and will never voluntarily consent to my medical history/information being shared with others.” Some respondents who would not donate data also expressed that they would only possibly donate to specific entities that have already proven trustworthiness to them, as P357 explained that ”regardless of privacy policy, the probability of a leak or misuse is high.
+unless it is an organization I have had personal interaction with, or am very familiar with, it is unlikely I would donate my medical records.”.
 Another reason was that the participant was concerned that their medical data included too much or too specific information.
-P155 highlighted this concern: "I would not feel comfortable with so much of my personal health information being shared with an organization that is not involved with my direct medical care, regardless of whether there is an advisory board or not"; as did P229: "It’s a lot of detailed information that I’m worried if it somehow gets in the wrong hands, it could reveal a lot of private information about me." Contrastingly, some respondents felt comfortable sharing data due to a lack of sensitive information (e.g., "I don’t have mayor illnesses or nothing to hide so Im ok with that") or because they already share information with other entities (e.g., "Considering I already disclose this information to other organizations (for example, data collection on phone, data compiled through search history), other agencies likely have the information on hand already and so another organization having it is no different").
-5 Discussion
-Analyzing the effect of privacy statements on users’ perceptions, trust, and willingness to donate in the context of a survey is particularly useful because the survey mimics the actual process of consenting that recipient entities use when requesting data.
-These entities typically provide an introduction, present privacy guarantees, and ask for consent to donate data, much like our survey.Our study reveals several key findings about privacy expectations, their influence on willingness to donate health data, and the implications for both researchers and regulators.
+P155 highlighted this concern: ”I would not feel comfortable with so much of my personal health information being shared with an organization that is not involved with my direct medical care, regardless of whether there is an advisory board or not”; as did P229: ”It’s a lot of detailed information that I’m worried if it somehow gets in the wrong hands, it could reveal a lot of private information about me.” Contrastingly, some respondents felt comfortable sharing data due to a lack of sensitive information (e.g., ”I don’t have mayor illnesses or nothing to hide so Im ok with that”) or because they already share information with other entities (e.g., ”Considering I already disclose this information to other organizations (for example, data collection on phone, data compiled through search history), other agencies likely have the information on hand already and so another organization having it is no different”).
+5. Concluding Discussion
+Like our survey, when soliciting data donations, recipient entities usually provide an introduction motivating their need to collect the individual’s data, outline privacy guarantees, and seek consent for data donation.
+Privacy statements function similarly to advertising (O’Flaherty, 2022) by shaping and influencing consumers’ expectations of privacy (Schaub et al., 2015).
+Yet, these statements do not always have impact.
+Prior work has highlighted the ways in which such statements can fail due to being vague or inaccurate.
+Our work highlights how the impact of privacy statements is moderated by people’s pre-existing privacy expectations.
+Our participants reported strong pre-existing privacy expectations, even in the absence of explicit privacy statements.
+We observe that these expectations vary strongly based on the profit model of the entity.
+For for-profit entities, where baseline privacy expectations are lower, strong privacy statements can effectively elevate these expectations among users.
+In contrast, the impact on non-profit organizations is minimal because users already hold high privacy expectations for these entities.
 […]
-Pre-existing Privacy Expectations.Pre-existing Privacy ExpectationsPre-existing Privacy Expectations.
-We find that people have strong pre-existing privacy expectations even when no specific privacy statements are shown.
-This expectation is particularly high for non-profit entities, which are perceived as more ethical, thus prioritizing public good over profit.
-In contrast, for-profit entities are seen as driven by financial gain, leading to skepticism on their commitment to protecting user privacy.
-Non-profit entities foster greater trust and confidence in their data handling practices to the point that respondents often expect more stringent privacy measures, such as purpose use restriction.
-However, such measures may not be technically enforced or commonly practiced; indeed, many health organizations employ broad (not purpose restricted) consent [58].
-This gap between expectations and reality can lead to privacy expectation violations, which are a significant source of privacy concerns according to the theory of contextual integrity [77].
+The Halo Effect.The Halo EffectThe Halo Effect.
+Participants qualitative responses suggest that their strongly positive privacy expectations for non-profit entities are rooted in arguments of morality: non-profits are perceived as more ethical and dedicated to serving the public good because they are not pursuing profit.
+This findings are consistent with a phenomenon termed “the halo effect”, which describes how an overall positive impression of a particular entity (e.g., the good work that a non-profit does) can lead to ungrounded assumptions about how the organization operates (Nisbett and Wilson, 1977; de Bruin Cardoso et al., 2024).
+While prior work in the business literature focuses on how the halo effect influences expectations about business practices such as donation management, our work illustrates that the halo effect also influences people’s expectations about the way non-profits handle data privacy.
+These positive expectations do not always align with reality.
+Privacy safeguards are not consistently enforced or widely practiced; for example, many health organizations rely on broad consent frameworks rather than adopting stricter purpose-restricted consent measures (Kaye et al., 2015).
+This gap between expectations and reality can result in privacy expectation violations, a major source of privacy concerns as explained by the theory of contextual integrity (Nissenbaum, 2009).
+Indeed, a well-known non-profit mental health helpline shared it’s data with a commercial entity, leading to significant public outcry (Levine, 2022).
+Future work may seek to explore how to design privacy communications that set appropriate expectations in the presence of such effects, as a prior work has sought to do in other domains (e.g., (Burton et al., 2015)).
 […]
-Impact of Privacy Statements.Impact of Privacy StatementsImpact of Privacy Statements.
-Privacy statements play a role similar to advertising, such as those used by companies like Apple [79].
-For for-profit entities, where baseline privacy expectations are lower, privacy statements raise these expectations.
-(In contrast, for non-profits, the impact is minimal because the expectations are already high.) Thus, we argue that privacy expectations of for-profits can be effectively influenced by transparently implementing and communicating about PETs.
-[…]
-Privacy Expectations and Willingness to Donate.Privacy Expectations and Willingness to DonatePrivacy Expectations and Willingness to Donate.
-Privacy expectations significantly influence people’s willingness to donate health data: individuals are more willing to donate when their privacy expectations are met or exceeded, as presented in Figure 6.
-Thus, we argue that entities wishing to receive people’s data should not only focus on employing PETs but also on effectively communicating these measures to potential data donors.
-However, there is a risk of "privacy washing" [67, 26], where information about PETs and “vague claims” are used to raise expectations and mislead individuals into donating their personal data.
+Privacy Washing.Privacy WashingPrivacy Washing.
+Conversely, our participants’ had far lower initial expectations of the privacy practices of for-profit entities due to their perceived singular focus on profit (a “horn” effect).
+As a result, information about data protection guarantees significantly raised their privacy expectations, and in turn, their willingness to donate their data.
+This means that for-profit entities have the opportunity to gain “competitive advantage” (Schaub et al., 2015) through influencing and enhancing user privacy expectations via transparently implementing and effectively communicating about PETs.
+However, this effect also posses a vulnerability to ”privacy washing” (MainWP, 2023; Cirucci, 2024), in which organizations make vague or exaggerated claims about PETs to falsely raise privacy expectations.
+Such deceptive practices can mislead individuals into sharing their data, erode trust, and discourage future data donations if individuals later feel deceived.
 […]
 The Role of Auditing.The Role of AuditingThe Role of Auditing.
-Auditing had a limited effect on raising privacy expectations, except in the case of purpose use restrictions in the for-profit scenario.
-Participants recognize that PETs are not flawless and that auditing does not practically mitigate the risk of data breaches.
-This result highlights a disconnect between the expert view, which often sees auditing as a fundamental guarantee to offer, and the participant view, which is more concerned with the overall protection of data rather than specific PETs implementations (e.g., a respondent writes “they can guarantee privacy all they want, things still get hacked”).
+In theory, auditing could offer verification that combats both the halo effect and privacy washing by grounding assumed or stated privacy guarantees in technical reality.
+Unfortunately, we find that auditing has a limited effect on elevating privacy expectations among participants, except in scenarios involving purpose use restrictions by for-profit entities.
+Participants understand that while PETs and auditing processes are in place, they are not infallible, and auditing does not substantially mitigate the practical risks associated with data breaches or unauthorized access.
+This perception shows a disconnect between the perspectives of experts and participants, as shown in (Dechand et al., 2019; Oates et al., 2018; Wu and Zappala, 2018).
+Experts often regard auditing as a fundamental assurance mechanism that reinforces trust and compliance.
+In contrast, participants are more concerned with the overall protection of their data, regardless of specific technical implementations or compliance processes.
+As one respondent noted, ”They can guarantee privacy all they want; things still get hacked.” True privacy is doubted to exist (Lerner et al., 2022).
 […]
-Recommendations for Researchers.Recommendations for ResearchersRecommendations for Researchers.
-If researchers aim to foster trust in entities that collect personal health data, they may want to focus on effectively communicating about PETs that protect against data breaches.
-Technologies like multi-party computation, although not yet scalable, could enhance trust in data protection: they specifically distribute sensitive information across different entities to ensure that no single entity can reconstruct the information on its own, thereby minimizing the impact of a breach.
+Building Trust.Building TrustBuilding Trust.
+Our respondent’s prior experiences with data leakage often informed this fatalistic perspective on privacy.
+Technologies like multi-party computation (MPC), although not yet scalable, have the potential to minimize the impact of a breach by specifically distributing sensitive information across different entities to ensure that no single entity can reconstruct the information on its own.
 However, deploying such systems requires careful engineering to prevent distributed data breaches on all the entities involved.
-We believe that convincing the public of the effectiveness of these measures presents an even greater challenge, as there is existing doubt and any failure of these systems could severely damage trust.
+Existing skepticism about data security means that any failure of these systems could severely damage trust, potentially more so than if simpler, well-understood methods were compromised.
 […]
-Recommendations for Regulators.Recommendations for RegulatorsRecommendations for Regulators.
-Our findings suggest the necessity of policies that ensure privacy practices align with public expectations.
-Given the strong assumptions people have about privacy protections, especially for non-profit entities, regulations should be designed to ensure that the protections assumed are consistently applied.
+Future Work.Future WorkFuture Work.
+In this work, we explicitly examined the influence of four privacy guarantees and two auditing guarantees.
+Future research could broaden this scope by incorporating a more diverse and comprehensive range of PETs, with a particular focus on understanding how multiple guarantees interact in real-world scenarios where they often coexist.
+Additionally, researchers can also conduct targeted studies to understand the perspectives of diverse demographic groups, including individuals with a strong distrust of data-handling practices.
+Identifying and understanding these individuals is critical in fostering confidence in PETs and practices.
+Our study focused exclusively on individuals’ perspectives regarding formal medical data donation scenarios.
+However, medical data are also frequently collected through smart devices and donated with a single click via in-app prompts.
+Prior work emphasizes the role of modality and timing in the design space of privacy information (Schaub et al., 2015).
+Future research could explore how such factors intersect with biases such as the halo effect to influence privacy expectations, trust, and willingness to share sensitive information.
+Finally, prior work (Seberger and Patil, 2021) studying COVID-19 contract tracing apps – many created through a partnership between for-profits and non-profits – found that people assumed these apps would “operate according to the economics of surveillance capitalism…[making it] unreasonable to expect them to trust the system that profits from the sale of their data.” Future work may seek to further interrogate how such partnerships, which are common in the medical domain, further complicate our understanding of people’s privacy expectations and their sensitivity to statements of data protection.
+Appendix A Survey
+In this section, we report the full text of our questionnaire.
+We first ask for a participant themselve or their close relatives have been diagnosed with a disease.
+In the rest of the questionnaire, we will choose at random one disease the participant has selected, if any, or one disease the participant has not selected.
+This is labeled in Italic as disease.
+The entity, also in Italic, may be either non-profit entity or for-profit entity, chosen at random.
+Egocentricity
+  1. (1)
+    Have you ever been diagnosed with any of these diseases?
+    • Cancer (Breast, Colorectal, Lung, and Prostate)
+    • Diabetes
+    • Heart Failure
+    • Hypertension (High blood pressure)
+    • Stroke
+    • None of the above
+  2. (2)
+    Have your partner and/or close relatives (parents, children, siblings) ever been diagnosed with any of these diseases?
+    • Cancer (Breast, Colorectal, Lung, and Prostate)
+    • Diabetes
+    • Heart Failure
+    • Hypertension (High blood pressure)
+    • Stroke
+    • None of the above
+Survey Introduction
+Next, we present you with a scenario and ask you how you would respond.
+Please do your best to imagine yourself in this scenario.
+There is no right or wrong answer to these questions: we are interested to hear how you would feel about being in this scenario and what you would do.
+Imagine that an entity entity wants to develop a new treatment for disease.
+They need medical data from people with and without disease to develop the treatment.
+They ask you to donate your medical record to help develop the treatment.
+Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
+Questions
+We will now ask you a series of questions about this scenario. For your reference, the scenario text will appear at the top of each page.
+  1. (1)
+    How would you rate your understanding of the above scenario?
+    • Fully understand
+    • Mostly understand
+    • Partially understand
+    • Not understand
+  2. (2)
+    In this scenario how likely would you be to donate your medical record?
+    • Very likely
+    • Likely
+    • Unlikely
+    • Very unlikely
+  3. (3)
+    Why your are willing (or unwilling) to share your medical record with this entity?
+  4. (4)
+    Please rate your agreement with the following statement:
+    I trust the entity will handle my data as described.
+    • Strongly agree
+    • Somewhat agree
+    • Somewhat disagree
+    • Strongly disagree
+  5. (5)
+    Please explain why you do (or do not) trust that the entity will handle your data as described.
+  6. (6)
+    In this scenario, how likely do you think it is that each of the following will occur?
+    • My full name or other personal identifiable information will be linked to the donated medical record.
+      • Very likely
+      • Likely
+      • Unlikely
+      • Very unlikely
+    • The donated medical record will be deleted at a set point in time.
+      • Very likely
+      • Likely
+      • Unlikely
+      • Very unlikely
+    • Any employee at the recipient entity will be able to access the donated medical records.
+      • Very likely
+      • Likely
+      • Unlikely
+      • Very unlikely
+    • The donated medical records will be used for another purpose without my consent.
+      • Very likely
+      • Likely
+      • Unlikely
+      • Very unlikely
+    • A group of independent experts will verify whether the privacy-preserving technology works and publish a report on their findings.
+      • Very likely
+      • Likely
+      • Unlikely
+      • Very unlikely
+    • I will be able to hire someone to verify that my medical record is protected as described.
+      • Very likely
+      • Likely
+      • Unlikely
+      • Very unlikely
+    • If I donate my data, I will meet Albert Einstein.
+      • Very likely
+      • Likely
+      • Unlikely
+      • Very unlikely
+  7. (7)
+    History In real life, have you ever donated your medical record?
+    • Yes, I have donated my medical record for disease
+    • Yes, I have donated my medical record for research of other illnesses.
+    • Yes, I have donated my medical record for non-medical purposes.
+    • No, I was asked to donate my medical record but I declined to participate.
+    • No, I have never been asked to donate my medical record.
+    • I prefer not to answer.
+Internet Skills
+  1. (1)
+    How familiar are you with the following terminology? Please choose a number between 1 and 5 where 1 represents “no understanding” and 5 represents “full understanding” of the term.
+    • Advanced search
+    • PDF
+    • Spyware
+    • Wiki
+    • Cache
+    • Phishing
+  2. (2)
+    Which of the following best describes your educational background or job field?
+    • I have an education in, or work in, the field of computer science, computer engineering or IT
+    • I DO NOT have an education in, nor do I work in, the field of computer science, computer engineering or IT.
+    • Prefer not to answer.
+  3. (3)
+    Have you ever heard any of the following techniques? (select all that apply)
+    • Differential Privacy
+    • End-to-end encryption
+    • Public-key cryptography
+    • Trusted execution environment (TEE)
+    • Pseudo-party computation
+    • None of the above
+Demographics
+  1. (1)
+    What year were you born in? (4 digits)
+  2. (2)
+    To which gender do you most identify? (Select all that may apply)
+    • Woman
+    • Man
+    • Non-binary
+    • Prefer not to answer
+    • Prefer to self-describe
+  3. (3)
+    What is the highest level of education you have completed?
+    • Less than high school degree (1)
+    • High school graduate (high school diploma or equivalent including GED) (2)
+    • Some college but no degree (3)
+    • Associate’s degree (4)
+    • Bachelor’s degree (5)
+    • Advanced degree (e.g., Master’s, doctorate) (6)
+    • Prefer not to answer (7)
+Appendix B Code Book
+[…]
+Data Donation.Data DonationData Donation.
+  1. (1)
+    Irrelevant response
+  2. (2)
+    Opt to donate regardless
+  3. (3)
+    Opt not to donate regardless
+  4. (4)
+    Require compensation
+  5. (5)
+    Only donate to a non-profit entity
+  6. (6)
+    No donation to a for-profit entity
+  7. (7)
+    Need more information
+  8. (8)
+    The provided protection is persuasive
+  9. (9)
+    Generally, the donation scenario looks good
+  10. (10)
+    Not satisfied with the provided protection
+[…]
+Trustworthy.TrustworthyTrustworthy.
+  1. (1)
+    Trust without a specific reason
+  2. (2)
+    Irrelevant response
+  3. (3)
+    Perfect protection does not exist
+  4. (4)
+    The recipient entity has no reason not to protect the data
+  5. (5)
+    The recipient entity has no reason to protect the data
+  6. (6)
+    A non-profit entity is trustworthy
+  7. (7)
+    A for-profit entity is not trustworthy
+  8. (8)
+    Need more information
+  9. (9)
+    The provided protection is persuasive
+  10. (10)
+    Generally, the donation scenario looks good
+  11. (11)
+    Not satisfied with the provided statement

### v3

@@
 The Role of Privacy Guarantees in Voluntary Donation of Private Health Data for Altruistic Goals
-Abstract.
-The voluntary donation of private health information for altruistic purposes, such as supporting research advancements, is a common practice.
-However, concerns about data misuse and leakage may deter people from donating their information.
-Privacy Enhancement Technologies (PETs) aim to alleviate these concerns and in turn allow for safe and private data sharing.
-This study conducts a vignette survey ($N=494$) with participants recruited from Prolific to examine the willingness of US-based people to donate medical data for developing new treatments under four general guarantees offered across PETs: data expiration, anonymization, purpose restriction, and access control.
-The study explores two mechanisms for verifying these guarantees: self-auditing and expert auditing, and controls for the impact of confounds including demographics and two types of data collectors: for-profit and non-profit institutions.
+Abstract
+While voluntary donation of private health information enables valuable research, privacy concerns often deter potential donors.
+Privacy Enhancing Technologies (PETs) aim to address these concerns, yet their effectiveness in encouraging data sharing remains unclear.
+This study conducts a vignette survey ($N=494$) with participants recruited from Prolific to examine the willingness of US-based people to donate medical data for developing new treatments.
+It investigates four general guarantees offered across PETs: data expiration, anonymization, purpose restriction, and access control and two mechanisms for verifying these guarantees: self-auditing and expert auditing.
+This study also controls for the impact of confounds, including demographics and two types of data collectors: for-profit and non-profit institutions.
 Our findings reveal that respondents hold such high expectations of privacy from non-profit entities a priori that explicitly outlining privacy protections has little impact on their overall perceptions.
@@
 We emphasize the risks associated with these findings and underscore the critical need for future interdisciplinary research efforts to bridge the gap between the technical community’s and end-users’ perceptions regarding the effectiveness of auditing PETs.
-Privacy Guarantee, Health Data Donation, Privacy Expectation, Auditing
-ccs: Security and privacy Social aspects of security and privacy
-1. Introduction
+I Introduction
 The altruistic use of personal health data holds immense potential for societal benefit.
-For example, the public health sector can leverage individuals’ medical histories and health data to analyze epidemiological trends, enhance disease surveillance and risk prediction, diagnose rare and emerging diseases, and accelerate the development of new treatments (Bietz et al., 2015; Aitken et al., 2016; Bartlett et al., 2018; Greene et al., 2019; Kim et al., 2015; Trinidad et al., 2020).
-Beyond these applications, there are significant economic advantages.
-The European Commission estimates that innovations enabling greater data sharing and availability could save billions of euros within the EU health sector (Commission, [n.
-d.]).
-As a result, the demand for personal health data continues to grow (Wilbanks and Friend, 2016; Bietz et al., 2015), particularly in light of the recent surge in large language models (LLMs), whose effectiveness heavily depends on the data they are trained on.
-Upon the acceptance of data-sharing agreements, personal health data is routinely collected by healthcare providers (Bartlett et al., 2018; Kim et al., 2015) as well as personal mobile apps and devices (Greene et al., 2019).
-This data is often accessed by third parties through data-sharing agreements that govern data donation practices (Trinidad et al., 2020).
-However, people frequently self-report that privacy concerns deter them from donating their sensitive information (Redmiles, 2020; Voigt et al., 2020; Silber et al., 2022; Greene et al., 2019; Kim et al., 2015).
-For instance, many individuals express discomfort with hospital policies that permit data-sharing agreements with third parties without obtaining explicit consent from the data owners (Lewis, 2019).
-In an effort to foster data collection and analysis while protecting the privacy of data contributors, there has been an increasing focus on deploying privacy-enhancing technologies (PETs) for data storage (Popa et al., 2011; Pappas et al., 2014), data processing (Arasu et al., 2013; Bajaj and Sion, 2011; Baumann et al., 2015; Gupta et al., 2019; Chowdhury et al., 2019; Al-Rubaie and Chang, 2019; Gaballah et al., 2024), and machine learning (ML) training (Wu et al., [n.
-d.]; De Viti et al., 2022).
-For example, CoVault (De Viti et al., 2022) enables multiparty analytics for epidemics-related datasets while preserving data confidentiality, and Anonify (Gaballah et al., 2024) supports decentralized, dual-level anonymity for secure medical data donation.
-However, in order for such PETs to impact the availability of data, the guarantees they offer must impact user privacy concerns significantly enough to alter their willingness to donate their data.
+For example, the public health sector can leverage individuals’ medical histories and health data to analyze epidemiological trends, enhance disease surveillance, improve risk prediction models, diagnose rare and emerging diseases, and accelerate the development of novel treatments [1, 2, 3, 4, 5, 6].
+Beyond direct medical applications, health data sharing offers economic benefits.
+A unified data-sharing infrastructure could eliminate redundant clinical trials and research efforts, thereby reducing healthcare costs [7].
+According to the European Commission, innovations that enhance health data accessibility could generate savings of billions of euros within the European healthcare sector alone [8].
+These demands are further intensified by recent advances in artificial intelligence, particularly machine learning models, whose effectiveness heavily depends on the quality and quantity of data they are trained on.
+Recognizing these converging medical, scientific, and economic imperatives, the National Academy of Medicine (NAM) has advocated for more open and collaborative approaches to health data sharing [9], emphasizing both the ethical obligation and practical necessity of broader data access frameworks.
+However, despite these compelling benefits and institutional support for expanded health data sharing, privacy concerns remain a significant barrier to widespread data donation.
+Personal health data is routinely collected by healthcare providers [3, 5], mobile applications, and wearable devices [4], often flowing to third parties through data-sharing agreements [6] with limited transparency to the individuals who generated it.
+Studies consistently show that privacy concerns are a primary deterrent to individuals contributing their sensitive health information [10, 11, 12, 4, 5], particularly when hospitals permit data sharing with third parties without explicit patient consent [13].
+In an effort to foster data collection and analysis while protecting the privacy of data contributors, there has been an increasing focus on deploying privacy-enhancing technologies (PETs) for data storage [14, 15], data processing [16, 17, 18, 19, 20, 21, 22], and machine learning (ML) training [23, 24].
+Some systems explicitly consider data donation scenarios: Waldo [25] enables privacy-preserving queries over medical time-series data for remote patient monitoring; CoVault [24] supports expressive queries on sensitive data at scale under a strong threat model, and considers a national-scale epidemic analytics scenario; Anonify [22] provides decentralized dual-level anonymity for medical data donation; and Mycelium [26] supports differentially-private queries over large distributed graphs (e.g., disease-spread data) across millions of user devices.
+However, despite these technological advances, implementing PETs alone does not guarantee increased data availability.
+Privacy protections must be expressed to potential donors in comprehensible terms that address their fundamental concerns, or these sophisticated mechanisms remain ineffective at encouraging data sharing.
 In order to evaluate whether the guarantees offered by PETs ultimately impact people’s willingness to donate their data, at least three steps are necessary.
 First, we must be able to effectively explain guarantees (either general guarantees or those of a particular PET) to the people they aim to protect: end-users.
-Prior work has focused on doing so for particular PETs, chiefly differential privacy (DP) and end-to-end encryption (E2EE) (Cummings et al., 2021; Abu-Salma et al., 2018; Akgul et al., 2021; Bai, 2019; De Luca et al., 2016; Dechand et al., 2019; Abu-Salma et al., 2017; Wu and Zappala, 2018; Xiong et al., 2020; Nanayakkara et al., 2023).
+Prior work has focused on doing so for particular PETs, mainly differential privacy (DP) and end-to-end encryption (E2EE) [27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
 Second, using these explanations, we must evaluate whether the guarantees impact people’s privacy concerns (or expectations), and in turn, their willingness to share data.
-Prior work has done such evaluations using self-report studies focused on individual PETs, again primarily DP (Cummings et al., 2021; Bullek et al., 2017; Nanayakkara et al., 2023) or E2EE (Dechand et al., 2019; Bai et al., 2020; Abu-Salma et al., 2018)).
-Third, while rare, we must validate those results in the field to confirm that the effects observed replicate when studying actual vs.
-intended behavior (e.g., (Dooley et al., 2022b; Acquisti et al., 2013)).
-Instead of focusing on a singular PET and it’s accompanying guarantees, we focus on addressing steps one and two for the general security and privacy guarantees that recur across PETs:
-  1. PG(1):
-    Anonymization (Murthy et al., 2019; Shahid et al., 2022): data is not linkable to its owner (as defined in (Chevrier et al., 2019));
-  2. PG(2):
-    Access control (Greene et al., 2019; Hu et al., 2006; Ni et al., 2010; Tourani et al., 2017): data is accessible only by authorized people, specified in the data collection agreement;
-  3. PG(3):
-    Data expiration (Politou et al., 2018; Sarkar et al., 2018; Perlman, 2005): data is discarded or inaccessible after a given expiration time, specified in the data collection agreement;
-  4. PG(4):
-    Purpose restriction (De Viti et al., 2022; Byun et al., 2005): data is only used for the stated collection purpose, specified in the data collection agreement;
-While previous research has explored how the combined presence of guarantees PG(1)-PG(4) affects data-sharing intentions (Aitken et al., 2016; Mori, 2016), our study takes a different approach by focusing on the impact of each guarantee individually.
-Our objective is to identify which specific protections or factors most significantly influence privacy expectations and, in turn, reported data-sharing intentions.
-Additionally, we investigate the role of auditing or verification processes increasingly introduced to produce public auditing results for PETs, specifically:
-  1. AG(1):
-    Expert auditing (Li et al., 2015; Wang et al., 2011): engaging an aggregator-selected external advisory board to audit the system to verify that the PET functions as described in the data collection agreement;
-  2. AG(2):
-    Self auditing (Holt and Seamons, 2005; Chen et al., 2022): granting anyone, including donors or external advisors appointed by them, the ability to perform such audits.
-We examine the role of these PET guarantees and their audits in shaping privacy expectations and influencing data-sharing intentions within a specific context used in prior work (see e.g., (Cummings et al., 2021; Nanayakkara et al., 2023)): medical data donation.
+Prior work has done such evaluations using self-report studies focused on individual PETs, again primarily DP [27, 37, 36] or E2EE [32, 38, 28].
+Third, while rare, we must validate those results in the field to confirm that the effects observed replicate when studying actual versus intended behavior (e.g., [39, 40]).
+In this work, we focus on addressing the first two steps outlined above: effectively explaining privacy guarantees and evaluating their impact on data-sharing intentions.
+We follow real-world medical consent frameworks [41, 42] and privacy regulations [43, 44], and individually investigate four fundamental guarantees across different technologies: anonymization, access control, data expiration, and purpose restriction (PG(1)–PG(4)), along with two verification processes: expert and self auditing (AG(1)–AG(2)).
+This approach helps identify which specific protections most influence privacy expectations and data-sharing decisions.
+We examine the role of these PET guarantees and their audits in shaping privacy expectations and influencing data-sharing intentions within a specific context used in prior work (see e.g., [27, 36]): medical data donation.
 We ask:
   1. RQ(1):
-    How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?
+    How well do people understand and expect what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?
   2. RQ(2):
     How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?
-To address these questions, we conduct a vignette survey ($n=494$) following the best practice methodology established by prior work (Cummings et al., 2021; Abu-Salma et al., 2018; Xiong et al., 2020).
-Each survey respondent is presented with a hypothetical opportunity to donate their health data to help develop a treatment for a specific chronic disease, along with information on how their data will be protected (PG(1)–PG(4) enforced (or not) by AG(1)–AG(2)).
-Furthermore, we isolate the effect of PG(1)–PG(4) and AG(1)–AG(2) from external confounding factors identified by prior work:
-  • The data-collection entity (Trinidad et al., 2020; Hendricks-Sturrup and Lu, 2023): whether the collecting entity is a for-profit or a non-profit organization;
-  • Socio-demographics (Hargittai et al., 2020; Dooley et al., 2022a; Trinidad et al., 2020): age, gender, education, donation history, and technical background.
+To address these questions, we conduct a vignette survey ($n=494$) following the best practice methodology in prior work [27, 28, 35].
+Each survey respondent is presented with a hypothetical opportunity to donate their health data to help develop a treatment for a specific chronic disease, along with how their data will be protected (PG(1)–PG(4) enforced (or not) by AG(1)–AG(2)).
+We control for confounding factors identified in prior work, including data-collection entity [6, 45] and socio-demographics [46, 47, 6].
 We find that even when told nothing about PETs implemented by the entity, participants are 23% more likely, on average, to expect a non-profit to implement PG(1)–PG(4) and AG(1)–AG(2).
@@
 Privacy expectations, in turn, influence the willingness to donate.
-Furthermore, while the technical community has suggested external audits as a mechanism to increase trust in PET implementation, our initial inquiry suggests that more work is needed to align the perceptions of the experts and end users about the effectiveness of such audits.
+Furthermore, while the technical community has suggested external audits as a mechanism to increase trust in PET implementation, our initial inquiry suggests that more work is needed to explain the purpose and effectiveness of such audits to end-users.
 In fact, the effect of audit statements on people’s willingness to donate is limited to a specific scenario involving for-profit entities and auditing to check that purpose restriction (PG(4)) is correctly implemented.
 We argue that it is critical for non-profit entities to rigorously implement data privacy measures , as any future data leak could lead to a significant loss of trust.
-Prior research (Acquisti et al., 2013) underscores that users place greater value on maintaining their expected privacy than on gaining additional privacy they did not initially anticipate.
-In contrast, we highlight the risk of for-profit entities engaging in ”privacy washing” (Cirucci, 2024), where statements about PETs are used to artificially raise privacy expectations to encourage data collection.
+Prior research [40] underscores that users place greater value on maintaining their expected privacy than on gaining additional privacy they did not initially anticipate.
+In contrast, we highlight the risk of for-profit entities engaging in “privacy washing” [48], where statements about PETs are used to artificially raise privacy expectations to encourage data collection.
 At the same time, our findings reveal that respondents are perceptive of the limitations in general PET guarantees, particularly concerning protections against data breaches.
 This underscores the need for future research to explore how to effectively communicate the potential of stronger, emerging PET guarantees and address the skepticism of end-users.
-2. Related Work
-In this section, we examine prior work on people’s willingness to donate personal data, privacy concerns, the role of privacy-enhancing technologies (PETs), and educational or explanatory strategies related to PETs.
+II Related Work
+In this section, we examine prior work on people’s willingness to donate personal data, privacy concerns, the role of PETs, and educational or explanatory strategies for PETs.
 […]
@@
 The analysis of personal health data is crucial for medical research, particularly in diagnosing emerging or rare diseases and developing new treatments.
-Indeed, the COVID-19 pandemic has further intensified the demand for personal health data (Scott et al., 2020).
-Despite the importance of this data, its sensitive nature poses significant obstacles to its collection (Fahey and Hino, 2020).
-Prior work indicates that individuals are generally willing to donate their data for altruistic purposes (Pilgrim and Bohnet-Joschko, 2022; Aitken et al., 2016; Seberger and Patil, 2021), though there is a notable reluctance when it comes to their health data specifically (Middleton et al., 2020; Mori, 2016).
-However, willingness to donate personal health data increases if the donor or their close family members are directly affected by the disease under investigation (Goodman et al., 2017; Bartlett et al., 2018).
+Indeed, the COVID-19 pandemic has further intensified the demand for personal health data [49].
+Despite the importance of this data, the sensitive nature of health data poses significant obstacles to its collection [50].
+Prior work indicates that individuals are generally willing to donate their data for altruistic purposes [51, 2, 52], though there is a notable reluctance when it comes to their health data specifically [53, 54].
+This reluctance diminishes when the donor or their close family members are directly affected by the disease under study [55, 3], suggesting that non-privacy-related factors influence donation decisions.
 […]
 Privacy concerns.Privacy concernsPrivacy concerns.
-Concerns about the privacy and the misuse of donated data are prominent among potential donors (Mori, 2016).
-These concerns include the risk of being identified, discriminated against, or having their personal sensitive data misused or leaked (Weitzman et al., 2012; Aitken et al., 2016; Kim et al., 2015; McCormick and Hopkins, 2021).
-A common source of these concerns is distrust of the receiving entity (Weitzman et al., 2012; Howe et al., 2018; Aitken et al., 2016; Stockdale et al., 2019), often due to fears that the entity might violate consent terms by sharing data with unauthorized third parties (Kacsmar et al., 2022; Middleton et al., 2020).
-This distrust is intensified when participants are unfamiliar with the recipient entity (Weitzman et al., 2012; Howe et al., 2018), especially if it is a governmental or for-profit organization (Aitken et al., 2016; Stockdale et al., 2019; Garrison et al., 2016; Hendricks-Sturrup and Lu, 2023; Khanra, 2023; American Medical Association, 2022).
-Beyond this intentional data misuse, there are concerns about data leakage caused by hackers or unintentional mishandling (Seh et al., 2020).
-People are also reluctant to interact with entities with any history of data breaches, doubting their ability to safeguard sensitive information (Seh et al., 2020).
+Concerns about privacy and the misuse of donated data are prominent among potential donors [54].
+These concerns include the risk of being identified, discriminated against, and having personal sensitive data misused or leaked [56, 2, 5, 57].
+A common source of these concerns is distrust of the receiving entity [56, 58, 2, 59], often due to fears that the entity might share data with unauthorized third parties [60, 53].
+This distrust is intensified when participants are unfamiliar with the recipient entity [56, 58], or if it is a governmental or for-profit organization [2, 59, 61, 45, 62, 63].
+Beyond this intentional data misuse, there are concerns about data leakage caused by hackers or unintentional mishandling [64].
+People are reluctant to interact with entities having any history of data breaches, doubting their ability to safeguard sensitive information [64].
 […]
 Mental models on PETs.Mental models on PETsMental models on PETs.
-PETs aim to address user privacy concerns.
-While various PETs are available for health data donation environments (Jordan et al., 2022), prior work indicates that participants with greater online privacy literacy tend to have more trusting beliefs in PETs (Harborth and Pape, 2020).
-Studies focusing on sensitive personal data, including health data, reveal similar findings.
-For instance, Dechand et al.
-(Dechand et al., 2019) found that most participants do not trust encryption because they do not understand how it works: e.g., many believe that service providers can bypass encryption to gain unauthorized data access.
-Furthermore, Lerner et al.
-(Lerner et al., 2022) noted that some participants doubt the existence of true privacy.
-Non-experts may misunderstand privacy technologies (Wu and Zappala, 2018; Oates et al., 2018; Dechand et al., 2019), e.g., viewing encryption merely as restrictive access control (Wu and Zappala, 2018) or confusing it with data encoding (Abu-Salma et al., 2017).
-Additionally, non-expert end users may struggle to understand the consequences of inadequate privacy protections (Tang et al., 2021; Velykoivanenko et al., 2022; Abu-Salma et al., 2018; Gerber et al., 2019; Pham and Castro, 2019; Vaniea et al., 2014).
+PETs aim to address user privacy concerns by providing technical safeguards for sensitive data.
+While various PETs are available for health data donation contexts [65], their effectiveness depends largely on users’ trust and understanding.
+Prior work indicates that participants with greater online privacy literacy tend to have more trusting attitudes in PETs [66], while those with limited knowledge often remain skeptical.
+Studies have also found that non-experts systematically misunderstand privacy technologies [34, 67, 32].
+For instance, many incorrectly conceptualize encryption as merely a form of access control [34] or confuse it with simple data encoding [33].
+Additionally, non-expert end users may struggle to understand the consequences of inadequate privacy protections [68, 69, 28, 70, 71, 72].
+Lerner et al.
+[73] identified an even more fundamental barrier: some participants express inherent skepticism about the existence of “true” privacy, illustrating how deep-seated misconceptions can fundamentally undermine the reassurance PETs are designed to provide.
 As a result, even when privacy protections are present, people may have risk expectations that are misaligned with reality.
@@
 Explaining PETs.Explaining PETsExplaining PETs.
-To address these misunderstandings and evaluate the impact of PETs on downstream factors such as privacy concerns (or expectations) and willingness to share data, technologists and researchers have worked to explain PETs to the public to address their unfamiliarity with privacy concepts (Cranor et al., 2006).
-Prior work focuses heavily on E2EE and DP (Abu-Salma et al., 2018; Bai, 2019; Cummings et al., 2021; De Luca et al., 2016; Akgul et al., 2021; Dechand et al., 2019; Abu-Salma et al., 2017; Wu and Zappala, 2018; Xiong et al., 2020; Nanayakkara et al., 2023).
+To address these misunderstandings and evaluate the impact of PETs on downstream factors such as privacy concerns (or expectations) and willingness to share data, technologists and researchers have worked to explain PETs to the public to address their unfamiliarity with privacy concepts [74].
+Prior work focuses heavily on E2EE and DP [28, 30, 27, 31, 29, 32, 33, 34, 35, 36].
 These efforts remain ongoing, as effectively and scalable setting privacy expectations remains a challenge.
-Methods found effective to explain PETs include visualizations (Stransky et al., 2021), mental models (Stewart and Lacey, 2012; Lin et al., 2012), nutrition labels (Kelley et al., 2010), metaphors (Karegar et al., 2022), short statements (Papacharissi and Fernback, 2005), and privacy games (Tupsamudre et al., 2018; Karl and Tao, 2023).
-In this work we use short statements to briefly explain the relevant guarantees in order to evaluate the impact of these guarantees on participants’ privacy expectations and willingness to donate medical data.
+Methods found effective to explain PETs include visualizations [75], mental models [76, 77], nutrition labels [78], metaphors [79], short statements [80], and privacy games [81, 82].
+However, there is a gap in the literature regarding techniques to explain auditing guarantees, despite the importance of auditing in verifying compliance with privacy promises.
 […]
 Privacy guarantees (PGs) and auditing.Privacy guarantees (PGs) and auditingPrivacy guarantees (PGs) and auditing.
-Prior work on the effect of PETs and their accompanying PGs typically focuses on the effect of a combination of guarantees offered by a specific PET (e.g., DP (Valdez and Ziefle, 2019; Cummings et al., 2021) or E2EE (Dechand et al., 2019)) on privacy concerns .
-A more limited subset of studies evaluate the impact on willingness to share data (Valdez and Ziefle, 2019).
-Instead of focusing on a specific PET and it’s combined accompanying PGs we seek to examine and compare four PGs that are shared across many PETs: data expiration, data anonymization, use restriction, access control, and auditing (see § 1).
+Prior research on PETs and their accompanying PGs has primarily focused on evaluating specific technologies (e.g., DP [83, 27] or E2EE [32]) and their combined impact on privacy concerns.
+A smaller subset of studies has assessed how these technologies affect willingness to share data [83].
+Additional research has examined public understanding of specific guarantees such as data retention [84], data anonymization [85], and secondary use permissions [86].
+Our approach differs by systematically comparing four core PGs implemented across many PETs: data expiration, data anonymization, use restriction, and access control (see § I).
+We also separately investigate auditing mechanisms, which serve as verification procedures rather than direct guarantees.
 In particular, we examine whether offering a given PG influences people’s willingness to donate health data to the recipient entity.
 Furthermore, we focus on the effect of two different auditing processes – expert auditing and self auditing – on enhancing privacy expectations and willingness to share data.
-Other studies have examined public understanding of data retention (Murillo et al., 2018) and general data anonymization (Haddow et al., 2011), and one study examines willingness to donate medical data if it is allowed to be used for secondary purposes (Juga et al., 2021).
-To our knowledge, no prior work focuses on comparing the effects of the four core PGs we study nor the effects of auditing PETs on user privacy expectations or willingness to donate data.
-3. Methodology
-As mentioned in § 1, we address research questions RQ1 and RQ2 through a user survey.
-In this section, we discuss the ethical considerations of our study, outline the survey design (§ 3.2), and describe the cognitive interview process used to ensure participants’ understanding of the presented scenarios (§ 3.3).
-We also describe our participant selection strategy (§ 3.4) and the limitations of our methodology (§ 3.6).
-3.1. Ethics
-Before we began our research, our ethics review board approved all of the study methods, including the cognitive interviews and the survey.
-A consent form detailing the data storage, handling, and survey compensation is presented at the beginning of the survey.
-Participants can opt out of disclosing any demographic data without influencing their compensation in the survey.
-Each participant received $1.20\$$ upon completing the survey.
+To our knowledge, this is the first study to systematically compare these four core PGs and evaluate how different auditing mechanisms affect user privacy expectations and willingness to donate health data.
+III Methodology
+As mentioned in § I, we address research questions RQ1 and RQ2 through a user survey.
+In this section, we discuss the ethical considerations of our study (§ III-A), outline the explored statements (§ III-B) and survey design (§ III-C), and describe the cognitive interview and pilot study process used to refine and improve participants’ comprehension of the presented scenarios (§ III-D).
+We also describe our participant selection strategy (§ III-E), analysis procedure (§ III-F), and the limitations of our methodology (§ III-G).
+III-A Ethics
+Our study was conducted under approval from our university’s Ethics Review Board (ERB).
+The approval covered both the cognitive interviews and the survey.
+We implemented several measures to ensure ethical treatment of participants:
+All survey data was collected anonymously through Qualtrics [87], with IP address collection disabled.
+Participants were presented with a consent form at the beginning of the survey that detailed the study purpose, data handling procedures, and compensation details.
+Participants could opt out of answering any demographic questions without affecting their compensation.
+For cognitive interviews conducted via Zoom [88], participants were asked to avoid using their real names when joining the meetings to maintain anonymity.
+Only anonymized transcripts were retained for analysis after removing any personal identifiable information.
+All research participants were recruited anonymously through Prolific [89], providing an additional layer of separation between researchers and participants’ identities.
+The collected data was accessible only to the research team, and all data will be deleted after the completion of the research project in accordance with our ERB protocol.
+Each participant received 1.20\$1.20\\$ upon completing the survey.
 To determine this figure, we ran a test of 20 participants, who took 5m56s (median) to complete it.
-Thus, the actual compensation was $12\$$/hr, which aligns with Prolific recommendations.
-For the signup survey of the cognitive interview, we set the study name to “Sign Up for Paid Interview on Data Donation” with a payment of $1.25\$$ ($13.68\$$/hr based on a loose completion estimate of 5 minutes).
-However, most respondents refused our interview request.
-Thus, in order to encourage participants interested in attending the interview to complete the screener, we lowered the compensation for the screening survey to $9.96\$$/hr.
-The interview reward is $15\$$, and all interviews lasted at most thirty minutes.
-The survey is conducted through Qualtrics (Qualtrics, 2020), with the IP address collection feature disabled to protect respondents’ privacy.
-Interviews are conducted via Zoom (zoom, 2024), and participants are instructed to avoid using their real names when joining the meeting to ensure anonymity.
-Research participants are anonymously recruited through Prolific (Prolific, 2024), as described in greater detail below.
-3.2. Survey Structure
-Figure 1.
-Survey design.
+Thus, the actual compensation was 12\$12\\$/hr, which aligns with Prolific recommendations.
+Interview participants were first recruited through a screening survey that collected demographic information and assessed willingness to participate in interviews.
+Initially, we compensated participants 1.25\$\\$ for completing this 5-minute screener (approximately 13.68\$\\$/hr), titled “Sign Up for Paid Interview on Data Donation.” However, this above-average compensation led to only 8% interview attendance, as many users were motivated primarily by the screening fee.
+To better align incentives and improve participation rates, we adjusted the screener compensation to 0.83\$\\$ (approximately 9.96\$\\$/hr), which better reflected platform norms for screening surveys.
+This adjustment increased our interview attendance rate to 21%.
+Participants who completed the 30-minute interviews were additionally compensated 15\$\\$.
+III-B Donation Scenarios
+PG(1) Anonymization: The privacy-preserving technology removes any personal identifiable information at the time of data collection, so that the data stored by the recipient entity is not linkable to its data owner (i.e., the data is anonymous).
+PG(2) Access control: The privacy-preserving technology restricts data access to the authorized scientists (within the recipient entity) which are working on the treatment for the given disease.
+PG(3) Data expiration: The privacy-preserving technology discards the donated data, or makes it inaccessible after a given expiration time, which can be chosen by the data donor in the data collection agreement.
+PG(4) Purpose restriction: The privacy-preserving technology ensures that the recipient entity can only use the donated data to develop a treatment for the given disease, and not in the context of any other research they may be working on (e.g., different disease).
+Baseline: N/A: (No privacy guarantee is at all mentioned.)
+Figure 1: Privacy statements PG(1)–PG(4). One of the five statements (including the empty baseline statement) is randomly presented in the donation scenario immediately after the survey introduction.
+AG(1): Expert auditing: An external advisory board of scientists and software engineers appointed by the recipient entity will regularly verify that the privacy-preserving technology is working as described.
+The results of this verification will be made public.
+AG(2): Self-auditing: Anyone interested, including the respondent and the experts the respondent trusts, will be able to verify that the privacy-preserving technology is working as described.
+Anyone, including the respondent and the experts the respondent trusts, can make their verification results public.
+Baseline: N/A: (No auditing statement is at all mentioned.)
+Figure 2: Auditing statements AG(1)–AG(2). One of the three statements (including the empty baseline statement) is randomly presented in the donation scenario after a non-control privacy statement.
+Many PETs involve complex mechanisms that cannot be adequately explained in brief consent forms without compromising informed consent principles.
+Additionally, consent forms must satisfy legal and ethical requirements while remaining accessible, as participants typically spend just minutes reviewing them [90, 91, 92].
+To address this challenge, we align with real-world consent practices [42, 41] by measuring reactions to the high-level guarantees that PETs provide rather than explaining their technical implementations.
+This approach allows us to investigate fundamental privacy guarantees that appear consistently across different technologies—guarantees that are both meaningful to users and reflective of actual implementations:
+  1. PG(1):
+    Anonymization [93, 94]: data is not linkable to its owner (as defined in [95]);
+  2. PG(2):
+    Access control [4, 96, 97, 98]: data is accessible only by authorized people, specified in the data collection agreement;
+  3. PG(3):
+    Data expiration [99, 100, 101]: data is discarded or inaccessible after a given expiration time, specified in the data collection agreement;
+  4. PG(4):
+    Purpose restriction [24, 102]: data is only used for the stated collection purpose, specified in the data collection agreement;
+In practice, implementing all guarantees simultaneously can be costly and challenging for real-world systems [43, 44].
+This constraint is reflected in many consent frameworks, which often focus on providing a single guarantee—such as anonymization [42] or access control [41].
+Following this practical approach, our study investigates the impact of each guarantee individually, rather than presenting them in combination.
+This targeted analysis enables us to identify which specific protections most significantly influence privacy expectations and, consequently, reported data-sharing intentions.
+Beyond these privacy guarantees, the technical community has developed methods to verify that PETs function as intended.
+We also investigate how these auditing mechanisms affect user trust and willingness to donate data, specifically examining two approaches [^1] We considered but excluded government auditing as a separate category due to overlap with “expert auditing” and potential political bias (see § II).:
+  1. AG(1):
+    Expert auditing [103, 104]: engaging an aggregator-selected external advisory board to audit the system to verify that the PET functions as described in the data collection agreement;
+  2. AG(2):
+    Self auditing [105, 106]: granting anyone, including donors or external advisors appointed by them, the ability to perform such audits.
+We present all statements in plain text format, avoiding metaphors and visual aids due to their mixed effectiveness in privacy communications [107, 108] and to align with standard practice in medical and survey consent forms [41, 42].
+Figure 1 and Figure 2 show the exact text presented (e.g., the factor levels).
+Justification on selecting the included PETs is in § -D.
+III-C Survey Structure
+Figure 3: Survey design.
 Each participant is presented with a donation scenario and a survey introduction with two variations: recipient entity and disease.
@@
 A sample screen that participants would encounter is provided in Figure 6.
-Our survey design is shown in Figure 1, and the full survey can be found in Appendix A. First, we present to the respondents a suvery introduction, which we detail in Figure 2.
+Our survey design is shown in Figure 3, and the full survey can be found in § -A. First, we present to the respondents a suvery introduction, which we detail in Figure 4.
 Survey introduction
@@
 Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
-Figure 2. Survey Introduction. The entity type (for-profit or non-profit) and the disease (selected from a list of common chronic diseases) are randomly selected.
+Figure 4: Survey Introduction. The entity type (for-profit or non-profit) and the disease (selected from a list of common chronic diseases) are randomly selected.
 We present a donation scenario using this introduction: we want to assess respondents’ willingness to donate their health data to a recipient entity developing a new treatment for a specific disease.
-The type of entity (for-profit or non-profit) and the disease (cancer, diabetes, heart failure, high blood pressure, and stroke) are randomly selected; the disease options are taken from a list of common chronic diseases published by the Centers for Medicare and Medicaid Services (CMS) (for Medicare and Services, 2024).
+The type of entity (for-profit or non-profit) and the disease (cancer, diabetes, heart failure, high blood pressure, and stroke) are randomly selected; the disease options are taken from a list of common chronic diseases published by the Centers for Medicare and Medicaid Services (CMS) [109].
 Then, we present to the respondents either a control (no statement) or an experimental statement.
@@
 Privacy statements are uniformly selected at random from a pool of four privacy statements; auditing statements are selected at random from two auditing statements.
-We have written these statements to align with PG(1)–PG(4) (see § 1), and we list them in Figure 3.
-We further distinguish the auditing guarantee into two different auditing statements to align with AG(1)–AG(2), and we list them in Figure 4.
-PG(1): Anonymization — personally identifiable data will not be collected.
-The privacy-preserving technology removes any personal identifiable information at the time of data collection, so that the data stored by the recipient entity is not linkable to its data owner (i.e., the data is anonymous).
-PG(2): Access control — only relevant scientists can read data.
-The privacy-preserving technology restricts data access to the authorized scientists (within the recipient entity) which are working on the treatment for the given disease.
-PG(3): Data expiration — data will be deleted later.
-The privacy-preserving technology discards the donated data, or makes it inaccessible after a given expiration time, which can be chosen by the data donor in the data collection agreement.
-PG(4): Purpose restriction — data will not be used for other purposes.
-The privacy-preserving technology ensures that the recipient entity can only use the donated data to develop a treatment for the given disease, and not in the context of any other research they may be working on (e.g., different disease).
-Baseline: Control
-No privacy-preserving guarantee is at all mentioned.
-Figure 3. Privacy statements PG(1)–PG(4). One of the five statements (including the control statement) is randomly presented in the donation scenario immediately after the survey introduction.
-AG(1): Expert auditing — the implementation of PET will be verified by experts.
-An external advisory board of scientists and software engineers appointed by the recipient entity will regularly verify that the privacy-preserving technology is working as described.
-The results of this verification will be made public.
-AG(2): Self-auditing — the implementation of PET can be verified by anyone.
-Anyone interested, including the respondent and the experts the respondent trusts, will be able to verify that the privacy-preserving technology is working as described.
-Anyone, including the respondent and the experts the respondent trusts, can make their verification results public.
-Baseline: Control
-No auditing is at all mentioned.
-Figure 4. Auditing statements AG(1)–AG(2). One of the three statements (including the control statement) is randomly presented in the donation scenario after the privacy statement.
+Specifically, one statement from Figure 1 and one from Figure 2 are presented to each respondent.
 Our survey questions assess respondents’ self-reported:
-  • scenario understanding, used to filter out respondents that report not understanding the scenario (see Section 3.4): “How would you rate your understanding of the above scenario?” (4 point Likert scale: Fully understand - Not understand).
+  • scenario understanding, used to filter out respondents that report not understanding the scenario (see § III-E): “How would you rate your understanding of the above scenario?” (4 point Likert scale: Fully understand - Not understand).
   • willingness to donate their health data to the recipient entity: “In this scenario how likely would you be to donate your medical record?” (4 point Likert scale: Very likely - Very unlikely).
   • privacy expectations regarding the specific privacy guarantees we investigate, measured by their agreement with the statements presented in Figure 5 using the same Likert scale as willingness to donate.
-[^1] This question included an additional attention check with a sixth statement: “If I donate my data, I will meet Albert Einstein.” Respondents who did not answer ‘Very Unlikely’ were removed from the experiment, as detailed in Section 3.4.
+[^2] This question included an additional attention check with a sixth statement: “If I donate my data, I will meet Albert Einstein.” Respondents who did not answer ‘Very Unlikely’ were removed from the experiment, as detailed in § III-E.
 We assess participants’ expectations about all guarantees, regardless of what statement they were presented.
   • trust that the recipient entity will implement the protection described by the privacy statement: “I trust the entity will handle my data as described.” (4-point Likert scale Strongly agree - Strongly disagree).
-  • demographics and experiences, including age, gender, education level, donation history, technical background (i.e., education background and job field relationships with computer technologies), and egocentricity(i.e., whether respondents or their close relatives have the disease, as defined in (Grgić-Hlača et al., 2020; Plane et al., 2017)).
+  • demographics and experiences, including age, gender, education level, donation history, technical background (i.e., education background and job field relationships with computer technologies), and egocentricity(i.e., whether respondents or their close relatives have the disease, as defined in [110, 111]).
 Anonymization
@@
 Purpose restriction
-The donated medical records will be used for another purpose without my consent
+The donated medical records will be used for another purpose without my consent.
 Expert auditing
@@
 I will be able to hire someone to verify that my medical record is protected as described.
-Figure 5. To measure privacy expectations respondents reported their agreement with each statement listed above on a 4-point Likert Scale from “Very Likely” to “Very Unlikely”.
+Figure 5: To measure privacy expectations respondents reported their agreement with each statement listed above on a 4-point Likert Scale from “Very Likely” to “Very Unlikely”.
+Figure 6: Screenshot of the survey, which consists of three segments.
+The first segment presents the survey introduction, varying the disease and the entity across participants.
+The second segment presents the randomized privacy statement and auditing statement, highlighted with color coding for emphasis.
+The third segment poses the survey question that participants are required to answer.
 We also ask respondents to explain their responses to the willingness to donate and trust questions.
 These explanations are collected through open-ended questions: “Why you are willing (or unwilling) to share your medical record with this entity?” and “Please explain why you do (or do not) trust that the entity will handle your data as described.” We note that the scenario understanding and trust questions are only asked when a non-control privacy statement is presented.
-A sample screen that participants would encounter is provided in Figure 6.
-Imagine that a non-profit organization (e.g., university, non-profit research institute) wants to develop a new treatment for Hypertension (High blood pressure).
-They need medical data from people with and without Hypertension (High blood pressure) to develop the treatment.
-They ask you to donate your medical record to help develop the treatment.
-Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
-The non-profit organization uses a privacy-preserving technology to ensure that your data is deleted after a set period of time.
-You will be able to choose from a list of time period options how long your data will be stored.
-Anyone interested, including you and experts you trust, will be able to verify that the privacy-preserving technology is working as described.
-Anyone, including you and experts you trust, can post their verification results publically.
-How would you rate your understanding of the above scenario? Fully understand Mostly understand Partially understand Not understand
-Figure 6.
-Sample survey screen that composed of three segments.
-The first segment presents the survey introduction, which varies the disease and the entity among all participants.
-The second segment includes the randomized privacy statement and auditing statement, with color coding applied to emphasize these elements.
-The third segment contains the survey question that participants are required to answer.
-3.3. Survey Refinement
-The final survey was refined through multiple iterations.
-Before launching the survey, we prescreened 49 users on the Prolific online platform (Prolific, 2024) by asking them to complete a short demographics survey and indicate their willingness to participate in a cognitive interview.
-Then, we conducted 17 cognitive interviews to ensure participants understood the survey correctly.
+A screenshot of the survey is presented in Figure 6 to illustrate the survey design.
+III-D Survey Refinement
+We refine our survey thourgh a pilot study and multiple iterations of cognitive interviews, following the design of [90]. Inividuals are only allowed to participate in one of the studies.
+We use the pilot study to test the survey design and the donation scenario.
+We found, for example, that no significant relationship between different diseases and donation willingness($p=0.32$, $\chi^{2}=17.00$) in the pilot survey ($N=81$ in the control condition).
+We thus present five common chronic diseases in the final survey (see § -A) to ensure participants can relate to the donation scenario and catch positive egocentricity measurement.
+We list the major takeaways in § -C.
+We conducted 17 cognitive interviews with participants recruited through Prolific [89], selected from 49 users who completed our screening survey and indicated their willingness to participate in the interview.
+These interviews evaluated participants’ understanding of the survey content and helped facilitate consistent comprehension across respondents.
 Each interview involved presenting the survey to participants and actively seeking feedback on their interpretation of the data donation scenario, and the privacy and auditing statements.
@@
 We continued refining the survey through cognitive interviews until no further constructive feedback was received.
-In the final interview, we observe a clear understanding of the statements and the donation scenario.
-For example, the participant described anonymization (PG(1)) as ”eliminating any sort of demographic or personal information affiliated with your data that could be associated with you if the data were to somehow be leaked” and access control (PG(2)) as ”some sort of database that only certain individuals have access to by using a passcode and certain credentials to to log in.” Our final survey can be found in Appendix A.
-3.4. Survey Sampling
-Following the recommendation of (Tang et al., 2022), we recruited participants through Prolific, gathering responses from 560 individuals.
+In the final interview, we observed generally good understanding of the statements and the donation scenario.
+For example, the participant described anonymization (PG(1)) as “eliminating any sort of demographic or personal information affiliated with your data that could be associated with you if the data were to somehow be leaked” and access control (PG(2)) as “some sort of database that only certain individuals have access to by using a passcode and certain credentials to to log in.”
+III-E Survey Sampling
+Following the recommendation of [112], we recruited participants through Prolific [89] and collected 560 responses.
 To ensure focused results, we restricted respondents to adults residing in the U.S.
 and requested a gender-balanced distribution.
-Our recruitment included 272 men, 275 women, 9 nonbinary individuals, and 4 participants who chose not to disclose their gender.
-Of these 560 respondents, we sequentially excluded the following from the analysis: 35 that submitted incomplete responses; 20 that failed our attention check (§ 3.2); and 11 that indicated they did not understand the survey scenario (using scenario understand).
+Our recruitment included 272 men, 275 women, 9 nonbinary individuals, and 4 participants who chose not to disclose.
+Of these 560 respondents, we sequentially excluded the following from the analysis: 35 that submitted incomplete responses; 20 that failed our attention check (§ III-C); and 11 that indicated they did not understand the survey scenario (using scenario understand).
 Thus, our final dataset comprises 494 respondents.
 To enhance statistical power, we limit the levels of age and education to binary groups based on the data distribution among our participants.
-We report their demographics in Table 1 and the number of respondents assigned to each condition in Table 2.
-Description Category $n$
+We report their demographics in Table I and the number of respondents assigned to each condition in Table II.
+TABLE I: Participant demographics.
+We note that the second category of each demographical attribute (e.g., no technical background) is considered the baseline scenario during analysis, except for “age”, which is treated numerically.
+Description Category $n$ %
 Age 18 - 29 120 24.3%
@@
 No 293 59.31%
-Table 1.
-Participant demographics.
-We note that the second category of each demographical attribute (e.g., no technical background) is considered the baseline scenario during analysis, except for “age”, which is treated numerically.
+TABLE II: Number of respondents assigned to each condition (i.e., who saw each privacy statement and auditing statement, or no privacy statement (bottom row) or no auditing statement (columns 4 and 7).
 For-Profit Non-Profit
@@
 Ctrl. - - 19 - - 21
-Table 2. Number of respondents assigned to each condition (i.e., who saw each privacy statement and auditing statement, or no privacy statement (bottom row) or no auditing statement (columns 4 and 7).
-3.5. Analysis
-We analyzed the open-text questions about willingness to donate and trust (see § 3.2) using inductive-thematic open coding.
+III-F Analysis
+We analyzed the open-text questions about willingness to donate and trust (see § III-C) using inductive-thematic open coding.
 Two researchers independently coded each entry and generated a codebook from a random sample of at least 100 (20.2%) responses.
 Then, they composed a final codebook and double-coded all responses.
-Since all responses were double-coded and inconsistencies were resolved, we do not report inter-rater reliability (IRR) (McDonald et al., 2019).
+Since all responses were double-coded and inconsistencies were resolved, we do not report inter-rater reliability (IRR) [113].
+Ultimately, six responses did not fit this scheme and classified as ‘Other’.
+The codebook of the open-text responses is presented in § -B.
 In addition to presenting a descriptive analysis detailing the distribution of responses on the survey items that address our research questions, we construct logistic regression models to analyze factors related to two dependent variables: privacy expectations (RQ1) and willingness to donate (RQ2).
@@
 We built separate regression models for the two recipient entities: for-profit and non-profit.
-3.6. Limitations
-Although the four privacy-preserving guarantees examined in this study (§ 1) cover a broad range of PETs, they are not fully comprehensive and do not capture the full complexity of PETs or real-world threats, which may involve privacy issues beyond the scope of our investigation.
-The randomly presented disease may also influence the results.
-Furthermore, our approach may not fully reflect the intricacies of real-world donation scenarios, where multiple guarantees often coexist.
-While such complexities lie outside the primary focus of our research, qualitative investigations have been conducted in prior work (Aitken et al., 2016; Mori, 2016), and quantitative exploration of these aspects remains an interesting direction for future research.
+III-G Limitations
+Although the four privacy-preserving guarantees examined in this study (§ I) cover a broad range of PETs, they are not fully comprehensive and do not capture the full complexity of PETs or real-world threats, which may involve privacy issues beyond the scope of our investigation.
+Furthermore, our approach may not fully reflect the intricacies of real-world donation scenarios, where multiple guarantees may coexist.
+While such complexities lie outside the primary focus of our research, qualitative investigations have been conducted in prior work [2, 54], and quantitative exploration of these aspects remains an interesting direction for future research.
 Despite our efforts to mitigate misunderstandings—through rigorous cognitive interviews, filtering respondents who reported not fully understanding the scenarios, and controlling for privacy expectations in our statistical analyses—the brief descriptions of each guarantee might still have led to partial or incorrect understanding among respondents.
 This limitation may have influenced our results.
-Additionally, while our study employs vignette scenarios that effectively mirror real-world behaviors (Redmiles et al., 2017), there may still be discrepancies between participants’ expressed privacy concerns and their actual behaviors, a phenomenon commonly referred to as the privacy paradox (Barth and de Jong, 2017).
+Additionally, our study relies on self-reported data, which is subject to well-documented biases, known as the privacy paradox, where individuals’ expressed privacy concerns often differ from their actual behaviors [114].
+To address these limitations, we implemented several methodological safeguards.
+Firstly, we employed vignette-based scenarios, which research has shown to effectively mirror real-world behaviors and reduce hypothetical bias [115].
+Secondly, our work is informed by prior research from 2019 [116] and replicated in 2022 [117] that demonstrated crowdsourced samples, specifically from Prolific [117], well approximate the security and privacy behavior, expectations, and knowledge of the general population for US adults between 18-50 who have at least some college education.
+While there may still be discrepancies between participants’ expressed privacy concerns and their actual behaviors, research confirms that this overambitiousness is systematic [118, 92].
+It establishes survey data as a reliable upper-bound approximation of real-world behavior.
 Furthermore, statistically significant results do not inherently eliminate the risk of underlying biases that could skew the findings.
@@
 population, nor does it capture the perspectives of individuals from other countries, highlighting a key limitation of this work.
-4. Results
-4.1. RQ1: Understanding of PGs
-Next, we answer the first research question (RQ1, see § 1): “How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?”.
-We analyze and report the survey responses and evaluate the rationales for the responses.
-4.1.1. Analysis Setup
+IV Results
+IV-A RQ1: Understanding of PGs
+We first address RQ1 (see § I): “How well do people understand and expect privacy guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?” We analyze survey responses and their underlying rationales.
+IV-A1 Analysis Setup
 We compare privacy expectations on PG(1)—PG(4) and AG(1)–AG(2) among groups of respondents who received different privacy statements (listed in Figure 1) and auditing statements (listed in Figure 2).
 In Figure 7, we show the privacy expectations of respondents who were shown different privacy statements.
-To assess the statistical significance of the descriptive quantitative results presented above, we used binomial logistic regression to analyze the relationship between the presence of a privacy statement in the scenario and respondents’ privacy expectations.
+To assess the statistical significance of the descriptive quantitative results presented above, we used logistic regression to analyze the relationship between the presence of a privacy statement in the scenario and respondents’ privacy expectations.
 We summarize results, which we distinguish for non-profit and for-profit entities, in Figure 8.
-Figure 7.
-Percentage of respondents who had a positive privacy expectation when shown a particular privacy statement in their donation scenario.
+Figure 7: Percentage of respondents who had a positive privacy expectation when shown a particular privacy statement in their donation scenario.
 The values in the table are the percentage of respondents in a given condition who had the privacy expectation (see Figure 5) that corresponded to the privacy statement they were presented.
@@
 The agreements are binarized here to present the overall tendency.
-Figure 8.
-Odd ratios and 95% confidence intervals of presenting different privacy statements on the respondents’ privacy expectations toward the two kinds of recipient entities, for-profit and non-profit.
-4.1.2. Response Distribution and Statistical Analysis Results
+Figure 8: Odd ratios and 95% confidence intervals of presenting different privacy statements on the respondents’ privacy expectations toward the two kinds of recipient entities, for-profit and non-profit.
+IV-A2 Response Distribution and Statistical Analysis Results
 We find that privacy expectations differ based on the data-collecting entity.
-In the control group, which was not shown any privacy statement, 26%-37% of respondents expected for-profit entities to provide anonymization, data expiration, access control, use restriction to a specific purpose, expert auditing, and self-auditing.
+In the control group, which was not shown any privacy statement, 26%-37% of respondents expected for-profit entities to provide anonymization, data expiration, access control, purpose restriction, expert auditing, and self-auditing.
 In contrast, a higher percentage of respondents (43%-76%) expected non-profit entities to employ these mechanisms, except for self-auditing (AG2).
-For instance, we note that respondent P430 (who received a for-profit scenario) expected privacy protections even though they were in the control group (and thus received no statement about privacy protection): ”I also assume that the data is looked at in the aggregate and likely no one at the company knows me”.
-In the same situation, respondent P53 (who received a non-profit scenario instead) made a stronger assumption: ”I trust that the organization will uphold strict privacy and ethical standards”.
-In the for-profit scenario, all privacy statements, except for the access-control statement (PG2), significantly increase the corresponding privacy expectations ($p$-value ¡ $0.05$).
+For instance, we note that respondent P430 (who received a for-profit scenario) expected privacy protections even though they were in the control group (and thus received no statement about privacy protection): “I also assume that the data is looked at in the aggregate and likely no one at the company knows me”.
+In the same situation, respondent P53 (who received a non-profit scenario instead) made a stronger assumption: “I trust that the organization will uphold strict privacy and ethical standards.”
+In the for-profit scenario, all privacy statements, except for the access-control statement (PG2), significantly increase the corresponding privacy expectations ($p$-value < $0.05$).
 In the non-profit scenario, where expectations are already high, only the privacy statement about data expiration (PG3) significantly increases expectations (from 57% to 100%).
@@
 It also reduced the gap in expectations for data expiration (PG3, 67% and 100%)) and purpose restriction (PG4, 74% and 84%)).
-In both scenarios, the auditing statements increase the expectation that a given privacy guarantee will be audited ($p$-value ¡ $0.05$).
-However, we note that incorporating an audit statement does not markedly change privacy expectations.
+Auditing statements significantly increased expectations of corresponding audits in both for-profit and non-profit scenarios ($p$-value < $0.05$).
+However, we note that incorporating an audit statement does not significantly change privacy expectations.
 The statistical significance of the auditing statements in Figure 8 indicates that respondents shown the audit statement correctly expected the corresponding type of audit, while respondents that were not shown the auditing statement did not expect this protection.
 As observed in  Figure 7, presenting an auditing statement does not increase the number of respondents who expect the privacy guarantee implied by the privacy statement they were shown, despite the stronger assurances that audits imply.
-4.1.3. Qualitative Analysis of Privacy Expectations via Open-Answer Responses
+IV-A3 Qualitative Analysis of Privacy Expectations via Open-Answer Responses
 Across for-profit and non-profit scenarios, 69% and 85% of respondents, on average, expressed that their positive privacy expectations were formed based on the presence of the privacy statement, belief in privacy obligations of the entity, trust in the entity, or trust in the auditing process.
@@
 16.3% and 20.8% of respondents in the for-profit and non-profit scenarios, respectively, were convinced by the received privacy statement.
-Some respondents, like P56, vaguely stated that the policy looks promising: ”I see their policy, and they have to follow their own policy”.
-Others felt explicitly safer and trusted the statement.
-For example, P494 noted: ”Trust is very important when it comes to medical data.
+Some respondents, like P56, vaguely stated that the policy looks promising: “I see their policy, and they have to follow their own policy.” Others felt explicitly safer and trusted the statement.
+For example, P494 noted: “Trust is very important when it comes to medical data.
 I believe the organization has privacy policies that outline that they collect and will use my data.
-I also believe the organization will employ security measures to safeguard data”.
-Similarly, P555 highlighted the presence of the privacy statement: ”(…) purposely states there is software in place to conserve privacy”.
-and P516 felt their information was safe because of it: ”(…) with the privacy protection in place, they are isolating the data they need while basically ”throwing out” the rest by putting in under that protection.
-In essence, my information is safe, and they’re only using what they said they’d use”.
-P208 also felt safer contributing to research knowing the privacy mechanisms were in place: ”I want to be able to contribute to research to better improve cancer treatment, and I feel safe if my data is protected through the mechanisms above”.
+I also believe the organization will employ security measures to safeguard data.” Similarly, P555 highlighted the presence of the privacy statement: “(…) purposely states there is software in place to conserve privacy.” and P516 felt their information was safe because of it: “(…) with the privacy protection in place, they are isolating the data they need while basically ‘throwing out’ the rest by putting in under that protection.
+In essence, my information is safe, and they’re only using what they said they’d use.” P208 also felt safer contributing to research knowing the privacy mechanisms were in place: “I want to be able to contribute to research to better improve cancer treatment, and I feel safe if my data is protected through the mechanisms above.”
 […]
-Belief in legal and reputational obligations..Belief in legal and reputational obligations.Belief in legal and reputational obligations..
-16.3% and 10.6% of respondents in the for-profit and non-profit scenarios, respectively, believed that entities are forced to protect the donated data due to legislative requirements and reputational concerns.
-For example, P28 stated regarding for-profits: ”A for-profit organization wouldn’t want to violate HIPAA, HITECH laws”.
-Additionally, P117 wrote ”I would trust them to do the right thing so they won’t face lawsuits”.
-Regarding non-profits, P141 noted: ”I trust that the law will restrict any data leaks to third parties”.
+Belief in legal or reputational obligations..Belief in legal or reputational obligations.Belief in legal or reputational obligations..
+16.3% and 10.6% of respondents in the for-profit and non-profit scenarios, respectively, believed that entities are forced to protect the donated data due to legislative requirements or reputational concerns.
+For example, P28 stated regarding for-profits: “A for-profit organization wouldn’t want to violate HIPAA, HITECH laws.” Additionally, P117 wrote “I would trust them to do the right thing so they won’t face lawsuits.” Regarding non-profits, P141 noted: “I trust that the law will restrict any data leaks to third parties.”
 […]
@@
 13.9% and 17.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed a general trust in the recipient entity without specifying reasons.
-For instance, P70 succinctly said ”I feel they are reliable and trustworthy” and P50 stated that ”I assume they take their research seriously, so they would handle the data carefully”.
-Other respondents, like P145, reported having no reason not to trust it: ”I have no reason to think they would do anything nefarious with my medical data”.
+For instance, P70 succinctly said: “I feel they are reliable and trustworthy” and P50 stated that “I assume they take their research seriously, so they would handle the data carefully.” Other respondents, like P145, reported having no reason not to trust it: “I have no reason to think they would do anything nefarious with my medical data.”
 […]
@@
 7.8% and 7.9% of respondents in the for-profit and non-profit scenarios, respectively, found the auditing statement to add at least some reliability.
-For example, P202 showed some reservation but found confidence in the public nature of audits: ”I don’t fully trust them, but I somewhat do, particularly if audits and verification of results are made public.
-That said, claiming that the advisory board is external is only partly reassuring, as it’s appointed by the institute”.
-Additionally, P540 had trust in the entity’s data handling because ”they let outsiders audit them” and P51 (who received the expert-auditing statement) because of ”safelocks and checks in place”.
-On the other hand, one interviewee doubted the expert auditing process and stated: ”somebody else says something doesn’t mean that it’s real”.
+For example, P202 showed some reservation but found confidence in the public nature of audits: “I don’t fully trust them, but I somewhat do, particularly if audits and verification of results are made public.
+That said, claiming that the advisory board is external is only partly reassuring, as it’s appointed by the institute.” Additionally, P540 had trust in the entity’s data handling because “they let outsiders audit them” and P51 (who received the expert-auditing statement) because of “safelocks and checks in place”.
+On the other hand, one interviewee doubted the expert auditing process and stated: “somebody else says something doesn’t mean that it’s real”.
 No survey respondent explicitly raised concern towards the auditing process.
@@
 21.2% and 10.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed general distrust in the feasibility of the privacy statement.
-In the non-profit scenario, P34 stated that ”no privacy technology is foolproof”.
-Similarly, in the for-profit scenario, P45 wrote: ”I don’t trust that the privacy-preserving tech would work.” More precisely, privacy statements were considered too ideal to be fully enforced.
+In the non-profit scenario, P34 stated that “no privacy technology is foolproof”.
+Similarly, in the for-profit scenario, P45 wrote: “I don’t trust that the privacy-preserving tech would work.” More precisely, privacy statements were considered too ideal to be fully enforced.
 Some respondents believed that unauthorized employee access would be unavoidable.
-For instance, P71 said: ”Why would I trust someone other than my doctor with my medical records?
+For instance, P71 said: “Why would I trust someone other than my doctor with my medical records?
 These days especially, I don’t trust anyone.
 There could be a breach or simply people I don’t know from a hole in the wall will then have access to all my med records.
-Insane!”.
+Insane”!
 Some others felt that data breaches and cyberattacks were inevitable.
-For instance, P83 stated that ”the primary reason for my distrust is due to past news of companies being hacked by people and their data getting leaked”, while P124 noted that ”corporate data breaches are very common”, and P80 echoed this sentiment, saying: ”I believe the [intentions] will be good, but data can be hacked”.
+For instance, P83 stated that “the primary reason for my distrust is due to past news of companies being hacked by people and their data getting leaked”, while P124 noted that “corporate data breaches are very common”, and P80 echoed this sentiment, saying: “I believe the [intentions] will be good, but data can be hacked”.
 We observe that the auditing statement did not substantially instill trust on all accounts — with the exception of expert auditing for access control in non-profit scenario, and data expiration in for-profit scenario.
-For example, P474 mentioned: ”They can check their privacy technology all they want but when there is a breach it is done and info is stolen.
-After it fails then they say sorry and offer monitoring but the info is still stolen”.
+For example, P474 mentioned: “They can check their privacy technology all they want but when there is a breach it is done and info is stolen.
+After it fails then they say sorry and offer monitoring but the info is still stolen.”
 […]
 Doubt on the recipient entity’s motivation to employ PETs..Doubt on the recipient entity’s motivation to employ PETs.Doubt on the recipient entity’s motivation to employ PETs..
-8.6% and 3.3% of the for- and non-profit scenarios, respectively, claimed that these entities are inherently self-serving and lack the motivation to uphold privacy-preserving guarantees or implement such measures at all.
+8.6% and 3.3% of participants in the for- and non-profit scenarios, respectively, claimed that these entities are inherently self-serving and lack the motivation to uphold privacy-preserving guarantees or implement such measures at all.
 Most criticism was directed at for-profit entities.
-For instance, P55 noted that for-profit entities ”will do what is profitable and not much more than that”.
-Similarly, P119 stated ”(…) because it is a for-profit organization.
-I expect them to cut corners” and P32 wrote that ”for-profit organization have low standard of morality”.
+For instance, P55 noted that for-profit entities “will do what is profitable and not much more than that”.
+Similarly, P119 stated “(…) because it is a for-profit organization.
+I expect them to cut corners” and P32 wrote that “for-profit organization have low standard of morality”.
 However, some respondents also expressed concerns about non-profit entities.
-For example, P190 in the non-profit scenario remarked: ”Medicine has become a business.
+For example, P190 in the non-profit scenario remarked: “Medicine has become a business.
 My data is only useful to them if it helps them make more money.
-Money comes first before the actual well being of humans”.
-4.2. RQ2: Willingness to donate health data
-Next, we answer the second research question (RQ2, see § 1): “How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?”.
-As in the previous subsection, we statistically analyze the responses and then report the qualitative rationales collected from the open-text questions.
-4.2.1. Analysis Set Up
-In Figure 9, we summarize the respondents’ willingness to donate their health data to a recipient entity.
-The 1st column (“overall”) reports results for all entity types, the 2nd column reports the results of the subset with for-profit entity, and the 3rd column reports the results of the subset with non-profit entity.
+Money comes first before the actual well being of humans.”
+IV-B RQ2: Willingness to donate health data
+Next, we answer the second research question (RQ2, see § I): “How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?” As in the previous subsection, we statistically analyze the responses and then report the qualitative rationales collected from the open-text questions.
+IV-B1 Analysis Set Up
+In Figure 9, we summarize the respondents’ willingness to donate their health data to the presented recipient entity.
+The 1st column (“overall”) reports results for both entity types, the 2nd column reports the results of the subset with for-profit entity, and the 3rd column reports the results of the subset with non-profit entity.
 The values represent the percentage of respondents willing to donate their heath data to the recipient entity.
@@
 We first analyze and confirm the influence of the recipient entity and then separate the participants based on the recipient entities and analyze them in separate models.
-The results are summarized in Table 3.
+The results are summarized in Table III.
 In Figure 10, we visualize the overlap in experimental group respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described.
-4.2.2. Response Distribution and Statistical Analysis Results
-Figure 9.
-Percentage of respondents willing to donate their personal health data to the recipient entity in each scenario (e.g., the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios that with PG(1) presented but no auditing statement were willing to donate.) The agreements are binarized here to present the overall tendency.
-Figure 10.
-Overlap in respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described (respondents in the control – no privacy statement – are excluded).
+IV-B2 Response Distribution and Statistical Analysis Results
+Figure 9: Percentage of respondents willing to donate their personal health data to the recipient entity in each scenario (e.g., the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios that with PG(1) presented but no auditing statement were willing to donate.) The agreements are binarized here to present the overall tendency.
+Figure 10: Overlap in respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described (respondents in the control – no privacy statement – are excluded).
 The numbers outside each circle summarize the total number of respondents who were e.g., willing to donate their data (310).
 $N=454$.
+TABLE III: Influences of privacy expectations and demographics on willingness to donate of the three participant groups. OR: odd ratio. CI: confidence interval.
 Entity Factor Levels OR 95% CI $p$-value
@@
    Donation History 2.25 [1.06, 4.75] 0.033
-Table 3. Influences of privacy expectations and demographics on willingness to donate of the three participant groups. OR: odd ratio. CI: confidence interval.
 There was no direct relationship between showing a privacy statement and respondents’ willingness to donate.
@@
 Combined with the results of RQ(1), the findings suggest that willingness to donate is influenced by privacy expectations, which are, in turn, impacted by privacy statements.
-Privacy statements and expectations are not the only reason for donation intentions however.
-While participants were generally less willing to donate to for-profit entities (OR=0.38, $p=\textless 0.001$), when analyzing only control group participants, we observe that the absolute difference in willingness to donate between for-profit and non-profit entities was 8.8% was smaller than the 20.3% difference in privacy expectations between the two entities.
-This indicates that other non-privacy-related factors, such as the perceived impact or benefits of the donations, also influence participants’ willingness to donate.
-Furthermore, we observe that respondents with bachelor’s degrees are less willing to donate to a non-profit entity (OR = 0.46, $p=0.007$), which may be related to their greater awareness of the complexities and potential vulnerabilities with privacy statement.
+Privacy statements and expectations influence donation intentions, but they are not the only factors.
+While participants were generally less willing to donate to for-profit entities (OR=0.38, $p$-value < $0.001$), other factors are also at play.
+Looking at control group participants (who received no privacy statements), we found that the difference in willingness to donate between for-profit and non-profit entities was only 8.8%, despite a much larger 20.3% difference in privacy expectations between these entities.
+This discrepancy suggests that non-privacy-related factors, such as perceived donation benefits, also influence participants’ willingness to donate.
+Furthermore, we observe that respondents with bachelor’s degrees are less willing to donate to a non-profit entity (OR = 0.46, $p$-value = $0.007$), which may be related to their greater awareness of the complexities and potential vulnerabilities with privacy statement.
 They have heightened privacy concerns that need to be addressed.
-For example, P42 wanted to verify that the PET worked visually; P97 thought that only the stored data was encrypted, but the data transaction was not protected the same way; and P236 argued that: “privacy-preserving technology that works today may not work in a few years.”
+For example, P42 wanted to verify that the PET worked visually; P97 thought that only the stored data was encrypted, but the data transaction was not protected the same way; and P236 argued that: “privacy-preserving technology that works today may not work in a few years”.
 Finally, we found that prior donors are more willing to donate to both entities.
-Respondents who have donated before are more willing to donate to both entities, with the OR value of 3.25 ($p=0.011$) and 2.25 ($p=0.033$) for the for and non-profit entities, respectively.These respondents expressed strong willingness of supporting science and recognized the importance of medical data in developing new treatments.
-For example, P40 claimed: “I feel that it is important information to share in hopes that they can find better ways to deal with diabetes.” and P208 responded: “Breakthroughs in science and moving forward in knowledge, our greatly benefited by such Data Collection”.
-4.2.3. Qualitative Analysis of Donation Intention Via Open-Answer Responses
+Respondents who have donated before were more willing to donate to both entities, with the OR value of 3.25 ($p$-value = $0.011$) and 2.25 ($p$-value = $0.033$) for the for- and non-profit entities, respectively.These respondents expressed strong willingness of supporting science and recognized the importance of medical data in developing new treatments.
+For example, P40 claimed: “I feel that it is important information to share in hopes that they can find better ways to deal with diabetes”.
+and P208 responded: “Breakthroughs in science and moving forward in knowledge, our greatly benefited by such Data Collection.”
+IV-B3 Qualitative Analysis of Donation Intention Via Open-Answer Responses
 Many respondents reasons for donating (or not) included a desire to support research, ego-centric connections, or desire for personal reward.
 In the for-profit scenario, 50.2% of respondents’ reasons for donating (or not) fell into these categories, in the non-profit scenario, 53.3% did.
-One non-privacy-related reason for donating included supporting research: e.g., P107 expressed that ”there are a number of things I could share to research such as money, time, data, and more.
-I would be willing to share my medical records to contribute to cancer research in hopes that people will be healthier and to enhance research future medicines.” Similarly, P97 stated: ”even if they are for-profit (which I don’t like), any closer we get to developing better treatments and/or a potential cure to cancer is something I’d be willing to assist in.”
+One non-privacy-related reason for donating included is helping research.
+For example, P107 wrote that “there are a number of things I could share to research such as money, time, data, and more.
+I would be willing to share my medical records to contribute to cancer research in hopes that people will be healthier and to enhance research future medicines”.
+P97 stated: “even if they are for-profit (which I don’t like), any closer we get to developing better treatments and/or a potential cure to cancer is something I’d be willing to assist”.
 Ego-centric reasons, particularly familial connections, also motivated donations.
-P154 mentioned donating because ”my dad died when his heart failed on him, and I’d have rather that not had happened”, and P233 noted that they ”would be willing to help to prevent someone else’s mom from dying of a stroke like [theirs] did.”
+P154 mentioned that the motivation was: “my dad died when his heart failed on him, and I’d have rather that not had happened”, and P233 noted that they “would be willing to help to prevent someone else’s mom from dying of a stroke like [theirs] did”.
 However, their existing (negative) experiences with for-profit healthcare entities also deter some of them from making donations.
-For example, P136 claimed they have heard a number of times that the entity says they won’t sell your information but can get around it by calling it something else, thus technically selling the information.
-P244 also mentioned that their information has been leaked before by a clinician.
-Some respondents desired personal reward for donating and thus were not interested in donating in our scenario: e.g., P92 explained that ”I would give them my data if I got paid for it.
-I would think twice if they want me to donate the data and make no money from it while the researchers will make money off the of research they compiled.”
-An alternative explanation could be the perceived imbalance in cost and profit between the donor and the recipient entity, where the donor provides data for free while the recipient entity profits from its use.
+For example, P136 claimed they have heard multiple instances where entities promise not to sell information but circumvent this by using different terminology, effectively still selling the data.
+P244 also mentioned that their information has been leaked by a clinician in the past.
+Some respondents desired personal reward for donating and thus were not interested in donating in our scenario: e.g., P92 explained that “I would give them my data if I got paid for it.
+I would think twice if they want me to donate the data and make no money from it while the researchers will make money off the of research they compiled.” An alternative explanation could be the perceived imbalance in cost and profit between the donor and the recipient entity, where the donor provides data for free while the recipient entity profits from its use.
 For instance, participants P495, P463, and P459 all share this perception.
 However, we note that some individual participants alternatively were motivated to donate due to egocentricity, even to for-profit entities.
-For example, P45 ”just lost my father to heart failure thirteen days ago.
-(…) I would be more than willing to donate my medical records if it helps develop medicines to make hearts function better.
-Respondents expressed negative sentiments toward for-profit entities collecting data for research.
+For example, P45 “just lost my father to heart failure thirteen days ago.
+(…) I would be more than willing to donate my medical records if it helps develop medicines to make hearts function better.”
+Respondents expressed negative sentiments toward for-profit entities collecting data.
 The desire for personal reward related closely to lack of donation intent for for-profit organizations.
 In the for-profit scenario, 11.4% explicitly refused due to negative perceptions of such organizations.
-For instance, P191 and P202 respectively stated: ”I have a negative connotation with for profit organization” and ”I feel like this type of organization already profits significantly off a large number of people (occasionally off of me as well) and as such, I do not want to give them direct permission to profit further off of me.” Furthermore, respondents felt that for-profits already benefit enough and were unwilling to contribute further.
-For instance, P36 wrote: ”For profit organization makes profit using the data.
+For instance, P191 and P202 respectively stated: “I have a negative connotation with for profit organization” and “I feel like this type of organization already profits significantly off a large number of people (occasionally off of me as well) and as such, I do not want to give them direct permission to profit further off of me.” Furthermore, respondents felt that for-profits already benefit enough and were unwilling to contribute further.
+For instance, P36 wrote: “For profit organization makes profit using the data.
 I’m never going to donate it.
 They can BUY it from me for a reasonable amount.”
-Alternatively, a small group of 1.3% respondents explicitly mentioned that they were willing to donate data to entities of non-profit nature, but even they expressed concerns about potential misuse or unethical behavior within these organizations.
-For instance, P122 mentions: ”it’s always going to be in the back of my head that there’s a possibility that it’s being sold or taken or used in some way I’m not okay with.
-Scares me a bit.” Then, P282 notes: ”it is difficult to police everyone in a non-profit organization.
+Alternatively, a small group of 1.3% respondents explicitly mentioned that they would donate to non-profits but remained concerned about potential data misuse.
+For instance, P122 mentions: ‘it’s always going to be in the back of my head that there’s a possibility that it’s being sold or taken or used in some way I’m not okay with.
+Scares me a bit.” Then, P282 notes: ‘it is difficult to police everyone in a non-profit organization.
 I have seen leaders in organizations act unethically at times, which makes me think that even a well-intended organization cannot fully control the actions of every employee or volunteer they have.”
 Privacy-related donation considerations focused on data sensitivity and leakage.
-Concerns for not donating were mostly about data leakage, with respondents expressing that they 12.9% and 13.9% participants in the for-profit and non-profit scenario, respectively, were protective of their data and hence would not donate at all, as P545 noted that ”there are so many data leaks from so called safe places.
-No one can anticipate what hackers can do in the future.” Others cited previous data leak experiences as to why they would not donate, as P290 said ”my medical data has been breached in the past by a clinician and I had to file a formal grievance against the health care system that employed her.
-For this reason, because my medical history was abused, I no longer have any trust and will never voluntarily consent to my medical history/information being shared with others.” Some respondents who would not donate data also expressed that they would only possibly donate to specific entities that have already proven trustworthiness to them, as P357 explained that ”regardless of privacy policy, the probability of a leak or misuse is high.
-unless it is an organization I have had personal interaction with, or am very familiar with, it is unlikely I would donate my medical records.”.
-Another reason was that the participant was concerned that their medical data included too much or too specific information.
-P155 highlighted this concern: ”I would not feel comfortable with so much of my personal health information being shared with an organization that is not involved with my direct medical care, regardless of whether there is an advisory board or not”; as did P229: ”It’s a lot of detailed information that I’m worried if it somehow gets in the wrong hands, it could reveal a lot of private information about me.” Contrastingly, some respondents felt comfortable sharing data due to a lack of sensitive information (e.g., ”I don’t have mayor illnesses or nothing to hide so Im ok with that”) or because they already share information with other entities (e.g., ”Considering I already disclose this information to other organizations (for example, data collection on phone, data compiled through search history), other agencies likely have the information on hand already and so another organization having it is no different”).
-5. Concluding Discussion
-Like our survey, when soliciting data donations, recipient entities usually provide an introduction motivating their need to collect the individual’s data, outline privacy guarantees, and seek consent for data donation.
-Privacy statements function similarly to advertising (O’Flaherty, 2022) by shaping and influencing consumers’ expectations of privacy (Schaub et al., 2015).
-Yet, these statements do not always have impact.
-Prior work has highlighted the ways in which such statements can fail due to being vague or inaccurate.
-Our work highlights how the impact of privacy statements is moderated by people’s pre-existing privacy expectations.
+Concerns for not donating were mostly about data leakage, with respondents expressing that they 12.9% and 13.9% participants in the for-profit and non-profit scenario, respectively, were protective of their data and hence would not donate at all, as P545 noted that “there are so many data leaks from so called safe places.
+No one can anticipate what hackers can do in the future.” Others cited previous data leak experiences as to why they would not donate, as P290 said: “my medical data has been breached in the past by a clinician and I had to file a formal grievance against the health care system that employed her.
+For this reason, because my medical history was abused, I no longer have any trust and will never voluntarily consent to my medical history/information being shared with others.” Some respondents also expressed that they would only possibly donate to specific entities that have already proven trustworthiness to them, as P357 explained that “regardless of privacy policy, the probability of a leak or misuse is high.
+unless it is an organization I have had personal interaction with, or am very familiar with, it is unlikely I would donate my medical records.”
+Some participants expressed concern about the level of detail in medical data.
+P155 highlighted this concern: “I would not feel comfortable with so much of my personal health information being shared with an organization that is not involved with my direct medical care, regardless of whether there is an advisory board or not”; as did P229: “It’s a lot of detailed information that I’m worried if it somehow gets in the wrong hands, it could reveal a lot of private information about me.” Contrastingly, some respondents felt comfortable sharing data due to a lack of sensitive information (e.g., “I don’t have mayor illnesses or nothing to hide so Im ok with that”) or because they already share information with other entities (e.g., “Considering I already disclose this information to other organizations (for example, data collection on phone, data compiled through search history), other agencies likely have the information on hand already and so another organization having it is no different”).
+V Concluding Discussion
+Health data donation decisions involve multiple complex factors beyond privacy considerations.
+Non-privacy factors, such as recipient reputation and perceived societal benefit, often influence willingness to donate.
+Moreover, the effectiveness of PETs depends on people’s belief on these protections.
+PETs’ impact on donation willingness diminishes substantially when individuals misunderstand guarantees or doubt implementation integrity.
+Prior work has highlighted the ways in which statements on explaining PETs can fail due to being vague or inaccurate.
+Our work extends this understanding by showing that even when people clearly understand the protection guarantees, the perceived effectiveness of these guarantees is often limited by non-privacy-related factors.
 Our participants reported strong pre-existing privacy expectations, even in the absence of explicit privacy statements.
@@
 In contrast, the impact on non-profit organizations is minimal because users already hold high privacy expectations for these entities.
+This disparity in privacy expectations between non-profit and for-profit entities reflects deeper psychological mechanisms that influence how people form judgments about organizational trustworthiness and privacy practices.
 […]
 The Halo Effect.The Halo EffectThe Halo Effect.
-Participants qualitative responses suggest that their strongly positive privacy expectations for non-profit entities are rooted in arguments of morality: non-profits are perceived as more ethical and dedicated to serving the public good because they are not pursuing profit.
-This findings are consistent with a phenomenon termed “the halo effect”, which describes how an overall positive impression of a particular entity (e.g., the good work that a non-profit does) can lead to ungrounded assumptions about how the organization operates (Nisbett and Wilson, 1977; de Bruin Cardoso et al., 2024).
+Participants’ qualitative responses suggest that their strongly positive privacy expectations for non-profit entities arise from basic moral judgments: non-profits are perceived as more ethical and dedicated to serving the public good because they are not pursuing profit.
+This findings are consistent with a phenomenon termed “the halo effect”, which describes how an overall positive impression of a particular entity (e.g., the good work that a non-profit does) can lead to ungrounded assumptions about how the organization operates [119, 120].
 While prior work in the business literature focuses on how the halo effect influences expectations about business practices such as donation management, our work illustrates that the halo effect also influences people’s expectations about the way non-profits handle data privacy.
-These positive expectations do not always align with reality.
-Privacy safeguards are not consistently enforced or widely practiced; for example, many health organizations rely on broad consent frameworks rather than adopting stricter purpose-restricted consent measures (Kaye et al., 2015).
-This gap between expectations and reality can result in privacy expectation violations, a major source of privacy concerns as explained by the theory of contextual integrity (Nissenbaum, 2009).
-Indeed, a well-known non-profit mental health helpline shared it’s data with a commercial entity, leading to significant public outcry (Levine, 2022).
-Future work may seek to explore how to design privacy communications that set appropriate expectations in the presence of such effects, as a prior work has sought to do in other domains (e.g., (Burton et al., 2015)).
+However, these positive expectations do not always align with reality.
+Privacy safeguards are not consistently enforced or widely practiced in non-profit entities; for example, many health organizations rely on broad consent frameworks rather than adopting stricter purpose-restricted consent measures [121].
+This gap between expectations and reality can result in privacy expectation violations, a major source of privacy concerns as explained by the theory of contextual integrity [122].
+Indeed, a well-known non-profit mental health helpline shared it’s data with a commercial entity, leading to significant public outcry [123].
+Future work may seek to explore how to design privacy communications that set appropriate expectations in the presence of such effects, as a prior work has sought to do in other domains (e.g., [124]).
 […]
 Privacy Washing.Privacy WashingPrivacy Washing.
-Conversely, our participants’ had far lower initial expectations of the privacy practices of for-profit entities due to their perceived singular focus on profit (a “horn” effect).
+Conversely, our participants had far lower initial expectations of the privacy practices of for-profit entities due to their perceived singular focus on profit (a “horn” effect).
 As a result, information about data protection guarantees significantly raised their privacy expectations, and in turn, their willingness to donate their data.
-This means that for-profit entities have the opportunity to gain “competitive advantage” (Schaub et al., 2015) through influencing and enhancing user privacy expectations via transparently implementing and effectively communicating about PETs.
-However, this effect also posses a vulnerability to ”privacy washing” (MainWP, 2023; Cirucci, 2024), in which organizations make vague or exaggerated claims about PETs to falsely raise privacy expectations.
-Such deceptive practices can mislead individuals into sharing their data, erode trust, and discourage future data donations if individuals later feel deceived.
+This means for-profit entities can gain competitive advantage [125] by transparently implementing and communicating PETs to counter their trust deficit.
+Yet this dynamic creates vulnerability to “privacy washing” [126, 48], where organizations make vague or exaggerated privacy claims to deceptively raise expectations.
+A recent incident involving GoodRx exemplifies this problem: despite promising never to share personal health information, it shared sensitive data with third-party advertisers [127].
+Such deception ultimately erodes trust and discourages future donations when discovered, deepening the very horn effect that motivated privacy washing.
+[…]
+Opt in Versus Opt out.Opt in Versus Opt outOpt in Versus Opt out.
+Switching to an “opt-out” model for health data donation might seem appealing to overcome donation hesitancy, particularly for for-profit entities struggling against lower baseline trust.
+However, such approaches face legal barriers in the U.S.
+HIPAA [128] sets a high bar for health data sharing, requiring explicit “opt-in” authorization for most research uses.
 […]
 The Role of Auditing.The Role of AuditingThe Role of Auditing.
-In theory, auditing could offer verification that combats both the halo effect and privacy washing by grounding assumed or stated privacy guarantees in technical reality.
-Unfortunately, we find that auditing has a limited effect on elevating privacy expectations among participants, except in scenarios involving purpose use restrictions by for-profit entities.
-Participants understand that while PETs and auditing processes are in place, they are not infallible, and auditing does not substantially mitigate the practical risks associated with data breaches or unauthorized access.
-This perception shows a disconnect between the perspectives of experts and participants, as shown in (Dechand et al., 2019; Oates et al., 2018; Wu and Zappala, 2018).
-Experts often regard auditing as a fundamental assurance mechanism that reinforces trust and compliance.
-In contrast, participants are more concerned with the overall protection of their data, regardless of specific technical implementations or compliance processes.
-As one respondent noted, ”They can guarantee privacy all they want; things still get hacked.” True privacy is doubted to exist (Lerner et al., 2022).
+In theory, auditing could serve as a powerful mechanism to counterbalance both the halo effect and privacy washing by providing objective verification of privacy guarantees, regardless of organizational structure.
+Auditing should transform abstract guarantees into verifiable technical reality.
+Yet our findings reveal that auditing has surprisingly limited impact on elevating privacy expectations among participants, with the exception of scenarios involving purpose use restrictions and for-profit entities.
+This limited effectiveness arises from participants’ fundamental skepticism about technological infallibility., and shows a disconnect between expert and public perspectives on verification mechanisms, consistent with findings in prior work [32, 67, 34].
+While security and privacy experts typically regard auditing as foundational to establishing trust and demonstrating compliance, our participants takes it as merely another layer of protection that could ultimately fail.
+As one respondent succinctly stated, “They can guarantee privacy all they want; things still get hacked.” This skepticism aligns with emerging research suggesting that many users have developed a resigned attitude toward privacy, doubting that true privacy protection is achievable in practice [73].
 […]
-Building Trust.Building TrustBuilding Trust.
-Our respondent’s prior experiences with data leakage often informed this fatalistic perspective on privacy.
-Technologies like multi-party computation (MPC), although not yet scalable, have the potential to minimize the impact of a breach by specifically distributing sensitive information across different entities to ensure that no single entity can reconstruct the information on its own.
-However, deploying such systems requires careful engineering to prevent distributed data breaches on all the entities involved.
-Existing skepticism about data security means that any failure of these systems could severely damage trust, potentially more so than if simpler, well-understood methods were compromised.
+Recommendations for Recipient Entities.Recommendations for Recipient EntitiesRecommendations for Recipient Entities.
+Given participants’ skepticism about privacy protections, recipient entities seeking health data donations should consider several approaches to address these concerns.
+First, they should recognize that users’ prior experiences with data leakage and exposure to media coverage of breaches create a sense of inevitability about privacy risks.
+This fatalistic perspective requires more than just technical solutions.
+For technical implementations, recipient entities should consider advanced technologies like multi-party computation (MPC) that can minimize breach impacts by distributing sensitive information across multiple entities, ensuring no single point of failure.
+However, these systems must be engineered carefully, as any failure could damage trust more severely than if simpler methods were compromised.
+When communicating privacy guarantees, recipient entities should move beyond explaining basic techniques like E2EE.
+Our findings show that participants struggle to understand more complex concepts like auditing, despite their effectiveness.
+Recipient entities should develop communication approaches that better explain these sophisticated cryptographic concepts in accessible ways that demonstrate their benefits.
+For-profit entities in particular should recognize their trust disadvantage and focus on transparent implementation and clear communication of robust privacy protections.
+At the same time, they should avoid “privacy washing” through vague or exaggerated claims, which can erode trust if discovered.
+[…]
+Recommendations for End-Users.Recommendations for End-UsersRecommendations for End-Users.
+Our research highlights several key points that individuals should consider before donating their health data.
+First, donors should evaluate recipient entities based on the data protections they explicitly state in policy documents rather than their organizational structure.
+The halo effect observed in our study demonstrates how implicit assumptions about certain types of entities’ business practices can lead to unfounded assumptions about privacy practices, particularly for non-profit entities.
+Second, individuals should avoid assuming privacy protections exist beyond those explicitly stated by the entity requesting data.
+While entities are legally bound to follow both their explicit promises and applicable privacy laws, the baseline legal protections may be less comprehensive than what users expect for sensitive health data [129].
+Finally, individuals should consider engaging privacy advocates before donating sensitive health data.
+Such consultation can help identify potential privacy risks that might not be immediately obvious and provide a more balanced perspective on the true privacy implications of donation.
 […]
 Future Work.Future WorkFuture Work.
-In this work, we explicitly examined the influence of four privacy guarantees and two auditing guarantees.
-Future research could broaden this scope by incorporating a more diverse and comprehensive range of PETs, with a particular focus on understanding how multiple guarantees interact in real-world scenarios where they often coexist.
-Additionally, researchers can also conduct targeted studies to understand the perspectives of diverse demographic groups, including individuals with a strong distrust of data-handling practices.
+While our study provides valuable insights into privacy expectations and willingness to donate health data, several promising directions remain for future research.
+First, researchers could expand beyond our four privacy guarantees and two auditing mechanisms to incorporate a more comprehensive range of PETs, with a particular focus on understanding how advanced and complex guarantees can be effectively communicated to non-technical audiences and how they shape privacy expectations and willingness to donate.
+This could include exploring the interplay between different privacy guarantees and how they are perceived by individuals.
+Second, future work should examine these privacy expectations across more diverse demographic groups, including individuals with a strong distrust of data-handling practices.
 Identifying and understanding these individuals is critical in fostering confidence in PETs and practices.
-Our study focused exclusively on individuals’ perspectives regarding formal medical data donation scenarios.
-However, medical data are also frequently collected through smart devices and donated with a single click via in-app prompts.
-Prior work emphasizes the role of modality and timing in the design space of privacy information (Schaub et al., 2015).
-Future research could explore how such factors intersect with biases such as the halo effect to influence privacy expectations, trust, and willingness to share sensitive information.
-Finally, prior work (Seberger and Patil, 2021) studying COVID-19 contract tracing apps – many created through a partnership between for-profits and non-profits – found that people assumed these apps would “operate according to the economics of surveillance capitalism…[making it] unreasonable to expect them to trust the system that profits from the sale of their data.” Future work may seek to further interrogate how such partnerships, which are common in the medical domain, further complicate our understanding of people’s privacy expectations and their sensitivity to statements of data protection.
-Appendix A Survey
-In this section, we report the full text of our questionnaire.
-We first ask for a participant themselve or their close relatives have been diagnosed with a disease.
-In the rest of the questionnaire, we will choose at random one disease the participant has selected, if any, or one disease the participant has not selected.
-This is labeled in Italic as disease.
-The entity, also in Italic, may be either non-profit entity or for-profit entity, chosen at random.
-Egocentricity
-  1. (1)
-    Have you ever been diagnosed with any of these diseases?
-    • Cancer (Breast, Colorectal, Lung, and Prostate)
-    • Diabetes
-    • Heart Failure
-    • Hypertension (High blood pressure)
-    • Stroke
-    • None of the above
-  2. (2)
-    Have your partner and/or close relatives (parents, children, siblings) ever been diagnosed with any of these diseases?
-    • Cancer (Breast, Colorectal, Lung, and Prostate)
-    • Diabetes
-    • Heart Failure
-    • Hypertension (High blood pressure)
-    • Stroke
-    • None of the above
-Survey Introduction
-Next, we present you with a scenario and ask you how you would respond.
-Please do your best to imagine yourself in this scenario.
-There is no right or wrong answer to these questions: we are interested to hear how you would feel about being in this scenario and what you would do.
-Imagine that an entity entity wants to develop a new treatment for disease.
-They need medical data from people with and without disease to develop the treatment.
-They ask you to donate your medical record to help develop the treatment.
-Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
-Questions
-We will now ask you a series of questions about this scenario. For your reference, the scenario text will appear at the top of each page.
-  1. (1)
-    How would you rate your understanding of the above scenario?
-    • Fully understand
-    • Mostly understand
-    • Partially understand
-    • Not understand
-  2. (2)
-    In this scenario how likely would you be to donate your medical record?
-    • Very likely
-    • Likely
-    • Unlikely
-    • Very unlikely
-  3. (3)
-    Why your are willing (or unwilling) to share your medical record with this entity?
-  4. (4)
-    Please rate your agreement with the following statement:
-    I trust the entity will handle my data as described.
-    • Strongly agree
-    • Somewhat agree
-    • Somewhat disagree
-    • Strongly disagree
-  5. (5)
-    Please explain why you do (or do not) trust that the entity will handle your data as described.
-  6. (6)
-    In this scenario, how likely do you think it is that each of the following will occur?
-    • My full name or other personal identifiable information will be linked to the donated medical record.
-      • Very likely
-      • Likely
-      • Unlikely
-      • Very unlikely
-    • The donated medical record will be deleted at a set point in time.
-      • Very likely
-      • Likely
-      • Unlikely
-      • Very unlikely
-    • Any employee at the recipient entity will be able to access the donated medical records.
-      • Very likely
-      • Likely
-      • Unlikely
-      • Very unlikely
-    • The donated medical records will be used for another purpose without my consent.
-      • Very likely
-      • Likely
-      • Unlikely
-      • Very unlikely
-    • A group of independent experts will verify whether the privacy-preserving technology works and publish a report on their findings.
-      • Very likely
-      • Likely
-      • Unlikely
-      • Very unlikely
-    • I will be able to hire someone to verify that my medical record is protected as described.
-      • Very likely
-      • Likely
-      • Unlikely
-      • Very unlikely
-    • If I donate my data, I will meet Albert Einstein.
-      • Very likely
-      • Likely
-      • Unlikely
-      • Very unlikely
-  7. (7)
-    History In real life, have you ever donated your medical record?
-    • Yes, I have donated my medical record for disease
-    • Yes, I have donated my medical record for research of other illnesses.
-    • Yes, I have donated my medical record for non-medical purposes.
-    • No, I was asked to donate my medical record but I declined to participate.
-    • No, I have never been asked to donate my medical record.
-    • I prefer not to answer.
-Internet Skills
-  1. (1)
-    How familiar are you with the following terminology? Please choose a number between 1 and 5 where 1 represents “no understanding” and 5 represents “full understanding” of the term.
-    • Advanced search
-    • PDF
-    • Spyware
-    • Wiki
-    • Cache
-    • Phishing
-  2. (2)
-    Which of the following best describes your educational background or job field?
-    • I have an education in, or work in, the field of computer science, computer engineering or IT
-    • I DO NOT have an education in, nor do I work in, the field of computer science, computer engineering or IT.
-    • Prefer not to answer.
-  3. (3)
-    Have you ever heard any of the following techniques? (select all that apply)
-    • Differential Privacy
-    • End-to-end encryption
-    • Public-key cryptography
-    • Trusted execution environment (TEE)
-    • Pseudo-party computation
-    • None of the above
-Demographics
-  1. (1)
-    What year were you born in? (4 digits)
-  2. (2)
-    To which gender do you most identify? (Select all that may apply)
-    • Woman
-    • Man
-    • Non-binary
-    • Prefer not to answer
-    • Prefer to self-describe
-  3. (3)
-    What is the highest level of education you have completed?
-    • Less than high school degree (1)
-    • High school graduate (high school diploma or equivalent including GED) (2)
-    • Some college but no degree (3)
-    • Associate’s degree (4)
-    • Bachelor’s degree (5)
-    • Advanced degree (e.g., Master’s, doctorate) (6)
-    • Prefer not to answer (7)
-Appendix B Code Book
-[…]
-Data Donation.Data DonationData Donation.
-  1. (1)
-    Irrelevant response
-  2. (2)
-    Opt to donate regardless
-  3. (3)
-    Opt not to donate regardless
-  4. (4)
-    Require compensation
-  5. (5)
-    Only donate to a non-profit entity
-  6. (6)
-    No donation to a for-profit entity
-  7. (7)
-    Need more information
-  8. (8)
-    The provided protection is persuasive
-  9. (9)
-    Generally, the donation scenario looks good
-  10. (10)
-    Not satisfied with the provided protection
-[…]
-Trustworthy.TrustworthyTrustworthy.
-  1. (1)
-    Trust without a specific reason
-  2. (2)
-    Irrelevant response
-  3. (3)
-    Perfect protection does not exist
-  4. (4)
-    The recipient entity has no reason not to protect the data
-  5. (5)
-    The recipient entity has no reason to protect the data
-  6. (6)
-    A non-profit entity is trustworthy
-  7. (7)
-    A for-profit entity is not trustworthy
-  8. (8)
-    Need more information
-  9. (9)
-    The provided protection is persuasive
-  10. (10)
-    Generally, the donation scenario looks good
-  11. (11)
-    Not satisfied with the provided statement
+Third, our study focused on formal medical data donation scenarios, yet health data is increasingly collected through passive means via consumer devices and casual in-app permissions.
+As prior work emphasizes the importance of modality, timing, and context in privacy information design [125], future research could explore how such factors intersect with biases such as the halo effect to influence privacy expectations, trust, and willingness to share sensitive information.
+Finally, the increasingly common partnerships between for-profit and non-profit entities in the medical domain warrant dedicated study.
+Building on prior work [52] that found deep skepticism toward hybrid entities in COVID-19 contact tracing, future work may seek to further interrogate how such partnerships, which are common in the medical domain, further complicate our understanding of people’s privacy expectations and their sensitivity to statements of data protection.
+Acknowledgment
+All authors were supported in full or in part by the Max Planck Society during the course of this research.
+The fourth author additionally acknowledges support from a Google Research Scholar Award “Aligning Technical Data Privacy Protections with User Concerns” and NSF Award # 2429838.